### v1

# Authorship Verification based on the Likelihood Ratio of Grammar Models

Abstract
Authorship verification (AV) is the process of analyzing a set of documents to determine whether they were written by a specific author. This problem often arises in forensic scenarios, e.g., in cases where the documents in question constitute evidence for a crime. Existing state-of-the-art AV methods use computational solutions that are not supported by a plausible scientific explanation for their functioning and that are often difficult for analysts to interpret. To address this, we propose a method relying on calculating a quantity we call $\lambda_{G}$ (LambdaG): the ratio between the likelihood of a document given a model of the Grammar for the candidate author and the likelihood of the same document given a model of the Grammar for a reference population. These Grammar Models are estimated using $n$-gram language models that are trained solely on grammatical features. Despite not needing large amounts of data for training, LambdaG still outperforms other established AV methods with higher computational complexity, including a fine-tuned Siamese Transformer network. Our empirical evaluation based on four baseline methods applied to twelve datasets shows that LambdaG leads to better results in terms of both accuracy and AUC in eleven cases and in all twelve cases if considering only topic-agnostic methods. The algorithm is also highly robust to important variations in the genre of the reference population in many cross-genre comparisons. In addition to these properties, we demonstrate how LambdaG is easier to interpret than the current state-of-the-art. We argue that the advantage of LambdaG over other methods is due to fact that it is compatible with Cognitive Linguistic theories of language processing.
1 Introduction
Texts serve numerous purposes and can be found in a variety of digital and non-digital forms such as books, social media posts, emails, letters, websites, academic papers, blogs, product reviews, poems and many others. These text types, in turn, can be categorized according to various criteria including language, genre, topic, readability or authorship. The latter is particularly important when it comes to investigating questioned documents, for example, in forensic settings, e. g., blackmail letters, suicide notes, letters of confession, wills or theses potentially written by ghostwriters.
Over the course of time, various research disciplines have emerged to deal with the analysis of the language of authors from different perspectives. This has led to an interdisciplinary branch of research that is often referred to as digital text forensics in the literature (e. g., [90, 10, 40, 95, 38]). Two well-known and closely related sub-areas of this branch of research are authorship attribution (AA) and authorship verification (AV). AA is concerned with the task of attributing an anonymous text to the most likely author based on a set of sample documents of known authors. AV, on the other hand, deals (in its simplest form) with the task of deciding whether two documents were written by the same author. Our focus in this paper is on AV, which represents a reformulation of AA, based on the fact that any AA problem can be decomposed into a series of AV problems [67].
1.1 Decision Problems
According to Stein et al. [105], Juola and Stamatatos [55], Koppel et al. [67] and Halvani [41] the following decision problems can be answered by AV:
  1. 1.
    $\textrm{AV}_{\,\textrm{Core}}$: Given two documents $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$: Were both documents written by the same author?
  2. 2.
    $\textrm{AV}_{\,\textrm{Batch}}$: Given two sets of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}^{\,(a)}_{1},\mathcal{D}^{\,(a)}_{2},\ldots\}$ and $\mathbb{D}_{\mathcal{B}}=\{\mathcal{D}^{\,(b)}_{1},\mathcal{D}^{\,(b)}_{2},\ldots\}$ each of which written by a single author: Are the documents in $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{D}_{\mathcal{B}}$ written by the same author?
  3. 3.
    $\textrm{AV}_{\,\textrm{Known}}$: Given a set of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots\}$ by the same author $\mathcal{A}$ as well as a document $\mathcal{D}_{\mathcal{U}}$ of an unknown author $\mathcal{U}$: Has $\mathcal{A}$ also written $\mathcal{D}_{\mathcal{U}}$ such that $\mathcal{A}=\mathcal{U}$ holds?
In a forensic context the process of assisting the trier-of-facts in finding answers to these problems is referred to as Forensic Text Comparison (FTC) [49]. The focus in this paper is on all three of these decisions problems. For the rest of this paper, we adopt the following notation: a document denoted by $\mathcal{D}_{\mathcal{U}}$, represents a document of an unknown author $\mathcal{U}$, whereas a document denoted by $\mathcal{D}_{\mathcal{A}}$ represents a document of a known author $\mathcal{A}$. In all decision problems, including $\textrm{AV}_{\,\textrm{Core}}$ where the denotation of which document is known or unknown can be arbitrary, an AV method seeks to answer the question of whether $\mathcal{A}=\mathcal{U}$ ($\mathcal{A}$ authored $\mathcal{D}_{\mathcal{U}}$) or $\mathcal{A}\neq\mathcal{U}$ ($\mathcal{A}$ did not author $\mathcal{D}_{\mathcal{U}}$) holds. In this context, a known verification case with a matching and non-matching authorship is denoted by Y- and N-case, respectively. Moreover, we denote $\neg\mathcal{A}$ as any author other than $\mathcal{A}$.
1.2 Applications of Authorship Verification
For all the mentioned forms of decision problems, numerous AV applications can be found across different domains. In forensic linguistics, for example, AV can be used to analyze malicious texts such as ransom notes, blackmail letters or other non-malicious texts that are used as evidence in investigatory, criminal or civil contexts. In the field of journalism and media, AV can play a vital role when it comes to confirming the authorship of news articles or opinion pieces or investigating cases of ghostwriting and unauthorized publication of content. In academia, AV can be used to analyze research articles, dissertations or other written content, for example, to resolve a plagiarism claim, or contract cheating from students, and thereby overall contribute to combating academic misconduct. In the context of social media, AV can also be employed to detect fake accounts or multiple accounts belonging to the same user. Moreover, as part of email and communication analysis, AV can be used to verify the authorship of emails in cases of fraud, harassment or cybercrime. Finally, in the Digital Humanities, AV can be used to resolve matters that deal with disputed historical evidence or to provide evidence in favour or against the origin of literary works.
1.3 Categories of AV Methods
According to Halvani et al. [43, 47], AV methods can be divided into three possible categories, namely unary, binary-intrinsic and binary-extrinsic AV methods [^1] A summarized visual representation of these three categories can be seen in [47, Figure 1]. All three rely on a decision criterion $\theta$, in order to accept or reject $\mathcal{A}$ as the author of $\mathcal{D}_{\mathcal{U}}$. The way in which $\theta$ is determined is what primarily distinguishes these three categories:
  • Unary AV methods rely solely on a set of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots\}$ of a known author $\mathcal{A}$ in order to determine $\theta$. These methods assume $\mathcal{D}_{\mathcal{U}}$ to be written by $\mathcal{A}$, if it is similar enough to the documents in $\mathbb{D}_{\mathcal{A}}$, where the decision is only based on these and $\theta$.
  • Binary-intrinsic AV methods require a given training corpus, consisting of verification cases with a ground truth regarding the Y- (same-authorship) and N-classes (different-authorship), in order to determine $\theta$. Binary-intrinsic AV methods treat verification cases that include the unknown and known document(s) of $\mathcal{A}$, as a single unit $X$ (e. g., a feature vector). If $X$ is more similar to the Y-cases, the method accepts $\mathcal{A}$ as the author of $\mathcal{D}_{\mathcal{U}}$. On the other hand, if $X$ is more similar to the N-cases, $\mathcal{D}_{\mathcal{U}}$ is assumed to be written by another author. In any case, the decision is made solely on the basis of $X$ and the training classification model that incorporates $\theta$.
  • Binary-extrinsic AV methods require external, so-called impostor [70] documents (which have not been written by $\mathcal{A}$) in order to determine $\theta$. To obtain impostor documents $\mathbb{I}=\{I_{1},I_{2},\ldots\}$, different strategies can be employed. Koppel and Winter [70], for example, used two approaches: (1) collecting documents using search engine queries and (2) using a fixed set of already crawled documents. Binary-extrinsic AV methods assume $\mathcal{D}_{\mathcal{U}}$ to be written by $\mathcal{A}$ only if $\mathcal{D}_{\mathcal{U}}$ is consistently more similar to $\mathcal{D}_{\mathcal{A}}$ (or alternatively, the documents in $\mathbb{D}_{\mathcal{A}}$) despite the presence of distractor impostor documents $\{I_{1},I_{2},\ldots\}$ after calculating a similarity score using a bootstrapping approach.
In a forensic context, the adherence to the likelihood ratio paradigm for forensic science implies that an AV method does not return $\theta$ but instead returns a likelihood ratio that quantifies the strength of the evidence in favour or against the two propositions. In addition to the three types above, therefore, it is also possible to classify an AV method as forensic when the method returns a likelihood ratio.
In a review of the field, Stamatatos [102] concluded that binary-extrinsic methods tend to outperform the other categories. In particular, a binary-extrinsic AV method that has been found to be very successful over several independent studies is the well-known Impostors Method (IM). The basic idea behind IM can be traced back to another successful AV method introduced by Koppel and Schler [65] and called Unmasking and later modified by Koppel et al. in [68]. The key insight of IM is to collect a set of texts that are plausible impostors for $\mathcal{D}_{\mathcal{A}}$ and then calculate a similarity score for $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ not once but a large number of times while withholding a random sub-set of the features at every iteration. The rationale is that only the real author of the disputed text is so similar to it to withstand random permutations of the features set, which, for IM, is the set of the most frequent character 4-grams. Koppel et al. [67] proposed that to choose the impostors the key principle should be to identify a relevant universe, choose linguistically similar texts, and then randomly sample from them.
IM found immediate external validation in the PAN 2013 competition, where a variant called GenIM (General Impostors Method), proposed by Seidman [98], outperformed all other submitted AV methods. As suggested by Koppel et al. [66], GenIM made use of collected impostors document from the web. The method then used the relative frequency of single words weighted by their document frequency (a quantity usually called tf-idf) and the Min-Max (aka Ružička or fuzzy Jaccard) similarity. Successful results were then found by Koppel and Winter [70] for the $\textrm{AV}_{\,\textrm{Core}}$ decision problem, again using the Min-Max similarity between vectors of character 4-grams tf-idf. For the impostors, Koppel and Winter [70] tried various configurations, including with impostors generated on the fly using web searches or sampling from genre-compatible corpora. They discovered that genre compatibility permits the use of fewer impostors but that, equally, a large number of impostors ($\approx 200$) generated using web searches can compensates for lack of genre control. Their results demonstrated that it is possible to solve this challenging two-text only verification problem even for texts as short as 500 words.
[…]
Although IM’s performance over the years is unquestionable, the four main drawbacks of the method are that it is not deterministic, that it suffers from a high run-time [41], that it relies on choosing impostors from the same genre, and that its results are often hard to interpret. The first issue is typically addressed by running several iterations and averaging the results but this additional procedure only leads to even longer run-times. The third issue is not strictly speaking an issue of IM but a well-known problem for both AV and AA, as virtually all methods to solve theses tasks tend to be highly affected by genre variation [5]. Regarding IM, Halvani [41] found that the performance significantly decreases when the set of impostors is chosen from another genre. For example, performing AV on a corpus of chat logs using a corpus of chat logs leads to a performance of 0.92 in accuracy. However, when using business emails or academic articles as impostors, this performance drops to almost chance level at 0.54. Only when the genres are relatively similar (e. g., chat logs compared to blogs) then, although a significant drop in performance is still evident, the results are still above chance (0.74).
The last and most significant limitation of IM and of any other AV methods using the frequency of short character $n$-grams is that the most important features of the analysis are not immediately obvious to the human analyst. In addition, unless some kind of content masking is implemented, the use of short character $n$-grams can also lead to the contamination of the data from topic or content noise and other unwanted idiosyncrasies that could influence the results. For example, a character 4-gram like “f th” is likely to capture grammatical information, such as the frequencies of grammatical sequences like “if the” or “of the” or “if that” but also information that has to do with the topic of a text, like “elf theme” or “grief therapy”. The danger of capturing content is that topics might correlate to authors and, especially in forensic settings, the AV method could, for example, return a positive verification outcome only because in the corpus used for training there is a correlation between content and authors (e. g., $\mathcal{A}$ is likely to be the author of $\mathcal{D}_{\mathcal{U}}$ because $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ are about “grief therapy”).
1.4 Effect of Content
This issue of the effect of content is one of a number of biases identified by Bevendorff et al. [11] on the basis of the 2014 and 2015 PAN competitions. These biases were taken in consideration by PAN starting from its 2020 edition [57]. However, the problem of using features that can potentially capture topic or content instead of authorship remains in many verification systems, including more recent ones.
One of the first approaches dealing with content bias in the context of authorship analysis was the topic-masking technique TextDistortion [103], which is simply based on replacing any word that is not among the top most frequent words in a reference corpus with an asterisk. TextDistortion was found to actually improve the results in cross-topic and cross-genre analyses. In 2021, Halvani and Graner [42] tested TextDistortion against another algorithm they introduced called POSNoise, which is a pre-processing method that consists in the replacement of words that are unlikely to be functional with their Part-Of-Speech (POS). They found that this method clearly improves the algorithms in cross-domain cases while often resulting in poorer performance in cases that are not cross-domain. These results are interpreted by them as evidence that this algorithm should be applied for forensic purposes because it is capable of removing the effect of content. Among the methods that benefited from this pre-processing was IM, which was also one of the most successful ones in their evaluations. A similar feature type to avoid content bias is the use of frame $n$-grams [82]. Similarly to POSNoise, this feature set consists of all function words and with all nouns, verbs, adjectives, adverbs, and pronouns being replaced with their corresponding POS.
As an alternative to these rule-based content removal methods, Bischoff et al. [13] proposed a solution using a neural network trained to differentiate content from style. In their study, they confirmed that the frequency of short unprocessed character $n$-grams is highly affected by content by analyzing a large corpus of cross-domain problems where they found that methods based on these features perform at random. Despite the success of their own neural network method, they discovered that a simpler approach such as TextDistortion works equally well at a fraction of the computational and data cost of a neural network.
Another approach to counteract content bias, called TAVeer, was proposed by Halvani et al. [44]. In contrast to TextDistortion and POSNoise, TAVeer represents a stand-alone AV method based on ensembles of topic-agnostic (TA) feature categories (namely punctuation $n$-grams, TA sentence and sentence-initial tokens, TA sentence endings, TA token $n$-grams and TA masking token $n$-grams). The TA feature categories themselves cover a total of $\approx 1000$ words and phrases that originate from 20 linguistic categories such as punctuation marks, conjunctions, determiners, prepositions, contractions, auxiliary verbs, transition words and others, which are known in the literature [86, 12, 106] to be content and topic independent.
1.5 Language Models and Compression Methods
In addition to IM and its variants, the most successful family of methods in AV are the ones that are based on language modeling and/or compression, which are mathematically related. A language model is here defined as a generative statistical model that assigns a probability distribution to stretches of language.
One of the very first successful examples for this class of methods was the winner of the PAN 2015 competition, Bagnall [8], whose AV method was based on a character-level recurrent neural network language model. Another notable neural network language modeling approach is Boenninghoff’s et al. [14] use of a Siamese neural network. Their approach was found to outperform IM by halving its error rate while also offering a way to visualize the importance of tokens. Their approach was designed to mitigate the effect of content but its limitation is that a large and structured comparison dataset is necessary, which is not always available. King and Cook [61] also proposed an approach based on language models which works by retraining the final layer of a model based on potentially short samples from an author and then comparing this model against a general model on the questioned text. They found that this system outperformed the use of traditional machine learning methods like Support Vector Machines. Although language models are also inevitably affected by the issue of content bias, attempts to mitigate this problem are scarce. A notable mention is Fourkioti’s et al. [37] approach consisting in training an $n$-gram language model on sequences of POS tags. The authors found that this kind of model is only valuable when fused with other features.
Related to language modeling, the other approach that has been found to be very successful is the use of compression algorithms. This family of methods is fundamentally very similar to language models because they both are probabilistic models of a training dataset. Language modeling can be thought of as a process of compression and one of the best compressor for text data, prediction by partial matching (PPM), could also be seen as a character-level language model. A particularly successful AV method which makes use of PPM as an underlying engine is COAV [46, 45]. The method was found to perform similarly to IM, but in a much shorter time and without relying on impostors. Halvani’s 2021 survey of the field [41] indeed concluded that COAV, together with IM, are the best AV methods. Both of them can be used together with POSNoise to avoid content bias. Although IM tended to be more robust over the various tests it was subjected to, the advantage of COAV over IM is that the method is deterministic and much faster (for comparison see, for example, [46, Table 4]). In addition, Halvani [41] was able to visualize the importance of character sequences for the prediction by producing a color-coded version of the texts, thus making the method more qualitatively explorable by the analyst.
2 The Likelihood Ratio of Grammar Models
The method proposed in this paper is a binary-extrinsic AV method influenced by the literature above which offers a general theoretically-motivated framework for AV that can help to better explain the other methods reviewed above. This section explains its theoretical framework and the algorithm, which is outlined in Algorithm 1, while most technical details are left to Section 5.
Input: $\mathbb{D}_{\mathcal{U}}=\{\mathcal{D}_{\mathcal{U}}^{1},\mathcal{D}_{\mathcal{U}}^{2},\dots\}$, $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}_{\mathcal{A}}^{1},\mathcal{D}_{\mathcal{A}}^{2},\dots\}$, $\mathbb{D}_{\text{ref}}=\{\mathcal{D}_{\text{ref}}^{1},\mathcal{D}_{\text{ref}}^{2},\dots\}$, $N$, $r$
1 Function sample($k$, $\mathbb{S}$):
2       Randomly sample a subset of $k$ sentences from the input set $\mathbb{S}$
3Function sent($\mathbb{D}$):
4       Construct a set $\mathbb{S}$ of tokenized sentences from the input set $\mathbb{D}$
// Apply POSNoise to each document set
$\mathbb{S}_{\mathcal{U}}\mbox{}\leftarrow\textnormal{{sent}}(\textsf{POSNoise}\mbox{}(\mathbb{D}_{\mathcal{U}}))$ $\mathbb{S}_{\mathcal{A}}\mbox{}\leftarrow\textnormal{{sent}}(\textsf{POSNoise}\mbox{}(\mathbb{D}_{\mathcal{A}}))$ $\mathbb{S}_{\textsf{ref}}\mbox{}\leftarrow\textnormal{{sent}}(\textsf{POSNoise}\mbox{}(\mathbb{D}_{\text{ref}}))$ // Build Grammar Model for the known author $\mathcal{A}$
$G_{\mathcal{A}}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\mathcal{A}},N)$ // Build reference Grammar Models
5 for $i\leftarrow 1$ to $r$ do
6       $\mathbb{S}_{\text{i}}\leftarrow$ sample($|\mathbb{S}_{\mathcal{A}}|$, $\mathbb{S}_{\text{ref}}$) $G_{i}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\text{i}},N)$
// Calculate $\lambda_{G}$ (the log-likelihood ratio of the Grammar Models) over $\mathbb{S}_{\mathcal{U}}\mbox{}$
7 $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow 0$ for $S\mbox{}\in\mathbb{S}_{\mathcal{U}}$ do
       // Decompose $S\mbox{}$ into a sequence of tokens
8       $(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{m})\leftarrow S\mbox{}$ for $j\leftarrow 1$ to $m$ do
             // Calculate mean $\lambda_{G}$ for $t\mbox{}_{j}$ over reference Grammar Models
9             for $i\leftarrow 1$ to $r$ do
10                   $c\leftarrow(t\mbox{}_{j-N+1},\ldots,t\mbox{}_{j-1})$ $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})+\frac{1}{r}\log\frac{P(t\mbox{}_{j}|c;G_{\mathcal{A}})}{P(t\mbox{}_{j}|c;G_{i})}$
11
12
return $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})$
Algorithm 1 Likelihood Ratio of Grammar Models (LambdaG)
[…]
$P(S\mbox{};G)=\prod_{j=1}^{m}P(t\mbox{}_{j}|t\mbox{}_{j-N+1},t\mbox{}_{j-N+2},\ldots,t\mbox{}_{j-1};G)$ (1)
For readability reasons, we rewrite the probability assigned by $G$ to each token $t\mbox{}$ in its context $c$ as $P(t\mbox{}|c;G)$ (as in line 18 in Algorithm 1, where $t\mbox{}_{j}$ is used instead of $t$). We estimate Grammar Models adopting Kneser-Ney smoothing [62, 26], represented in Algorithm 1 as $\textnormal{{KN}}()$, with a fixed single default discount parameter $D=0.75$, as explained in Section 5.3. We set $N=10$ as an order of the model to ensure that all grammatical relations among the tokens of a sentence are captured, even in cases of long sentences. However, the order of the model $N$ could be considered as a hyperparameter of the algorithm (lines 8 and 11 in Algorithm 1).
Assuming the Principle of Linguistic Individuality that at any time $\tau$, for any language $L_{\tau}$, there do not exist two individuals, $\mathcal{A}$ and $\mathcal{B}$, for whom their grammar is identical [82], we expect that in a large number of cases the two Grammar Models $G_{\mathcal{A}}$ and $G_{\mathcal{B}}$ would assign two different probabilities to the same token in context. $P(t\mbox{}|c;G_{\mathcal{A}})$ can be interpreted as the extent to which token $t\mbox{}$ is ’grammatical’ or, according to Cognitive Linguistic theories, entrenched for author $\mathcal{A}$ when in context $c$. For this reason, we would expect that, if the sequence $s=t\mbox{}_{1},t\mbox{}_{2},t\mbox{}_{3}$ was produced by $G_{A}$, then $P(s;G_{\mathcal{A}})>P(s;G_{\mathcal{B}})$ or, more generally according to the Principle of Linguistic Individuality, $P(s;G_{\mathcal{A}})>P(s;G_{\neg\mathcal{A}})$. The model $G_{\mathcal{A}}$ is estimated from $\mathbb{S}_{\mathcal{A}}\mbox{}$ (line 8 in Algorithm 1). Instead, to generate an approximation to $G_{\neg\mathcal{A}}$ we estimate a set of reference Grammar Models $\mathbb{S}_{\textsf{ref}}\mbox{}=\{G_{1},G_{2},\dots,G_{r}\}$ where each $G_{r}$ is generated by uniformly sampling randomly from $\mathbb{S}_{\textsf{ref}}\mbox{}$ a set of sentences $\mathbb{S}_{r}$ of the same size as the set of sentences in the candidate author’s corpus, $\mathbb{S}_{\mathcal{A}}\mbox{}$, so that $|\mathbb{S}_{r}|=|\mathbb{S}_{\mathcal{A}}\mbox{}|$ (line 10 in Algorithm 1). We set $r=100$ but this value can also be considered a hyperparameter of the algorithm. A comparison of hyperparameter settings is reported in Section 3.
Taking an approach consistent with the likelihood ratio framework adopted in forensic science, similarly to Ishihara [48], it is therefore possible to calculate (lines 16-18 in Algorithm 1) a mean log-likelihood ratio of the probability assigned to each token in context by $G_{\mathcal{A}}$ and by each $G_{r}$, which we call $\lambda_{G}$:
$\lambda_{G}(t\mbox{}|c)=\frac{1}{r}\sum_{i=1}^{r}{\log\frac{P(t\mbox{}|c;G_{\mathcal{A}})}{P(t\mbox{}|c;G_{i})}}$ (2)
In Eq. 2 the numerator can be seen as a measure of similarity, how much the token is grammatical or entrenched for $G_{\mathcal{A}}$ given its context, while the denominator is a measure of typicality, how ’ungrammatical’ or atypical this token in context is in the reference population. A value of $\lambda_{G}$ for each $S\mbox{}\in\mathbb{S}_{\mathcal{U}}$ can then be calculated (line 14-18 in Algorithm 1) as:
$\lambda_{G}(S\mbox{})=\sum_{j=1}^{m}\lambda_{G}(t\mbox{}_{j}|c)$ (3)
Since we model each $S\mbox{}\in\mathbb{S}_{\mathcal{U}}\mbox{}$ as independent from each other, we can then calculate (line 12-18 in Algorithm 1):
$\begin{split}\lambda_{G}(\mathcal{D}_{\mathcal{U}})=\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})=\sum_{S\mbox{}\in\,\mathbb{S}_{\mathcal{U}}\mbox{}}\lambda_{G}(S\mbox{})=\sum\limits_{S\mbox{}\,\in\,\mathbb{S}_{\mathcal{U}}\mbox{}}\sum\limits_{j=1}^{m}{\frac{1}{r}\sum\limits_{i=1}^{r}\log\frac{P(t\mbox{}_{j}|c;G_{\mathcal{A}})}{P(t\mbox{}_{j}|c;G_{i})}}\end{split}$ (4)
An advantage of this formula is that the final $\lambda_{G}$ can be decomposed into $\lambda_{G}$ values for each single sentence and each single token in each single sentence. This means that the analyst can reconstruct step by step what features contribute to the final score, for example, by ranking each sentence for importance and, for each sentence, color-code or highlight the most important tokens.
In addition to this property, the $\lambda_{G}$ value is also effectively an uncalibrated log-likelihood ratio score that includes information about both similarity and typicality and that can then be calibrated into a forensic log-likelihood ratio using a calibration method such as a logistic regression. We denote a calibrated $\lambda_{G}$ log-likelihood ratio that expresses the strength of evidence in a forensic context with the upper case lambda ($\Lambda_{G}$). To perform this calibration we use a training corpus containing Y- and N-cases to estimate a logistic regression calibration model [110, 49] that turns each $\lambda_{G}$ into $\Lambda_{G}$. Here, $\Lambda_{G}$ quantifies the strength of the linguistic evidence for each of the two alternative forensic propositions. More details on the Likelihood Ratio Framework for forensic science and its evaluation is described in Section 5.4. Based on $\Lambda_{G}$, we classify a verification case as Y if $\Lambda_{G}>0$, else as N.
Finally, we note that another important aspect that is not discussed in detail in this section is the fact that several of the key components of LambdaG are consistent with Cognitive Linguistics theories. More details about this aspect of LambdaG are left to Section 4.
3 Results
Algorithm 1 was applied to a set of twelve corpora that simulate various real-life AV problems and scenarios, described in Section 5. The full results are presented in Table 1, where the performance of each method is evaluated using five standard performance measures for classification: Accuracy, the Area Under the Curve (AUC), F1, Precision, and Recall. These are values that were calculated using a threshold $\theta$, which is derived from $\Lambda_{G}$. The split between training and test datasets for each corpus tested is shown in Table 4. To assess the performance of $\Lambda_{G}$, instead, the $C_{\text{llr}}\mbox{}$ measure is used, which expresses the amount of information in the likelihood ratio. This metric is explained in Section 5.4.
In Table 1, LambdaG is compared against four well-known and tested AV methods: IM, COAV, TAVeer and a neural language model using a Siamese network, SiamBERT, which is explained in Section 5.5. IM and COAV were run on POSNoise pre-processed corpora and they are therefore content-agnostic, similarly to LambdaG. TAVeer was run on the unprocessed corpora but because the method only measures topic-agnostic features, this is also comparable to LambdaG. Since SiamBERT uses a pre-trained model on data that it is not POSNoise-processed, fine-tuning it on POSNoise-processed data would be an unfair comparison. Equally, however, SiamBERT fine-tuned on unprocessed data risks using content information and is also, in a different way, an unfair comparison. For this reason, we present the results of both versions. The results of SiamBERT on POSNoise-processed data is indicated with ${}^{\,(\bm{\dagger})}$.
The results in Table 1 show that LambdaG is superior to the other AV topic-agnostic baselines in all cases. LambdaG is also the best performing method when compared to SiamBERT despite the fact that any correlation between author and topic could have been exploited in each corpora. The only exception is for $\mathcal{C}_{\mathrm{AllNews}}$, where instead LambdaG comes second to SiamBERT. We do not have an explanation for this result but we advance the plausible hypothesis that $\mathcal{C}_{\mathrm{AllNews}}$ contains data with a significant author/topic bias such that the task is made easier when content words are included. This hypothesis is supported by the extremely poor performance of ${}^{\,(\bm{\dagger})}$ on $\mathcal{C}_{\mathrm{Stack}}$, which is a corpus explicitly designed to penalize AV methods that exploit topic information.
Table 1: Evaluation results of LambdaG and the four baseline methods. ${}^{\,(\bm{\dagger})}$ indicates that the method was applied to the original (i. e., non-POSNoise-processed) corpora. In terms of accuracy and AUC, bold and underlined values represent the best and second-best results, respectively.
Method Acc. AUC F1 Prec. Rec. TP FN FP TN
LambdaG 0.906 0.938 0.903 0.875 0.933 42 6 3 45
IM 0.823 0.920 0.817 0.792 0.844 38 10 7 41
COAV 0.750 0.827 0.760 0.792 0.731 38 10 14 34
TAVeer 0.729 0.842 0.735 0.750 0.720 36 12 14 34
SiamBERT 0.646 0.673 0.721 0.595 0.917 44 4 30 18
$\bm{\mathcal{C}_{\mathrm{Enron}}}$ ${}^{\,(\bm{\dagger})}$ 0.750 0.841 0.778 0.700 0.875 42 6 18 30
LambdaG 0.876 0.938 0.872 0.841 0.905 95 18 10 103
IM 0.854 0.927 0.862 0.912 0.817 103 10 23 90
COAV 0.801 0.912 0.793 0.761 0.827 86 27 18 95
TAVeer 0.788 0.858 0.789 0.796 0.783 90 23 25 88
SiamBERT 0.580 0.627 0.676 0.550 0.876 99 14 81 32
$\bm{\mathcal{C}_{\mathrm{Wiki}}}$ ${}^{\,(\bm{\dagger})}$ 0.757 0.909 0.799 0.681 0.965 109 4 51 62
LambdaG 0.873 0.933 0.864 0.807 0.929 92 22 7 107
IM 0.689 0.738 0.684 0.675 0.694 77 37 34 80
COAV 0.561 0.640 0.554 0.544 0.564 62 52 48 66
TAVeer 0.693 0.763 0.685 0.667 0.704 76 38 32 82
SiamBERT 0.605 0.648 0.643 0.587 0.711 81 33 57 57
$\bm{\mathcal{C}_{\mathrm{Stack}}}$ ${}^{\,(\bm{\dagger})}$ 0.535 0.576 0.579 0.529 0.640 73 41 65 49
LambdaG 0.782 0.861 0.775 0.75 0.802 105 35 26 114
IM 0.775 0.851 0.755 0.693 0.829 97 43 20 120
COAV 0.750 0.835 0.751 0.757 0.746 106 34 36 104
TAVeer 0.707 0.748 0.689 0.650 0.734 91 49 33 107
SiamBERT 0.593 0.641 0.639 0.574 0.721 101 39 75 65
$\bm{\mathcal{C}_{\mathrm{ACL}}}$ ${}^{\,(\bm{\dagger})}$ 0.607 0.664 0.654 0.584 0.743 104 36 74 66
LambdaG 0.955 0.990 0.955 0.955 0.955 149 7 7 149
IM 0.920 0.980 0.916 0.878 0.958 137 19 6 150
COAV 0.901 0.960 0.901 0.910 0.893 142 14 17 139
TAVeer 0.885 0.939 0.889 0.923 0.857 144 12 24 132
SiamBERT 0.865 0.920 0.873 0.828 0.923 144 12 30 126
$\bm{\mathcal{C}_{\mathrm{PJ}}}$ ${}^{\,(\bm{\dagger})}$ 0.740 0.805 0.771 0.690 0.872 136 20 61 95
LambdaG 0.909 0.961 0.911 0.929 0.893 158 12 19 151
IM 0.885 0.939 0.878 0.824 0.940 140 30 9 161
COAV 0.832 0.910 0.840 0.882 0.802 150 20 37 133
TAVeer 0.829 0.891 0.822 0.788 0.859 134 36 22 148
SiamBERT 0.629 0.655 0.652 0.615 0.694 118 52 74 96
$\bm{\mathcal{C}_{\mathrm{Apric}}}$ ${}^{\,(\bm{\dagger})}$ 0.800 0.871 0.814 0.760 0.876 149 21 47 123
Method Acc. AUC F1 Prec. Rec. TP FN FP TN
LambdaG 0.760 0.822 0.758 0.750 0.766 180 60 55 185
IM 0.694 0.761 0.610 0.479 0.839 115 125 22 218
COAV 0.685 0.748 0.684 0.683 0.686 164 76 75 165
TAVeer 0.690 0.729 0.655 0.588 0.738 141 99 50 190
SiamBERT 0.529 0.551 0.555 0.526 0.588 141 99 127 113
$\bm{\mathcal{C}_{\mathrm{Trip}}}$ ${}^{\,(\bm{\dagger})}$ 0.581 0.586 0.623 0.567 0.692 166 74 127 113
LambdaG 0.765 0.839 0.769 0.783 0.755 188 52 61 179
IM 0.650 0.735 0.613 0.554 0.686 133 107 61 179
COAV 0.638 0.728 0.640 0.646 0.635 155 85 89 151
TAVeer 0.671 0.749 0.686 0.721 0.655 173 67 91 149
SiamBERT 0.571 0.594 0.611 0.559 0.675 162 78 128 112
$\bm{\mathcal{C}_{\mathrm{Yelp}}}$ ${}^{\,(\bm{\dagger})}$ 0.658 0.727 0.703 0.622 0.808 194 46 118 122
LambdaG 0.796 0.874 0.791 0.772 0.811 618 182 144 656
IM 0.778 0.858 0.783 0.798 0.768 638 162 193 607
COAV 0.730 0.802 0.730 0.729 0.731 583 217 215 585
TAVeer 0.691 0.752 0.668 0.621 0.722 497 303 191 609
SiamBERT 0.621 0.672 0.652 0.602 0.711 569 231 376 424
$\bm{\mathcal{C}_{\mathrm{IMDB}}}$ ${}^{\,(\bm{\dagger})}$ 0.739 0.817 0.755 0.712 0.804 643 157 260 540
LambdaG 0.891 0.954 0.891 0.892 0.89 803 97 99 801
IM 0.827 0.908 0.812 0.750 0.886 675 225 87 813
COAV 0.781 0.867 0.789 0.820 0.761 738 162 232 668
TAVeer 0.781 0.867 0.778 0.767 0.789 690 210 185 715
SiamBERT 0.749 0.843 0.770 0.711 0.840 756 144 308 592
$\bm{\mathcal{C}_{\mathrm{Blog}}}$ ${}^{\,(\bm{\dagger})}$ 0.701 0.776 0.719 0.679 0.763 687 213 325 575
LambdaG 0.921 0.977 0.921 0.922 0.919 1107 93 97 1103
IM 0.856 0.932 0.855 0.851 0.859 1021 179 167 1033
COAV 0.801 0.884 0.805 0.820 0.790 984 216 261 939
TAVeer 0.842 0.915 0.839 0.822 0.857 986 214 164 1036
SiamBERT 0.856 0.937 0.864 0.821 0.912 1094 106 239 961
$\bm{\mathcal{C}_{\mathrm{Amazon}}}$ ${}^{\,(\bm{\dagger})}$ 0.864 0.941 0.871 0.830 0.916 1099 101 225 975
LambdaG 0.705 0.777 0.689 0.652 0.729 868 463 322 1009
IM 0.676 0.739 0.616 0.520 0.756 692 639 223 1108
COAV 0.633 0.701 0.638 0.648 0.629 863 468 510 821
TAVeer 0.635 0.682 0.614 0.581 0.651 773 558 414 917
SiamBERT 0.603 0.647 0.631 0.590 0.677 901 430 626 705
$\bm{\mathcal{C}_{\mathrm{AllNews}}}$ ${}^{\,(\bm{\dagger})}$ 0.807 0.886 0.809 0.799 0.820 1092 239 275 1056
Table 2 instead shows the values of $C_{\text{llr}}$ and $C_{\text{llr}}^{\text{min}}$ for both $\lambda_{G}$ and $\Lambda_{G}$, since both are log-likelihood ratios. A significant discrepancy is observed between the $C_{\text{llr}}$ and the $C_{\text{llr}}^{\text{min}}$ values for $\lambda_{G}$, indicating that $\lambda_{G}$ is uncalibrated. This means that although higher values of $\lambda_{G}$ do correctly correspond to Y-cases, the scale of variation does not reflect the expectations of a perfectly calibrated system, where $\lambda_{G}=0$ means an inconclusive result, a positive value suggests a Y-case, and a negative value suggests an N-case. When $\lambda_{G}$ is turned into $\Lambda_{G}$ by fitting a logistic regression on training data, however, the lower values of $C_{\text{llr}}\mbox{}(\Lambda_{G})$ are evidence that $\Lambda_{G}$ is well calibrated.
All results in Table 1 and Table 2 were obtained using LambdaG with parameters $r=100$ and $N=10$. Figure 1 shows how changes in these hyper-parameters affects Accuracy, revealing that the method is robust in the choice of these hyper-parameters. It is clear from the graph that $r=1$, meaning that only one reference Grammar Model $G_{\text{ref}}$ is extracted from the reference corpus, is not enough to achieve optimal results. However, in most corpora $r=100$ is not needed as performance stabilize at $r=30$.
Regarding the other hyperparameter $N$, the order of the $n$-gram model, the value $N=10$ was chosen to ensure that grammatical dependencies are captured even for long sentences and Figure 1 does suggest that this assumption is correct. For example, for a corpus such as $\mathcal{C}_{\mathrm{PJ}}$ consisting of chat messages or $\mathcal{C}_{\mathrm{Enron}}$ consisting of emails, $N=2$ is enough. In contrast, for the academic papers in $\mathcal{C}_{\mathrm{ACL}}$ and the news in $\mathcal{C}_{\mathrm{AllNews}}$ a larger order of the model clearly has an advantage. In all corpora, however, there is no gain in increasing $N$ to 20.
Figure 1: Accuracy results for different numbers of repetitions $r$ and orders of the $n$-gram model $N$ for each corpus.
Figure 2: The loss in Accuracy (top) and $C_{\text{llr}}$ (bottom) results for cross-corpus comparison, i.e. evaluating on Base Corpus while using reference texts $\mathbb{D}_{\text{ref}}$ from Reference Corpus. Diagonal bold values denote the original Accuracy and $C_{\text{llr}}$, respectively. Darker shades denote a greater loss in performance.
Another important robustness test that we performed is the selection of the reference data. Reliance on external data is a potential limitation for any binary extrinsic AV methods, like IM. As explained in Section 1, IM, suffers a substantial loss in performance if the impostors are selected from another genre. LambdaG also has the crucial requirement that a reference corpus $\mathbb{D}_{\text{ref}}$ is collected and it is presumably important that $\mathbb{D}_{\text{ref}}$ is as compatible as possible, especially in genre but generally speaking in as many parameters as possible, with $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{D}_{\mathcal{U}}$. To test the robustness of LambdaG to variations in $\mathbb{D}_{\text{ref}}$ type we ran 144 tests by pairing each of the twelve corpora to every other corpora. For example, we performed the analysis for $\mathcal{C}_{\mathrm{Enron}}$ but we used as $\mathbb{D}_{\text{ref}}$ $\mathcal{C}_{\mathrm{ACL}}$. We then calculated Accuracy and $C_{\text{llr}}$ for all combinations and we produced a plot for each metric displaying the resulting drop in performances when using a different $\mathbb{D}_{\text{ref}}$ (Figure 2).
Corpus $C_{\text{llr}}\mbox{}(\lambda_{G})$ $C_{\text{llr}}\mbox{}(\Lambda_{G})$ $C_{\text{llr}}^{\text{min}}$
$\mathcal{C}_{\mathrm{Enron}}$ 3.630 0.432 0.315
$\mathcal{C}_{\mathrm{Wiki}}$ 2.026 0.457 0.405
$\mathcal{C}_{\mathrm{Stack}}$ 5.716 0.499 0.411
$\mathcal{C}_{\mathrm{ACL}}$ 18.038 0.885 0.636
$\mathcal{C}_{\mathrm{PJ}}$ 2.000 0.254 0.143
$\mathcal{C}_{\mathrm{Apric}}$ 2.811 0.397 0.337
$\mathcal{C}_{\mathrm{Trip}}$ 2.666 0.752 0.715
$\mathcal{C}_{\mathrm{Yelp}}$ 1.872 0.726 0.691
$\mathcal{C}_{\mathrm{IMDB}}$ 2.714 0.648 0.626
$\mathcal{C}_{\mathrm{Blog}}$ 2.969 0.424 0.403
$\mathcal{C}_{\mathrm{Amazon}}$ 1.604 0.302 0.286
$\mathcal{C}_{\mathrm{AllNews}}$ 9.920 0.812 0.797
Table 2: Values of $C_{\text{llr}}$ and $C_{\text{llr}}^{\text{min}}$ for $\lambda_{G}$ and for $\Lambda_{G}$ generated by calibrating $\lambda_{G}$ with the logistic regression calibration method. $C_{\text{llr}}^{\text{min}}$ is identical for both $\lambda_{G}$ and $\Lambda_{G}$.
The diagonal of Figure 2 (top) shows the same results reported in Table 1, which function as the benchmark of performance for each row of the plot, corresponding to the corpus being tested. Overall the results suggest that LambdaG is extremely robust in many cases of genre variation, including some unrealistic genre combinations. This result is unexpected given how IM is instead very much dependent on the set of impostors [41]. For example, cross-corpus comparisons such as using $\mathcal{C}_{\mathrm{Yelp}}$ as reference for problems in $\mathcal{C}_{\mathrm{Amazon}}$ lead to very small losses in Accuracy and this is clearly explainable by the fact that both corpora contain reviews. However, certain results are more surprising, such as that using the academic papers from $\mathcal{C}_{\mathrm{ACL}}$ as reference data for problems in a corpus of chat logs such as $\mathcal{C}_{\mathrm{PJ}}$ leads to a performance loss of only 0.27 Accuracy, which is substantial but much less than expected given the disparity. This is a comparison that a human analyst would be unlikely to perform and therefore it is a real test of the lower limits in terms of robustness to reference corpora.
Looking at Figure 2 (top) row-wise, the most notable pattern is that the two corpora that suffer the most performance loss overall are $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$. For $\mathcal{C}_{\mathrm{PJ}}$ the explanation is likely to be that it is the only corpus with sentences mostly corresponding to chat messages, and therefore any match of this corpus with other corpora results in at least this incompatibility. The chat logs in $\mathcal{C}_{\mathrm{PJ}}$ are also the texts with the lowest level of formality and highest degree of linguistic freedom among all corpora. For $\mathcal{C}_{\mathrm{Stack}}$ this result could be due to its hybrid genre character (forum posts on academic topics) or it could simply be due the use of several tags to mask mathematical expressions or computer code. When looking at Figure 2 column-wise, instead, the most notable pattern is that using $\mathcal{C}_{\mathrm{ACL}}$ or $\mathcal{C}_{\mathrm{AllNews}}$ as reference corpora tends to affect performance more drastically for the majority of corpora, which again could be explained by their relatively higher degree of formality compared to the other corpora. All in all this analysis suggests that, for $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$ the reference corpus has a moderate impact on performance while for all other corpora it is only the use of $\mathcal{C}_{\mathrm{ACL}}$ or $\mathcal{C}_{\mathrm{AllNews}}$ as reference where this modest impact can be observed. For all other cross-corpus comparisons for most corpora the effect of the reference corpus is tolerable.
We also present the results for the $C_{\text{llr}}$ metric in Figure 2 (bottom). As explained in Section 5.4, a $C_{\text{llr}}$ value close to or equal to one means that the system is not providing any useful information for the case, while a $C_{\text{llr}}$ value higher than 1 means that results tend to be misleading (e.g. negative $\Lambda_{G}$ for a Y-problem). None of the cross-genre comparisons produce $C_{\text{llr}}$ higher than one, meaning that in all cases the values of $\Lambda_{G}$ are not misleading the analyst but can be, at worst, inconclusive with $\Lambda_{G}\approx 0$. These results support the principle that an analyst should always compile a reference data set that is as close as possible to the case data [52]. However, the results do also suggest that LambdaG is stable to small variations in genre.
The final test we carried out is a comparison of LambdaG as implemented in Algorithm 1 vs. LambdaG without applying the POSNoise function. The rationale for this test is to understand whether more authorial information can be extracted and exploited using this method beyond just grammatical structure. The results in Table LABEL:tab:EvalResultsPosnoise show that, regarding Accuracy and AUC, in all cases the proposed version of Algorithm 1 with the POSNoise pre-processing is better performing than a version of the algorithm without this pre-processing. The difference in Accuracy is not large but the method is clearly superior when the focus is entirely on grammar to the extent that adding information about meaning instead leads to adding noise in the authorial signal.
As explained in Section 5.3, the advantage of LambdaG is that it can be better interpreted than other methods. As a first step, the analyst can produce a ranking of all sentences in $\mathcal{D}_{\mathcal{U}}$ and then, additionally, each token can be color-coded according to its value of $\lambda_{G}$, similarly to COAV. Some of these visualizations explaining the features discovered are shown in Section 6.
$\Delta$Acc. $\Delta$AUC $\Delta$F1 $\Delta$Prec. $\Delta$Rec. $\Delta$TP $\Delta$FN $\Delta$FP $\Delta$TN
$\mathcal{C}_{\mathrm{Enron}}$
0.031 0.039 0.039 -0.017 0.083 4 -4 1 -1
$\mathcal{C}_{\mathrm{Wiki}}$
0.026 0.003 0.029 0.022 0.036 4 -4 -2 2
$\mathcal{C}_{\mathrm{Stack}}$
0.005 0.001 -0.002 0.047 -0.044 -5 5 -6 6
$\mathcal{C}_{\mathrm{ACL}}$
0.057 0.050 0.031 0.106 -0.050 -7 7 -23 23
$\mathcal{C}_{\mathrm{PJ}}$
0.009 0.005 0.010 0.001 0.019 3 -3 0 0
$\mathcal{C}_{\mathrm{Apric}}$
0.077 0.042 0.084 0.038 0.129 22 -22 -4 4
$\mathcal{C}_{\mathrm{Trip}}$
0.037 0.049 0.048 0.022 0.071 17 -17 -1 1
$\mathcal{C}_{\mathrm{Yelp}}$
0.057 0.046 0.063 0.043 0.083 20 -20 -7 7
$\mathcal{C}_{\mathrm{IMDB}}$
0.080 0.101 0.091 0.070 0.108 87 -87 -42 42
$\mathcal{C}_{\mathrm{Blog}}$
0.063 0.038 0.066 0.048 0.083 75 -75 -38 38
$\mathcal{C}_{\mathrm{Amazon}}$
0.046 0.030 0.046 0.041 0.050 60 -60 -49 49
$\mathcal{C}_{\mathrm{AllNews}}$
0.043 0.049 0.059 0.034 0.076 101 -101 -15 15
4 Discussion
Our proposed LambdaG approach matches many desiderata of an AV method:
  1. 1.
    It works on both short and long documents;
  2. 2.
    It is not affected or minimally affected by topic or content;
  3. 3.
    It is stable to small genre variations in the reference corpus;
  4. 4.
    It is more interpretable by an analyst;
  5. 5.
    It is not computationally intensive and does not have a high run time;
  6. 6.
[…]
  7. 7.
    It has a theoretical foundation based on Cognitive Linguistic theories of language processing.
Section 3 showed evidence of the first six of these points as well as evidence that LambdaG outperforms other AV state-of-the-art methods. We believe that a possible reason to explain these results, especially considered the simplicity of the method compared to approaches such as SiamBERT, is precisely that LambdaG has a number of properties that make it consistent with knowledge about language processing gained from studies in Cognitive Science and Cognitive Linguistics.
Since the pioneering study of Mosteller and Wallace [79], it is well known that the frequency of functional items distinguish authors and this conclusion is also evident in the case of LambdaG. The reason why this is the case, however, has always been seen as a mystery in the field of authorship analysis and stylometry [56]. A connection between the unreasonable effectiveness of these functional items and Cognitive Linguistics has been proposed by Nini [82] and we believe that the success of LambdaG is additional evidence for these explanations.
Cognitive Linguistics models language as a complex adaptive system [9], where regularities emerge from the complexity of language used in interactions, like a phenomenon of the third kind [33]. This is partially evidenced by decades of linguistics studies on corpora which have revealed that the traditional conception of a grammar as an abstract set of phrase structure rules is untenable and that, instead, a better model of grammar is a lexicogrammar continuum that goes from fixed lexicalized expressions (e. g., “I don’t know”) to flexible templates (e. g., “the” X-er “the” Y-er) up to completely abstract structures (e. g., Doer Action Object).
This continuum roughly corresponds to the continuum between declarative memory, which is semantic/encyclopedic memory for facts and which tends to be conscious, and procedural memory, which is instead memory for the probabilistic prediction of patterns of sequences and that deals with subconscious knowledge of habits and skills [108]. The roles that these two memories play in language is clearly evidenced in cases of aphasia. When an impairment affects circuits of declarative memory, this gives rise to fluent aphasia, resulting in speech that is mostly constituted by function words like articles, prepositions, pronouns, and morphological forms but with lack of content words that contain meaning, leading to fluent but nonsensical speech comparable to POSNoise pre-processed data such as the one in Table 5. Instead, when the impairment is on procedural memory, then speech is only consisting of content words that are not connected through grammar [109]. These two memory systems work together and, typically, once a habit is formed through repetition then procedural memory takes over, leading to fluency of expression. This transition explains why languages tend to be heavily formulaic. For example, in English one can say big mistake but not large mistake, even though the formal grammar would seemingly have no problems with the latter expression, which is an adjective modifying a noun. What explains this reliance on formulaic expressions is the cognitive tendency to recur to memory as much as possible rather than creating something new every time we use language [108].
The way these units of procedural memory are formed is through the process of chunking: when two items that often go together are found together, then these two items can be treated as one, thus saving space in working memory [39, 21, 28]. In all effects, chunking is a process of information compression and, if the units of grammar are chunks, then the logical inference is that grammar itself can be thought of as a compressed version of a language or, equivalently, as a probability model. A probability model of only functional items is therefore quite like the Grammar Models used in this study.
Cognitive Linguistics also explains why each individual has a different and possibly unique grammar, a statement encapsulated by the Principle of Linguistic Individuality [82]. The status of unithood given by chunking is not discrete (e. g., something is or is not a unit) but a continuum of strength of unithood, where this strength is called entrenchment [32, 97]. Because entrenchment is dependent on exposure and usage, it is fundamentally personal and idiosyncratic. It is for this reason that Langacker hypothesized that “the set of units mastered by all members of a speech community might be a rather small proportion of the units constituting the linguistic ability of any given speaker” [71], a hypothesis that is finding increasing empirical support [82]. For example, for most if not all speakers of English the entrenchment for a sequence such as “of the” is maximal, meaning that they can produce it effortlessly and unconsciously as a single unit. In contrast, a sequence such as “the of” would not be a unit for speakers of English because it is virtually never encountered or used. If a speaker of English encounters this sequence, they would decompose it as a structure made up of two units, “the” and “of”. In between these two extreme examples there is a lot of freedom as to what counts as a unit for a person or another, depending on what each individual has automatized in their procedural memory. For example, an individual $\mathcal{A}_{1}$ might have a very entrenched sequence such as “it is known that” when writing academic articles while an individual $\mathcal{A}_{2}$ might have “it is well established that”. These sequences will then have open slots ready to be filled in with the appropriate meaning. This kind of phenomenon theorized by Langacker can be seen at play in this method, especially when looking at the examples of idiosyncratic structures in Section 6. As briefly mentioned in Section 5.3, therefore, the likelihood of a token given a context assigned by an author $\mathcal{A}$’s Grammar Model, $P(t|c;G_{\mathcal{A}})$, can be seen as a way to measure the level of entrenchment of token $t$ in context $c$ for $\mathcal{A}$ The higher the entrenchment, the more likely it is that $\mathcal{A}$ is using procedural memory and that it is therefore performing an unconscious routine that constitutes a unit for this individual.
There have been other attempts at modeling the grammar of an author in the past for AV or AA purposes [24, 7, 37] but these methods relied too much on Part of Speech tags or on syntactic trees which are often not compatible with predictions made in Cognitive Science and Cognitive Linguistics. As explained above, the picture of grammar painted by Cognitive Linguistics is much more lexically-based and word-idiosyncratic. For example, the preposition “of” in English behaves very oddly compared to other prepositions [100]. Giving it the same Part of Speech label as to other prepositions is therefore going to obscure a lot of important grammatical information. Instead, a topic-masking method such as POSNoise is much more consistent with the predictions of Cognitive Linguistics because it leaves the function items unaffected while only obscuring those items are more likely to change from context to context depending on the communicative needs.
To sum up, the model of an individual’s grammar that emerges from Cognitive Science and Cognitive Linguistics is one that is very much compatible with the Grammar Models used for LambdaG. Although language modeling is seen in computational linguistics as a tool and not necessarily a realistic representation of language, evidence from Cognitive Linguistics suggests that, in reality, a language model, while not being realistic as a model of linguistic production, is instead a plausible representation of how knowledge of grammar is stored in the mind. The mechanisms of such a sequential processing can be efficiently represented through the short-term correlations captured by $n$-gram models as done in this work.
A deeper connection between LambdaG and approaches to AV based on language modeling or, as for COAV, on compression algorithms can also be drawn, thus offering a plausible explanation for their success. Through the lenses of Information Theory, the average $P(s;G)$ for each $s\in\mathcal{D}$ can be also seen as an estimate of the cross-entropy rate of the Grammar Model $G$ against the probability distribution of grammatical tokens that generates $\mathcal{D}$. In language modeling literature, the cross-entropy is often exponentiated to calculate the perplexity of a document given a model. Because for LambdaG the two alternative Grammar Models are compared in relation to the same $\mathcal{D}$, for the purposes of this AV application, $\lambda_{G}$ is equivalent to comparing the perplexities of the candidate author’s model vs. the reference model given the same $\mathcal{D}$ and assign the document to the candidate author if this perplexity is lower. The cross-entropy rate can also be interpreted as the average length in bits of a binary encoding of grammatical tokens for a data-compression scheme that is optimized for the probability distribution $P(\cdot;G)$. For this reason, a language model can be seen as a text-compression scheme. This connection thus links compression and language model AV methods like LambdaG and COAV to the Cognitive Linguistics conception of an individual’s grammar as a compressed representation of real language usage encountered by said individual.
In addition to the successful results of LambdaG, we note that the qualitative explorations reported in Section 6 also constitute evidence that LambdaG captures linguistic behavior compatible with these Cognitive Linguistic explanations. Although it is impossible to know for certain the effective mental status of the sequences found in the examples in Section 6, we propose that it is plausible to believe that the authors of these texts do not know that they use the sequences of tokens highlighted in red or that these sequences identify them in the general population. These sequences, which are often seemingly unremarkable (e. g.,enough to be ADVERB VERB” or “so PRONOUN cant”), can be extremely idiosyncratic and uniquely used only by the candidate author. Given the fact that the problems analyzed in this paper involve relatively short texts (the longest texts are just around 2000 tokens), we must conclude that these idiosyncratic units are easy to find and this conclusion therefore strongly supports Langacker’s proposal that most units in a language are not shared by many individuals.
We believe that additional evidence towards this assertion is given by the results of the cross-reference experiments. Controlling for genre compatibility is probably the hardest problem in AV and AA. LambdaG is very robust to this variation in many cases, much more than expected given the results of similar tests carried out on IM [41]. This is despite the incredible challenges in some of the comparisons, such as in the analysis of the $\mathcal{C}_{\mathrm{PJ}}$ chat logs using as reference the academic papers in $\mathcal{C}_{\mathrm{ACL}}$. The genre differences between these two corpora cannot be greater and therefore it is hard to believe that the method is not performing at chance level. A possible explanation for the success of LambdaG in these circumstances is that the reference corpus effectively does not play a major role. If individual grammars are so different from each other, then the role of the reference corpus could be to just weight down the those units that are well-entrenched for most individuals (e. g., of the) and that therefore carry almost no authorial information. If this explanation is correct, then this means that once these common units are taken into account, the rest is idiosyncratic enough to be identifying.
Finally, due to the fact that these idiosyncratic units are easy to find within as little as 1,000 to 2,000 tokens, we add that a consequences of the present findings and of the underlying Cognitive Linguistic theory is that the grammar of a person can be seen as a behavioral biometric [82]. Similarly to handwriting, gait, typing patterns, computer mouse usage, and most features of human voice, grammar is also a cognitive process relying on procedural memory that strengthens with habit and repetition. We propose that this is the explanation underlying most if not all AV and AA methods that, by measuring the frequency of functional items, are indirectly analyzing the idiosyncratic grammar of an individual. LambdaG tends to be more successful probably because it is measuring the same phenomenon more directly.
4.1 Limitations and future research
We argue that future research should therefore continue in this direction of extracting better representations of the grammar of an individual. The POSNoise algorithm used here is language dependent and suffers from the use of general part of speech categories that are not compatible with Cognitive Linguistics, which instead predicts the existence of categories that are more semantic in nature and at lower degrees of abstractness. New future methods should embrace the close relationship between Cognitive Linguistics and Information Theory to gain more insight into grammar, which could consequently lead to further breakthroughs in AV.
In addition, LambdaG is still affected by genre differences and performs less well in registers where there is less variation, such as academic prose. These two last limitations affect more or less all existing algorithms which are outperformed by LambdaG, which could mean that these limitations are intrinsic to AV and, indeed, there would be good linguistic reasons for why this would be the case. Despite these limitations, future methods based on a better understanding of grammar could still lead to more progress. In this respect, the reliance on a reference corpus could also be seen as a limitation of any binary AV method. The tendency for future AV methods should be to increase the power of unary methods that only need a test dataset.
Finally, another limitation of the present paper is that the research was carried out entirely on English. The theory underlying LambdaG generalizes to other languages and for this reason the prediction is that it would also outperform the same baseline methods in other languages. The way Grammar Models are trained in languages with a rich morphology is likely to be different, however. Different tokenizers might have to be employed as well as different algorithms such as POSNoise to focus the model on the grammatical aspects of the text. Another goal for the future would therefore be to find a language-independent variant of LambdaG.
4.2 Final Remarks
The advances in AV made in this paper were possible because of the collaboration between Linguistics and Natural Language Processing. In the 2022 edition of the PAN competition on AV, Stamatatos et al. [104] concluded that “all submissions, despite their increased level of sophistication in most of the cases, were outperformed by a naive baseline based on character $n$-grams and cosine similarity” thus noticing a situation of stall in the progress on AV. Knowledge from Linguistics could resolve these stalls, as demonstrated in this paper, by shifting the point of view rather than adding more computational sophistication or making algorithms more complex. The method and philosophy introduced in the present work offer significant advancements in integrating Linguistics theory and Computer Science methods towards a new paradigm in AA and AV where these two fields are better harmonized. The task of identifying a writer from the language used in a text is really an application of Linguistics and there are various reasons for computational methods to maintain a close relationship with Linguistics. Firstly, these methods directly or indirectly exploit linguistic patterns that require knowledge of Linguistics to be described and understood. Secondly, knowledge of Linguistics is needed to understand and potentially explain in a court of law the reasons why these patterns identify or exclude a certain author. We believe that the work in this paper has made some significant advances in AV while also proposing and stimulating a paradigm change to move on from the impasse described by Stamatatos [104] towards new AV methods that are not only well-performing but also scientifically meaningful.
5 Methods
In this section, we present the more technical details of our work, first with an introduction to the corpora we have compiled to evaluate LambdaG, then by providing more details about the POSNoise content-masking algorithm as well as the details of the Kneser-Ney smoothing algorithm, the likelihood ratio framework for forensic sciences and, finally, the Siamese neural network baseline method, SiamBERT, developed for this analysis.
5.1 Corpora
To evaluate LambdaG, we compiled a collection of 12 corpora covering a variety of challenges. These include, for instance, very short texts ($\approx 5$ sentences), texts with a lack of coherence, texts with excessive use of slang, texts (of the same author) written at different times, texts with opposing topics, texts with distorted writing style caused by post-editing, texts with truncated sentences as well as non-standard spellings and missing spaces. In total, the corpora comprise 17,306 verification cases, were each corpus $\mathcal{C}=\{c_{1},c_{2},\ldots\}$ is split into author-disjunct training and test partitions based on a 40/60 ratio, resulting in 6,402 training and 10,904 test verification cases. Each $c\in\mathcal{C}$ denotes a verification case $(\mathcal{D}_{\mathcal{U}},\mathbb{D}_{\mathcal{A}})$, where $\mathcal{D}_{\mathcal{U}}$ is an unknown document and $\mathbb{D}_{\mathcal{A}}$ is a set of example documents of the known author $\mathcal{A}$. To counteract population homogeneity, a known bias in authorship verification described by Bevendorff [11], we ensured that for each author $\mathcal{A}$ there are exactly two (one same-authorship (Y) and one different-authorship (N)) verification cases. Furthermore, we constructed all corpora so that the number of Y- and N-cases is equal, in other words, all training and test corpora are (strictly) balanced. In what follows, we describe all corpora in detail, underline their respective challenges and explain the pre-processing [^2] The pre-processing of the documents was carried out using TextUnitLib https://github.com/Halvani/TextUnitLib steps we have performed. A compact overview of the 12 corpora, including their partitioning into training and test corpora, as well as further information can be found in Table 4.
Corpus Genre Topic Partition $|\mathcal{C}|$ avg$(\mathbb{D}_{\mathcal{A}})$ avg$(\mathcal{D}_{\mathcal{U}})$
$\mathcal{C}_{\mathrm{Enron}}$ E-mails Related $\mathcal{C_{\,\textrm{train}}}$ 64 783 (3866) 906 (4432)
$\mathcal{C_{\,\textrm{test}}}$ 96 763 (3808) 882 (4391)
$\mathcal{C}_{\mathrm{Wiki}}$ Wikipedia talk pages Related $\mathcal{C_{\,\textrm{train}}}$ 150 467 (1266) 614 (1654)
$\mathcal{C_{\,\textrm{test}}}$ 226 492 (1328) 626 (1699)
$\mathcal{C}_{\mathrm{Stack}}$ Q & A posts Cross $\mathcal{C_{\,\textrm{train}}}$ 150 2204 (11247) 1972 (9956)
$\mathcal{C_{\,\textrm{test}}}$ 228 2306 (11803) 2114 (10700)
$\mathcal{C}_{\mathrm{ACL}}$ Scientific papers Related $\mathcal{C_{\,\textrm{train}}}$ 186 1707 (9497) 2553 (14196)
$\mathcal{C_{\,\textrm{test}}}$ 280 1479 (8250) 2492 (13913)
$\mathcal{C}_{\mathrm{PJ}}$ Chat logs Related $\mathcal{C_{\,\textrm{train}}}$ 208 594 (2329) 752 (2946)
$\mathcal{C_{\,\textrm{test}}}$ 312 592 (2324) 754 (2967)
$\mathcal{C}_{\mathrm{Apric}}$ Forum posts Related $\mathcal{C_{\,\textrm{train}}}$ 228 796 (3900) 825 (4023)
$\mathcal{C_{\,\textrm{test}}}$ 340 799 (3921) 817 (4020)
$\mathcal{C}_{\mathrm{Trip}}$ Travel reviews Related $\mathcal{C_{\,\textrm{train}}}$ 120 658 (3169) 249 (1186)
$\mathcal{C_{\,\textrm{test}}}$ 480 698 (3344) 252 (1202)
$\mathcal{C}_{\mathrm{Yelp}}$ Hotel/restaurant reviews Related $\mathcal{C_{\,\textrm{train}}}$ 320 139 (637) 167 (767)
$\mathcal{C_{\,\textrm{test}}}$ 480 139 (640) 167 (768)
$\mathcal{C}_{\mathrm{IMDB}}$ Movie reviews Mixed $\mathcal{C_{\,\textrm{train}}}$ 400 658 (3151) 374 (1796)
$\mathcal{C_{\,\textrm{test}}}$ 1600 690 (3291) 366 (1753)
$\mathcal{C}_{\mathrm{Blog}}$ Blogs posts Mixed $\mathcal{C_{\,\textrm{train}}}$ 1200 742 (3416) 881 (4086)
$\mathcal{C_{\,\textrm{test}}}$ 1800 740 (3385) 882 (4061)
$\mathcal{C}_{\mathrm{Amazon}}$ Product reviews Mixed $\mathcal{C_{\,\textrm{train}}}$ 1600 861 (4010) 869 (4041)
$\mathcal{C_{\,\textrm{test}}}$ 2400 860 (4009) 868 (4041)
$\mathcal{C}_{\mathrm{AllNews}}$ News articles Mixed $\mathcal{C_{\,\textrm{train}}}$ 1776 675 (3629) 941 (5064)
$\mathcal{C_{\,\textrm{test}}}$ 2662 680 (3665) 942 (5071)
Table 4: All used corpora and their key statistics. Notation: $|\mathcal{C}|$ denotes the number of verification cases in each corpus $\mathcal{C}$, while avg$(\mathbb{D}_{\mathcal{A}})$ and avg$(\mathcal{D}_{\mathcal{U}})$ denotes the average token count (and character count in parentheses) of all $\mathcal{D}_{\mathcal{A}}\in\mathbb{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$, respectively.
5.1.1 $\mathcal{C}_{\mathrm{Enron}}$
The $\mathcal{C}_{\mathrm{Enron}}$ corpus was derived from the well-known Enron Email Dataset333available at https://www.cs.cmu.edu/~enron, which has been used in many AV and AA studies, including [17, 18, 19, 27, 31, 72, 83, 112, 4], as well as in many other research areas. The Enron email dataset has gained considerable popularity in the authorship analysis community, likely due to the fact that it is (to the best of our knowledge) the only publicly available corpus of real-world emails with a guaranteed ground truth. Similar to previous AV studies, we also reduced the initial set to a smaller number of authors (in our case from 150 to 80), as only a few suitable emails were available for some of the authors. In total, $\mathcal{C}_{\mathrm{Enron}}$ consists of 362 documents, with each document representing a concatenation of several emails from the corresponding author in order to obtain a sufficient length. Since the emails in the Enron Email Dataset contain a wide range of noise, extensive preprocessing was necessary. To this end, we decided to manually preprocess all considered texts and instead create individual replacement rules for various types of noise. First, we removed URLs, email headers, greeting/closing formulas, email signatures, (phone) numbers, quotes as well as various non-letter repetitions. Next, we normalized UTF-8 symbols and, in a final step, replaced multiple consecutive spaces, line breaks and tabs with a single space. A similar preprocessing procedure was also described by Brocardo et al. [17, 18].
5.1.2 $\mathcal{C}_{\mathrm{Wiki}}$
The $\mathcal{C}_{\mathrm{Wiki}}$ corpus consists of 752 excerpts from 288 Wikipedia talk page editors taken from the Wikipedia sockpuppets dataset, which was released by Solorio et al. [101]. The original dataset contains two partitions comprising sockpuppets and non-sockpuppets cases, where for $\mathcal{C}_{\mathrm{Wiki}}$ we only considered the latter subset. In addition, we did not use the full range of authors within the non-sockpuppets cases, as there were not sufficiently long texts available for each of the authors. Regarding the selected texts, we removed Wiki markup, timestamps, URLs and other types of noise. Furthermore, we removed sentences containing many numbers, proper names and near-duplicate string fragments as well as truncated sentences.
5.1.3 $\mathcal{C}_{\mathrm{Stack}}$
The $\mathcal{C}_{\mathrm{Stack}}$ corpus consists of 567 posts from 189 users crawled from the question-and-answer (Q & A) network Stack Exchange [^4] https://stackexchange.com in 2019. The network comprises 173 Q & A communities, with each community focusing on a specific topic (e. g., linguistics, computer science, philosophy, etc.). For the compilation of $\mathcal{C}_{\mathrm{Stack}}$, which represents a cross-topic corpus, we collected questions and answers from users who were simultaneously active in the two thematically different communities Cross Validated ($\mathcal{S}_{1}$) and Academia ($\mathcal{S}_{2}$). The verification cases were constructed as follows. Each verification case consists of exactly two documents, a known and an unknown document $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$. Within the Y-cases, $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ were taken from $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ respectively, while within the N-cases $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ were taken from the same community $\mathcal{S}_{1}$. Due to the cross-topic nature of this corpus, AV methods that utilize topic-based features are more likely to provide inverse predictions (i. e., N-cases are more likely to be classified as Y and vice versa) while AV methods that use topic-agnostic features behave more robustly. Another challenge of $\mathcal{C}_{\mathrm{Stack}}$ is that the corpus is very unbalanced with respect to the length of the documents, with lengths between 3014 and 20476 characters. To avoid stylistic distortions from the outset, we only used the original posts of the respective authors for all texts in $\mathcal{C}_{\mathrm{Stack}}$, while subsequent edits by third parties (for the purpose of spelling or stylistic improvement) were excluded [^5] Stack Exchange provides metadata to show exactly which parts of the text have been edited, and also allows the original unedited post to be retrieved.. In addition, we replaced equations/formulas, numerical values (e. g., percentages and currency values), dates, references and URLs with appropriate placeholders. The reason why we did not omit the affected sentences and instead opted for this approach was that we wanted to keep as much text content per author as possible to ensure a sufficient number of authors.
5.1.4 $\mathcal{C}_{\mathrm{ACL}}$
The $\mathcal{C}_{\mathrm{ACL}}$ corpus comprises 466 excerpts from scientific papers by a total of 233 researchers, taken from the computational linguistics archive ACL Anthology [^6] https://aclanthology.org/. The corpus was constructed in such a way that there are exactly two papers for each author [^7] Based on the metadata of the papers (more precisely, the author fields within the corresponding BibTex entries), we ensured that each paper is from exactly one author., dating from different time periods. The average time span between the two documents of each author is approximately 15.6 years, while the minimum and maximum time spans are 8 and 31 years, respectively. In addition to the temporal aspect of the documents in $\mathcal{C}_{\mathrm{ACL}}$, another challenge of this corpus is the formal language, in which the use of stylistic devices (e. g., repetitions, metaphors or rhetorical questions) is more restricted in contrast to other text types such as chat logs, forum posts or product reviews. For the original papers, we have tried to limit the content of each text as much as possible to those sections that consist mainly of natural language text (e. g., abstract, introduction, discussion, conclusion or future work). To ensure that the extracted fragments are consistent with the remarks of Bevendorff et al. [11], we decided to preprocess each paper excerpt in $\mathcal{C}_{\mathrm{ACL}}$ (semi-)manually. Among other things, we removed tables, quotations and sentences that contained a large amount of non-linguistic content such as mathematical constructs as well as specific names of researchers, systems and algorithms. We also replaced (inline) mathematical expressions, references (e. g., “[12]”, “[SHL13]”), endophoric markers, numeric values and URLs with appropriate placeholders, as we did with the texts in $\mathcal{C}_{\mathrm{Stack}}$.
5.1.5 $\mathcal{C}_{\mathrm{PJ}}$
The $\mathcal{C}_{\mathrm{PJ}}$ corpus comprises a total of 738 chat logs from 260 sex offenders collected from the Perverted-Justice [^8] http://www.perverted-justice.com portal. The chat logs were obtained from older instant messaging clients (e. g., MSN, AOL and Yahoo), whereby for each conversation we ensured that only the offender’s chat lines were extracted based on the annotations provided by the portal. To maximize variability in terms of conversation content, we selected chat lines from different messaging clients as well as from different time spans (where possible). A major challenge of $\mathcal{C}_{\mathrm{PJ}}$ is that it contains texts characterized by excessive use of non-standard English and also contains many chat-specific abbreviations and other forms of noise. This in turn makes it difficult to extract syntax-based features such as POS tags from the texts, at least when using NLP models trained on texts written in standard English (e. g., news articles or scientific documents). Another challenge of $\mathcal{C}_{\mathrm{PJ}}$ is that the text fragments it contains are very short, which in turn limits stylistic variability. In contrast to $\mathcal{C}_{\mathrm{Enron}}$, we have automatically pre-processed the texts in this corpus. Among other things, we sorted out chat lines with less than 5 tokens and 10 characters or those containing usernames, timestamps, URLs and words with many repeating characters (e. g., “mmmhhhmm”, “heeheehee” and “:-*:-*:-*:-*:-*:-*:-*:-*”). To further increase the variability in the content of each chat log $X$, we made sure that all contained lines $(x_{1},x_{2},\ldots,x_{m})$ are different from each other. For this purpose, we applied a string similarity function to each pair $(x_{i},x_{j})$ with $i\neq j$. As a metric for string similarity, we chose the Levenshtein distance [^9] Regarding the fuzzy string matching, we used FuzzyWuzzy https://github.com/seatgeek/thefuzz based on a threshold of 0.25, which we considered appropriate for our purposes. All pairs of lines that exceeded this threshold were discarded with respect to the text fragments.
5.1.6 $\mathcal{C}_{\mathrm{Apric}}$
The $\mathcal{C}_{\mathrm{Apric}}$ corpus comprises 1,395 forum posts from 284 users crawled from the portal The Apricity - A European Cultural Community [^10] https://theapricity.com. To compile the $\mathcal{C}_{\mathrm{Apric}}$, we ensured that all posts within each verification case were collected from different sub-forums covering a variety of topics, including anthropology, genetics, race and society and ethno-cultural discussion. We then pre-processed each post and removed certain markup tags, URLs, forum signatures, quotations (as well as nested quotations that contained the original author’s content), and other types of noise, similar to what we did with the texts within the $\mathcal{C}_{\mathrm{Enron}}$ corpus. However, despite the pre-processing, some textual peculiarities remained in the texts, such as truncated sentences, spelling errors and missing spaces between adjacent words or between a punctuation mark and the following word. These can be seen as a challenge of $\mathcal{C}_{\mathrm{Apric}}$, especially with regard to AV methods that use features based on word lists.
5.1.7 $\mathcal{C}_{\mathrm{Trip}}$
The $\mathcal{C}_{\mathrm{Trip}}$ corpus comprises reviews from 300 bloggers extracted from the Webis TripAdvisor Corpus 2014 published by Trenkmann and Spiel [107]. As the original corpus was not intended for the purposes of AV, we reformatted the data accordingly. While inspecting the data, we found that most of the included reviews were already of reasonable quality, so we only carried out moderate pre-processing. Among other things, we adjusted spaces next to punctuation marks (e. g., “( $\rightarrow$(alight”)) and normalized many character repetitions to a maximum of three characters (e. g., “(!!!!$\rightarrow$!!!”). Besides, we excluded reviews that contained too many named entities and we discarded authors for which only one review was available.
5.1.8 $\mathcal{C}_{\mathrm{Yelp}}$
The $\mathcal{C}_{\mathrm{Yelp}}$ corpus comprises 2,400 reviews written by 400 so-called “Yelplers” extracted from the well-known Yelp dataset [^11] https://www.yelp.com/dataset. The corpus was developed with several challenges in mind. First, compared to all other corpora we have compiled, it contains the shortest texts, with each $\mathcal{D}_{\mathcal{U}}$ consisting of at most ten sentences. For this reason, the stylistic variability within these texts is correspondingly limited. Secondly, many reviews were written at different times (in several cases often about four years apart), so linguistic changes are to be expected in reviews written by the same person. Thirdly, most of the texts in the $\mathcal{C}_{\mathrm{Yelp}}$ are restaurant reviews, which focus heavily on similar topics such as food, atmosphere and taste experiences. To a lesser extent, however, other forms of reviews are also included in the corpus (e. g., about wedding or clothing stores), which are accompanied by linguistic and also stylistic diversity with respect to the authors. Fourth, some of the reviews contain a mixture of moderate and colloquial writing styles. Regarding the latter, we observed certain spelling variations in the texts (e.g. missing apostrophes within contractions), which may result in noticeable loss of features for AV methods that rely on word lists. For the review texts in $\mathcal{C}_{\mathrm{Yelp}}$, we performed the following pre-processing steps. Sentences containing capitalized words throughout were removed (e. g., “I WAITED 4 HOURS IN THE BEND AND COULDN’T GET IN!!!!!!!!”), as were sentences containing quotes and many repetitive characters (e. g., “wait for a heart attack!!!!! Incompetent staff!!!! will never come here again!!!!”). Moreover, we removed sentences that only consist of single words (e. g., “fraud. Fraud. Lies.”) as well as review-type summaries such as “Quality: 10/10; Service: 9/10; Ambience/Location: 10/10; Overall: 9.5/10”.
5.1.9 $\mathcal{C}_{\mathrm{IMDB}}$
[…]
5.1.10 $\mathcal{C}_{\mathrm{Blog}}$
The $\mathcal{C}_{\mathrm{Blog}}$ corpus comprises 6,000 documents written by 1,500 bloggers, taken from the well-known Blog Authorship Corpus released by Schler et al. [96]. The original dataset comprises 681,288 posts of 19,320 bloggers and was collected from blogger.com in 2004. However, a large proportion of the posts were not suitable for our purposes for various reasons. For example, many posts consisted of non-English content or contained too much noise (e. g., superfluous strings, duplicate phrases, quotes, etc.) that was difficult to preprocess without losing too much text for the respective authors. In addition, the amount of text was insufficient in itself for many authors. Therefore, we only considered a smaller fraction of the original corpus. To counteract various noise types, we first sampled sentences from the original posts and ensured that each sentence is solely restricted to (case-sensitive) English letters, spaces and standard punctuation marks. From the resulting texts we then removed sentences containing repeated characters (e. g., “?!??!”, “---” and “-.-;;”) and words (e. g., “FUN!! FUN!! FUN!!” and “kekeke”), as well as sentences with many capitalized words (e. g., “i am saying that WHOEVER HAS OPTUS PLEASE TELL ME”). Besides, we discarded sentences containing URLs, gibberish strings (e. g., “rak;db’dfth6yortyi95”, “aDr0” and “w00t”) as well as truncated sentences. Finally, we concatenated the preprocessed sentences into a single document, which in turn led to a disruption of the coherence of the resulting texts. Similar to $\mathcal{C}_{\mathrm{PJ}}$, the texts in $\mathcal{C}_{\mathrm{Blog}}$ also exhibit an excessive amount of non-standard English expressions such as “w8 4” (wait for), “b4” (before), “2day” (today), “sum1’s” (someone’s) and “jnr” (junior). This can be seen as another challenge of $\mathcal{C}_{\mathrm{Blog}}$, since the spelling of such words is not always consistently adhered to with respect to the authors.
5.1.11 $\mathcal{C}_{\mathrm{Amazon}}$
[…]
5.1.12 $\mathcal{C}_{\mathrm{AllNews}}$
The $\mathcal{C}_{\mathrm{AllNews}}$ corpus consists of 4,438 news articles written by 2,219 authors. The articles stem from the large well-known All the News 2.0 dataset [^12] The dataset can be accessed at https://components.one/datasets/all-the-news-2-news-articles-dataset, which consists of 2.7 million news articles and essays from 27 American news agencies and podcasting companies. From these, we limit ourselves to a single source, namely Reuters, which contains the largest number of news articles (840,094 in total). However, since not all articles were suitable for the creation of $\mathcal{C}_{\mathrm{AllNews}}$, it was necessary to preprocess the initial Reuters subset accordingly. First, we discarded all authors who wrote only a single article, as otherwise there is no suitable possibility to construct two verification cases for each author. Next, we sorted all articles of each author by their publication year and kept only four articles (the two oldest and the two newest). The main reasons for this were firstly to maximize the stylistic variation of each author and secondly to ensure sufficient length of the resulting documents, which in some cases were massively shortened by the pre-processing. Based on the four documents per author, the pre-processing was performed as follows. First, we removed all quotations (which are very common due to the nature of news articles) to mitigate the degree of stylistic distortion. Next, we removed headlines, sentences that consisted of more than 40 percent named entities and/or numeric tokens (e. g., dates, amounts of money, percentages, etc.), and article signatures that contained names of people who reported, wrote, and edited the articles. Finally, we normalized apostrophes (e. g., $\rightarrow$”) and replaced multiple consecutive spaces with a single space. The pre-processing procedure severely impaired the coherence of the resulting documents, as the natural order of many sentences was no longer given and some sentences also resulted in a truncated form. The impaired coherence can be seen as a challenge of this corpus, as AV methods that use implicit features (e. g., character/word $n$-grams) may capture “artificial features” that occur across sentence boundaries. Another challenge of $\mathcal{C}_{\mathrm{AllNews}}$ is that the documents often have been edited by other authors (e. g., the editor-in-chief) before online publication, so the original writing style of the articles may be distorted to some extent. Unlike the Stack Exchange network, All the News 2.0 does not provide metadata indicating which specific parts of the original article have been edited. Therefore, it was not possible for us to remove these afterwards.
5.2 Topic-Masking via POSNoise
Authorship verification constitutes a similarity detection problem in which the subject of the similarity determination is the language of the author rather than other document aspects such as the topic [42]. However, a large proportion of existing AV methods including [1, 17, 19, 22, 69, 70, 72, 80, 88] use features that are directly influenced by the topic and thus run the risk of biasing their predictions accordingly. To counteract the problem of biased predictions regarding LambdaG, which also employs topic-affected features, we apply topic masking to all documents in our corpora. Topic masking is an approach to pre-process texts in such a way that topic-dependent text units (e. g., words, phrases, etc.) are masked by appropriate placeholders, while grammatical text units (e. g., determiners, conjunctions, contractions, prepositions, pronouns, auxiliary verbs, quantifiers, transitional phrases, etc.) are retained within the texts. Of the few topic masking approaches published in the literature, we chose POSNoise presented by Halvani and Graner [42], which proved to outperform the previously published topic masking approach called TextDistortion [103] in 34 out of 42 cases, with an increase in accuracy of up to 10%.
Corpus Original / topic-masked sentences
$\mathcal{C}_{\mathrm{Wiki}}$ If they actually censor anything is another question.
If they © Ø anything is another #.
$\mathcal{C}_{\mathrm{Stack}}$ When I took the course, I ended up not being able to learn from the professor
When I took the #, I Ø up not being @ to Ø from the #
$\mathcal{C}_{\mathrm{PJ}}$ going to wash clothes ,call me,tim
going to Ø # ,call me,§
$\mathcal{C}_{\mathrm{Yelp}}$ Our meal started with a basket of freshly baked baguette and focaccia.
Our # started with a # of © Ø # and #.
$\mathcal{C}_{\mathrm{Blog}}$ Then, Anna, Peter and I all went to baker’s Square.
Then, §, § and I all went to #’s §.
$\mathcal{C}_{\mathrm{Amazon}}$ I got a new mascara that was impervious to makeup remover
I got a @ # that was @ to # #
Table 5: Comparison between the original texts and their resulting topic-masked representations generated by POSNoise [42, Algorithm: 1]. The semantics of the placeholders can be looked up in [42, Table: 2].
5.3 N-Gram Language Modelling with Kneser-Ney Smoothing
This section provides a self-contained introduction to (modified) Kneser-Ney smoothing [62, 25]. Our main purpose is to provide the reader with all necessary details required to reproduce our AV method. For a more in-depth discussion of probabilistic and language modeling aspects, as well as for comparisons with other smoothing algorithms, we refer the reader to the original references [62, 26].
5.3.1 Basic Definitions and Notation
We assume we are given a dictionary $\mathcal{T}\mbox{}$ and a corpus $\mathcal{C}$, the latter consisting of (pre-tokenized) sentences built out of tokens from $\mathcal{T}\mbox{}$. Sequences of tokens of length $n$, or $n$-grams, will be indicated by juxtaposition, such as in $t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n}$, and their observed frequencies will be denoted by:
$c(t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})=\text{Number of occurrences of $t\mbox{}_{1}t\mbox{}_{2}\dots t\mbox{}_{n}$ in $\mathcal{C}$.}$ (5)
[…]
$c(\langle\text{BOS}\rangle\langle\text{BOS}\rangle\cdots\langle\text{BOS}\rangle)=c(\langle\text{EOS}\rangle)=\text{Number of sentences in }\mathcal{C}$
[…]
$c(\langle\text{BOS}\rangle\langle\text{BOS}\rangle\cdots\langle\text{BOS}\rangle t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})\text{ and }c(t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n}\langle\text{EOS}\rangle)$
are given by the number of sentences that started and ended with the string $t\mbox{}_{1}t\mbox{}_{2}\dots t\mbox{}_{n}$, respectively. Finally, for notational convenience, we define the left trimming operator $L$ that maps $n$-grams into $(n-1)$-grams:
$L(t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})$ $=t\mbox{}_{2}t\mbox{}_{3}\cdots t\mbox{}_{n},$ (6)
$L(t\mbox{})$ =$\emptyset$= ∅ (7)
where $\emptyset$ denotes the empty 00-gram.
5.3.2 Prefix and Suffix Counts
[…]
$N_{r}(\bullet g)$ =|$=\left|\left\{t\mbox{}\in\mathcal{T}\mbox{}\cup\{\langle\text{BOS}\rangle\}\colon\,c(t\mbox{}g)=r\right\}\right|,$ (8)
$N_{r}(g\bullet)$ =|$=\left|\left\{t\mbox{}\in\mathcal{T}\mbox{}\cup\{\langle\text{EOS}\rangle\}\colon\,c(gt\mbox{})=r\right\}\right|,$ (9)
and, correspondingly:
$N_{r+}(\bullet g)$ =|$=\left|\left\{t\mbox{}\in\mathcal{T}\mbox{}\cup\{\langle\text{BOS}\rangle\}\colon\,c(t\mbox{}g)\geq r\right\}\right|,$ (10)
$N_{r+}(g\bullet)$ =|$=\left|\left\{t\mbox{}\in\mathcal{T}\mbox{}\cup\{\langle\text{EOS}\rangle\}\colon\,c(gt\mbox{})\geq r\right\}\right|,$ (11)
where ||$\cdot$|| ⋅ | specifies the cardinality of a set. In plain words, these functions provide the number of different prefixes or suffixes a $n$-gram has been observed with exactly, or at least, $r$ times. Notice that all definitions automatically extend to the case where $g=\emptyset$, and we will use the shorthand notations $N_{x}(\bullet)\equiv N_{x}(\emptyset\bullet)$ and $N_{x}(\bullet\bullet)\equiv N_{x}(\bullet\emptyset\bullet)$. The juxtaposition of the empty 00-gram to other $n$-grams is understood as no operation, e. g., $t\mbox{}\emptyset t\mbox{}^{\prime}\equiv t\mbox{}t\mbox{}^{\prime}$.
5.3.3 N-gram Language Models
$N$-gram language models provide a probabilistic description of language production that relies on two essential simplifying assumptions:
  • Sentences are independent of each other.
  • Tokens are conditionally independent of previous history, given a finite (fixed) number, say $N$, of preceding tokens.
According to these assumptions, the probability of sampling a document $\mathcal{D}_{m}$, consisting of $m$ sentences $S\mbox{}_{1},\,S\mbox{}_{2},\,\dots,\,S\mbox{}_{m}$, from a stream of natural language would be given by:
$P(\mathcal{D}_{m}|m)=\prod_{i=1}^{m}P(S\mbox{}_{i})$ (12)
where, in turn, the probability of a sentence $S\mbox{}_{i}=t\mbox{}_{i,1}t\mbox{}_{i,2}\cdots t\mbox{}_{i,n_{i}}$ is:
$P(S\mbox{}_{i})=\prod_{j=1}^{n_{i}+1}P(t\mbox{}_{i,j}|t\mbox{}_{i,j-N+1}t\mbox{}_{i,j-N+2}\cdots t\mbox{}_{i,j-1})$ (13)
and we make the conventions $t\mbox{}_{i,n_{i}+1}=\langle\text{EOS}\rangle$, $t\mbox{}_{i,-1}=t\mbox{}_{i,-2}=\cdots=t\mbox{}_{i,-N+1}=\langle\text{BOS}\rangle$. A detail that is often overlooked in the literature is that the probability in Eq. (12) is conditional on the number $m$ of sentences that compose the document $\mathcal{D}_{m}$, as reflected by our notation.
5.3.4 Kneser-Ney Smoothing
We are now ready to present the (modified) Kneser-Ney $n$-gram model of order $N$. The algorithm follows a general recursion scheme:
$p_{\text{KN}}(t\mbox{}|g)$ $=\alpha(t\mbox{}|g)+\gamma(g)\cdot p_{\text{KN}}(t\mbox{}|L(g)),$ (14)
$p_{\text{KN}}(t\mbox{}|\emptyset)$ $=\alpha(t\mbox{}|\emptyset)+\gamma(\emptyset)\cdot\dfrac{1}{|\mathcal{T}\mbox{}|+1}.$ (15)
The functions $\alpha$ and $\gamma$ are defined by:
$\alpha(t\mbox{}|g)=\begin{cases}\dfrac{\max(c_{\text{KN}}(gt\mbox{})-D(c_{\text{KN}}(gt\mbox{})),\,0)}{\sum_{t\mbox{}^{\prime}\in\mathcal{T}\mbox{}^{*}}c_{\text{KN}}(gt\mbox{}^{\prime})}&\text{if }c(g)>0,\\ 0&\text{otherwise},\end{cases}$ (16)
and:
$\gamma(g)=\begin{cases}\dfrac{\sum_{k=1}^{\infty}D(k)N_{n}(g\bullet)}{\sum_{t\mbox{}^{\prime}\in\mathcal{T}\mbox{}^{*}}c_{\text{KN}}(gt\mbox{}^{\prime})}&\text{if }c(g)>0,\\ 1&\text{otherwise},\end{cases}$ (17)
where $\mathcal{T}\mbox{}^{*}\equiv\mathcal{T}\mbox{}\cup\{\langle\text{EOS}\rangle\}$. In turn, the modified count functions $c_{\text{KN}}(\cdot)$ are given by:
$c_{\text{KN}}(t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})=\begin{cases}c(t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})&\text{if }n=N,\\ N_{1+}(\bullet t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})&\text{otherwise}.\end{cases}$ (18)
Finally, $D(1),\ D(2),\,\dots$ are free parameters satisfying $10<D(k)<$ 1, whose optimal values can be estimated from data. The parameter space is usually made finite by assuming $D(k)=D(r)$ for some $r$ and $k\geq r$, in which case we can rewrite the infinite sum in (17) as:
$\sum_{k=1}^{\infty}D(k)N_{n}(g\bullet)=\sum_{k=1}^{r-1}D(k)N_{n}(g\bullet)+D(r)N_{r+}(g\bullet).$
In our work we use $r=3$, as suggested by the original reference [26]. We conclude the Section with a few remarks:
  • The modified count (18) treats differently $n$-grams of top and lower orders, respectively. We can get some intuition about the usage of the prefix counts $N_{1+}(\bullet t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})$ for lower orders $n$-grams by noticing that these put more weights into tokens that are observed with many contexts, so that we may interpret $N_{1+}(\bullet t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})$ as (being proportional) to a kind of continuation probability[26]. If we want to insist with such a probabilistic interpretation, we may observe that e. g., $\frac{N_{1+}(\bullet t\mbox{})}{N_{1+}(\bullet\bullet)}$ estimates the probability that, out of all distinct bigrams $t\mbox{}_{1}t\mbox{}_{2}$ found in a text corpus, the second token $t\mbox{}_{2}=t\mbox{}$.
  • The conditional probability $p_{\text{KN}}(t\mbox{}|g)$ is defined for $t\mbox{}\in\mathcal{T}\mbox{}^{*}$, which is the reason why the denominator in the second term of (15) is $|\mathcal{T}\mbox{}|+1$. In particular $p_{\text{KN}}(\langle\text{EOS}\rangle|g)$ gives the probability that the $N-1$ gram $g$ terminates the sentence, whereas $p_{\text{KN}}(\langle\text{BOS}\rangle|g)$ is not defined (and indeed has no meaningful interpretation).
  • Kneser-Ney smoothing as originally formulated [62] uses a constant discount function $D(k)\equiv D$, corresponding to $r=1$ in the present formulation. The proposal of a non-constant discount function $D(k)$ was advanced by Chen and Goodman [26]. In the present paper we used a constant discount value of $D=0.75$.
5.4 The Likelihood Ratio Framework
The Likelihood Ratio Framework for forensic sciences approaches the forensic analysis as an assessment of two competing hypotheses. Typically, in authorship verification, these hypotheses are:
  • the Prosecution Hypothesis ($H_{p}$): The disputed document ($\mathcal{D}_{\mathcal{U}}$) was written by the author of the known document(s) (the defendant, $\mathcal{A}$).
  • the Defence Hypothesis ($H_{d}$): The disputed document ($\mathcal{D}_{\mathcal{U}}$) was not written by the author of the known document(s) (the defendant, $\mathcal{A}$), but was written by someone else, $\neg\mathcal{A}$
The police’s or stakeholders’ interest in the given evidence ($E$), comprising $\mathcal{D}_{\mathcal{U}}$ and $\mathcal{D}_{\mathcal{A}}$, is probabilistically expressed as the ratio of the following two Bayesian conditional probabilities:
${\text{odds}_{\text{post}}=\frac{P(H_{p}|E=\{\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}}\})}{P(H_{d}|E=\{\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}}\})}}$ (19)
The numerator of Equation 19 states Given the evidence, what is the probability of $H_{p}$ being true? The denominator of Equation 19 states Given the same evidence, what is the probability of $H_{d}$ being true? If the ratio is greater than one, the evidence supports $H_{p}$ more than $H_{d}$; the larger than one is the ratio, the stronger the support for $H_{p}$. Similarly, if the ratio is smaller than one, the evidence supports $H_{d}$ more than $H_{p}$; the smaller than one is the ratio, the stronger the support for $H_{d}$. The ratio provided in Equation 19 is known as the posterior odds in Bayes’ Theorem. These odds can be estimated using Bayes’ Theorem, which, in its odds form is given in Equation 20.
${\underbrace{\frac{P(H_{p}\ |E)}{P(H_{d}\ |E)}}_{\text{posterior\ odds}}\ =\underbrace{\frac{P(H_{p})}{P(H_{d})}}_{\text{prior\ odds}}\times\underbrace{\frac{P(E|H_{p})}{P(E|H_{d})}}_{\text{likelihood\ ratio}}}$ (20)
Equation 20 states that posterior odds can be estimated as the product of the prior odds and the likelihood ratio. The prior odds represents the trier-of-fact’s belief about the two competing hypotheses ($H_{p}$ and $H_{d}$) before observing the evidence (i.e., linguistic text evidence). The forensic scientist’s task is to evaluate the strength of the evidence, quantified in the form of a likelihood ratio. In other words, Equation 20 states that the belief of the trier-of-fact (prior odds), which was formed or influenced by other evidence presented prior to the linguistic text evidence, is updated through the evaluation of the linguistic text evidence (LR), resulting in an up-to-date belief (posterior odds).
It is important to note that the forensic expert cannot estimate and legally must not refer to the posterior odds. They cannot estimate it logically because they are not privy of the trier-of-fact’s belief regarding the hypotheses and do not have access to other pieces of evidence presented prior to the linguistic text evidence. Legally, they must not refer to it because the posterior odds represents the ultimate issue of being guilty or innocence [73, p. 96] and this determination is a specific responsibility reserved only for the trier-of-fact. The odds form of Bayes’ Theorem explicitly outlines the role of the forensic scientist. In the Likelihood Ratio framework, their responsibility lies in estimating the strength of evidence in the form of a likelihood ratio. As indicated by the rightmost term of Equation 20, the likelihood ratio is the ratio between the probability of the evidence given $H_{p}$ and the probability of the same evidence given $H_{d}$. The Likelihood Ratio serves as a gradient, quantifying how much more likely the evidence is given one hypothesis over the other [35]. When the LR is greater than one, it provides stronger support for $H_{p}$, whereas when it is smaller than one, it offers stronger support for $H_{d}$. The LR framework is now accepted as a logically and legally correct framework for interpreting forensic evidential analyses [3, 94], having been studied for various evidential types [15, 16, 29, 77, 81, 113]. The LR framework has also been supported by the relevant scientific and professional associations [2, 85, 23, 92, 6, 84].
5.4.1 The Log-Likelihood Ratio Cost ($C_{\text{llr}}\mbox{}$)
In numerous source-identification tasks, the system’s performance is commonly evaluated in terms of identification accuracy and/or identification error rate. Metrics such as precision, recall, and equal error rate (EER) are typical in this regard. While these metrics offer a quick overview of performance and are intuitively easy to comprehend, they are not suitable for assessing likelihood ratio-based inference systems [78, p. 93].
Primarily based on categorical decisions with specific thresholds, these metrics are inadequate for evaluating such systems. Firstly, metrics relying on categorical thresholding implicitly assess how effectively the system correctly identifies the source of evidence. In other words, these metrics implicitly pertain to the strength of the hypothesis rather than the strength of the evidence [78, p. 93]. Secondly, categorical metrics fail to capture the gradient nature of likelihood ratios [78, p. 93]. For instance, in a same-author comparison, if one system returns $\log_{10}(\text{LR})=3$ while another yields $\log_{10}(\text{LR})=2$, the former is superior as it supports the correct hypothesis more strongly. However, using a categorical metric would assess both systems in the same way, disregarding the differing strengths of evidence supporting the correct hypothesis.
The conventional metric employed for evaluating the performance of a system within the Likelihood Ratio Framework is the log-likelihood ratio cost ($C_{\text{llr}}\mbox{}$). The computation of the $C_{\text{llr}}\mbox{}$ involves utilising a set of LRs for same-source scenarios (e.g., same-author, Y) and a set for different-source scenarios (e. g., different-author, N). The computation is performed using the following equation [20, 91, 111]
$C_{\text{llr}}\mbox{}=\frac{1}{2}\left(\left[\frac{1}{n}\sum_{i}^{n}\log_{2}\left(1+\frac{1}{LR_{i}}\right)\right]+\left[\frac{1}{m}\sum_{j}^{m}\log_{2}\bigg{(}1+LR_{j}\bigg{)}\right]\right)$ (21)
where, $n$ and $m$ are the number of Y- and N-cases, and $LR_{i}$ and $LR_{j}$ are the LRs estimated for the Y- and N-cases, respectively. $C_{\text{llr}}\mbox{}$ considers the magnitude of calculated LR values, imposing appropriate penalties based on their values. Within the $C_{\text{llr}}\mbox{}$ metric, LRs supporting counter-factual hypotheses, or contrary-to-fact LRs (with $\log_{10}(\text{LR})<0$ for Y-problems and $\log_{10}(\text{LR})>0$ for N-problems), incur significant penalties. The extent of the penalty is proportional to the deviation of the LRs from unity. Optimal performance is attained when $C_{\text{llr}}\mbox{}$ equals 0, and performance gradually diminishes as $C_{\text{llr}}\mbox{}$ approaches and surpasses 1. Therefore, a lower $C_{\text{llr}}\mbox{}$ value corresponds to better performance.
$C_{\text{llr}}\mbox{}$ serves as an assessment metric for the overall performance of an LR-based inference system and can be decomposed into two distinct metrics: $C_{\text{llr}}^{\text{min}}\mbox{}$ and $C_{\text{llr}}^{\text{cal}}\mbox{}$. The former evaluates the discrimination performance and the latter the calibration performance of the system. The $C_{\text{llr}}^{\text{min}}\mbox{}$ is determined by computing the losses for the optimised LRs through the non-parametric pool-adjacent-violators algorithm. The disparity between $C_{\text{llr}}\mbox{}$ and $C_{\text{llr}}^{\text{min}}\mbox{}$ is expressed as $C_{\text{llr}}^{\text{cal}}\mbox{}$; in other words, $C_{\text{llr}}\mbox{}$ equals the sum of discrimination loss ($C_{\text{llr}}^{\text{min}}\mbox{}$) and calibration loss ($C_{\text{llr}}^{\text{cal}}\mbox{}$). More in-depth descriptions of $C_{\text{llr}}\mbox{}$, encompassing both $C_{\text{llr}}^{\text{min}}\mbox{}$ and $C_{\text{llr}}^{\text{cal}}\mbox{}$, can be found in work by Brümmer and du Preez [20], Drygajlo et al. [34], and Meuwly et al. [75]. The $C_{\text{llr}}\mbox{}$ has already been effectively utilized for evaluating LR-based forensic text comparison systems [49, 50, 51, 82].
5.5 SiamBERT: Neural Net Baseline
Recognising the paradigm shift towards neural network techniques that is also evident in the research field of AV, we consider it important to compare LambdaG with contemporary methods utilizing deep learning techniques. For this reason, we incorporate in our evaluation a baseline method based on a Siamese BERT neural network and functionally very similar to the Sentence-BERT architecture [93] and Ishihara’s et al [53] method, in which likelihood-ratios were estimated via cosine similarity and probabilistic linear discrimination analysis. We call this method from here on SiamBERT. Note that, in our experiments, we use the BERT-Large, Cased model with its corresponding tokenizer [30], which are pretrained on English text and considers character casing. Given a document $\mathcal{D}$, we tokenize it into a sequence of tokens. Since the BERT model we are using is only able to handle up to 512 tokens, we split the sequence of tokens into a set of adjacent chunks $\varsigma_{1},\ldots,\varsigma_{m}$, where $m$ is the number of chunks. Each $\varsigma_{i}$ comprises 512 tokens, except for $\varsigma_{m}$, which may comprise less. These chunks are then embedded as feature vectors $F_{1},\ldots,F_{m}$, by encoding each $\varsigma_{i}$ with the BERT model and taking the average of the resulting embedding vectors over the tokens in a chunk. Finally, the feature vectors are then again averaged, resulting in a final feature vector $F$, which represents the document $\mathcal{D}$ in an embedding space. Given a single verification case, we consider an unknown document $\mathcal{D}_{\mathcal{U}}$ and a known document $\mathcal{D}_{\mathcal{A}}$ for SiamBERT. Both are then embedded as described above, resulting in two feature vectors $F_{\mathcal{U}}$ and $F_{\mathcal{A}}$. Based on these vectors, we then calculate their cosine similarity $s$. If $s>\theta$, where $\theta$ is a threshold, we assume $\mathcal{U}=\mathcal{A}$ (i. e., same authorship). To train SiamBERT, we further split the training set $\mathcal{C_{\,\textrm{train}}}$ into $\mathcal{C_{\,\textrm{train}*}}$ and $\mathcal{C_{\,\textrm{val}}}$ based on a 80/20 ratio, in order to optimize hyper-parameters. We trained the parameters of the BERT model on $\mathcal{C_{\,\textrm{train}}}$ with MSE loss function (calculating the mean squared error between the cosine similarity and the binary label, where 1 represents a Y case), the AdamW optimizer, a batch size of $4$ and a learning rate of $0.00005$ for $100$ epochs (scheduled with a linear warmup in the first $10$ epochs and a cosine decay to zero in the remaining). In order to optimize the hyper-parameter epoch count, we measured the AUC on each epoch of the corresponding model on $\mathcal{C_{\,\textrm{val}}}$ and selected the model with the best AUC. Finally, we calculated the cosine similarities of the verification cases of $\mathcal{C_{\,\textrm{train}}}$ using the selected model and determined the optimal threshold $\theta$, which would yield the best accuracy on $\mathcal{C_{\,\textrm{train}}}$.
6 Appendix A: Qualitative Data Exploration
[…]
Ayas Aya AyN Ay, Ayi Ayca Ayn’t AyV Aytoo Aymany Ayof Aythem Aywanting Ayto Aygo Aythe AyJ AyN Ayto AyV Aya AyN Aytoo Ay. Ay[EOS]
Aythis AyN Ayseems Ayto Aybe Ayalready AyB AyV Aya Aylittle Aymore AyB Ayin Ayanother AyN Ay: AyP Ay. Ay[EOS]
AyV Aya AyJ AyN Ayfrom Ayyour AyJ AyN Aywould Ayactually AyV AyN Ayabout Ayyour AyN Ayin Ayyour AyJ AyN Ay, Aythus AyV Aythe AyN Ayof AyN Ayyou Ay’re Aylooking Ayfor Ay( Ayat AyJ Ayas AyV Ayin Aya AyJ AyN Ay) Ay. Ay[EOS]
Ayif Ayyou Aymoved Ayto Aya AyJ AyN Ay( Aybe Ayit AyN Ayor AyN Ay) Ayfor Aya AyN Ay, Ayyou Aywill AyB Ayrun Ayinto Aysome AyJ AyN Aywith AyN Ay. Ay[EOS]
Aythis Ayis AyV Ay- Ayand Ay- AyJ AyN Ay. Ay[EOS]
The first example is a unknown document $\mathcal{D}_{\mathcal{U}}$ taken from $\mathcal{C}_{\mathrm{Stack}}$ where the sentences above are the five sentences with the highest $\lambda_{G}$. Although the important tokens are highlighted following a color scheme, the interpretation of why they are important is left to the analyst using a manual analysis of the concordance lines. An exploration of the corpus reveals that it is the entire grammatical construction “as a N,” at the beginning of a sentence that is rare in the reference dataset but it is found in both the known and unknown documents. The use of “J N with N .” is only used by this author and another author in the entire corpus. Similarly, the phrase “you’re looking for” is rare and only used by four other authors in the corpus. The combination of these three grammatical constructions is therefore already uniquely identifying this individual.
Ayfinally Ay, Ayit Ayputs Ayan Ayever Ay- Ayso Ay- AyJ AyN Ayof AyN Ayon Aythe AyN Ayby AyV Aythe AyN Ayso Aythat Aythey Ay’re Aytrying Ayto AyV Ayyou Ay, Ayrather Aythan Aythe Ayother AyN Ayaround Ay. Ay[EOS]
Ayto Aybegin Aywith Ay, Ayit Ay’s Aynot AyN Ayif Ayyou AyB AyV Ayyour AyN Ay. Ay[EOS]
Ayit Aywould Aybe AyJ Ayto AyV Aythat Ayyou Ay’re Aythe Ayfirst Ayto AyV Aya AyN Aybetween AyN Ayand AyN AyN Ay, Ayfor AyN Ay, Aybut Ayit Ay’s AyB AyJ AyJ Ayother AyN Ayon Aythe AyN Ayand Aythen AyV Ayor AyV Aythem Aywith Ayyour AyJ AyN Ay. Ay[EOS]
Ayin Aythe Ayvery AyJ AyN Ayof AyJ AyN Ay, Ayyou Ay’d Ayneed Ayto AyV AyP Ay’s AyN Ayto Ayget Aya AyN Ayto Aydo Aywhat Ayyou Aywant Ay. Ay[EOS]
Ayyou Aycould AyV Aythe AyJ AyN Ayunder Ayyour AyN AyN Ayto AyV Ayan AyN Ayof Ayyour AyN Ayof AyN Ay, Aybut Aythe AyN Ayalone Aydoes Ayn’t AyV Ayany AyN Ay. Ay[EOS]
In this other $\mathcal{D}_{\mathcal{U}}$ from $\mathcal{C}_{\mathrm{Stack}}$, some of most important sequences highlighted by the model are the use of “finally,” and “to begin (with)” at the beginning of a sentence. Another important sequence that is hard to spot immediately and that becomes evident only after more manual investigation is the construction “get…to do what” in the fourth sentence, which is only attested for this author.
Ayi Ayam Aya Aylittle AyJ Aybetween Aymy AyN Ay[EOS]
Ayso Aywhat Aydid AyP Aydo Ay? Ay[EOS]
Ayhave Ayyou Aywanted Ayto Aybe Ay? Ay[EOS]
Ayhow Aydo Ayi Ayget Aythere Ay? Ay[EOS]
Ayshe Aydid Ayn’t Ayknow Aywhat Ayto Aythink Ayat Ayfirst Ay. Ay[EOS]
This is an example from $\mathcal{C}_{\mathrm{PJ}}$, which is the easiest of the corpora to interpret, probably because of the non-standard, unedited nature of the language used in this genre. In this example, the author uses the phrase “you wanted” frequently, as well as the 5-gram “i am a little J”, which is only found in one other author in the corpus. The most interesting pattern is the phrase “she did n’t”, which seems completely unremarkable but it is used only by two other authors in the entire corpus and never at the beginning of a sentence. This could be taken as an instance of content interference but, in reality, there are 417 occurrences of the pronoun “she” in the corpus. Instead, this is an example that reflects a larger unit for this author: beginning a sentence, which in this case is a new chat message, using the pronoun and the contracted form of “did not” with an apostrophe and then followed by a mental verb (“think” in the known sample) to describe a mental state.
Ayit Ayis AyJ Ayenough Ayto Aybe AyB AyV Ayin Ayany AyJ AyN Ay. Ay[EOS]
Ayhowever Ay, Aythis AyN Ayof AyN Aydoes Aynot AyV AyV Aythe AyN Ay, Ayfind Aythe AyN Ayfrom Aythe AyN Ay: Aythe AyN AyN Ay, Ayunlike Ayother AyN Ay, Ayis Aynot AyJ Ay. Ay[EOS]
Aythe Ayfirst Ayof Aythese AyN Ayis Aya Ayset Ayof AyV AyN AyV Aywith Aya AyN Ayof AyN AyN Ay: AyN Ayabyssal Aya Ay- AyN AyN AyV AyP Ay- AyN Aythe AyN Ayof Aythis AyN Ayof AyN Ayin AyJ AyP Ayis Ayquite AyJ Ay. Ay[EOS]
Aythe Ayonly AyN Aywe Ayset Ayis AyN Ay: Aythe AyN Ayof AyN Ay, Ayeven Ayif Ayit Ayis AyJ Ay, AyV Aythe AyN Ayof Ayother AyN Ay. Ay[EOS]
This is an example from $\mathcal{C}_{\mathrm{ACL}}$. Apart from the use of “however,” at the beginning of a sentence, this author uses the sequence “is (not) J enough to be B V”, which is only found in their writing.
Aythe AyV AyN Ayis Aya AyJ AyN AyN Aythat AyV Ayvarious AyJ AyJ AyN Aythat Aycan Aybe AyB AyV Ayto AyV AyN AyJ AyN Ay. Ay[EOS]
Aythe AyN AyV Aythe AyN Ay- AyN AyJ AyN AyP Ay, Aythe AyJ AyN Aybetween AyP Ayand AyP AyN AyN Ayof AyN Ay, AyJ AyN Ay, Ayand AyN AyN AyN Ay. Ay[EOS]
AyV AyN AyV Ayby AyN Ay, Aythe AyP Ay- AyP AyN AyV AyJ AyN Ayin Aythe Aygiven Ayset Ayof AyN Ay, Aywhether Aythey Ayme Aydue Ayto AyJ Ay, AyJ Ay, AyJ Ay, Ayor AyN AyN Ay. Ay[EOS]
Aythe Ayother AyN Ayis Aythe AyJ AyN Ayof Aythe AyN AyN Ay. Ay[EOS]
AyJ AyN AyP Ayand AyP Ayare AyB AyV Ayfor AyV AyJ AyJ AyN Ayof Ayeach AyN Ay. Ay[EOS]
This example, again from $\mathcal{C}_{\mathrm{ACL}}$, reveals the sequence “can be B V to”, which is only used by two other authors and the construction “the other N” at the beginning of a sentence.
Aythis AyN AyV AyB AyJ AyN Ay, Aynow Ay, Aythey AyV Aylike AyP Ayso Ayi Ayca Aynt Aysay Aythey Ayare Ayall AyJ Aybut Aythe AyJ Ayfew Ayi Aytried Aymade Ayit AyD AyN Ayin Aymy AyN Ayhad Aythe AyN AyN Aywith AyV AyN Ay( AyB AyV Ay) Ay. Ay[EOS]
AyD AyN AyD Ay) AyJ AyP Ayin AyJ AyN Ay- AyD AyN Ay, AyV AyJ AyN AyN Ayin Aya AyJ AyN Ay, Ayover AyV AyN AyD Ay) AyP AyP Ayover AyN Ay- AyD AyN Ay, AyJ Ayfor Aya AyN Ayunder AyS AyD Ay, Aycomes Aywith AyN Ayand AyV AyJ AyN AyD Ay) AyP AyP Ay, AyV AyN Ay- AyJ Aybut Aya AyN AyJ Ayfor Aya AyN Ayplus AyN Ay. Ay[EOS]
Ay1 Ay) Aythe AyN Ay- AyN AyN Ay, AyN AyN Ay, AyP AyP AyP Ay, Ayand Aya AyN Ayfor Ayless Aythan AyD AyN Ay. Ay[EOS]
Aynow Ay, Aythe AyN Ayset Ayis AyJ Ay, Aybut Ayi AyV Aythe AyN Ayand AyN Ayof Aythe AyJ Ay. Ay[EOS]
Aybut Aydefinitely AyJ Ay! Ay[EOS]
Ayit Aywas AyJ Ayas AyN Aycould Aybe Ayin AyP Ay. Ay[EOS]
Finally, the last example is from $\mathcal{C}_{\mathrm{Yelp}}$, and it is a $\mathcal{D}_{\mathcal{U}}$ where the most obvious evidence is constituted by punctuation mark usage, such as dashes and commas, and by the contracted form of “can’t” without an apostrophe. This token is marked as very important because the overarching sequence that identifies this author is the use of “can’t” in combination with “so” with an intervening pronoun, a sequence only found in one other author in the corpus.

### v2

@@
-Authorship Verification based on the Likelihood Ratio of Grammar Models
+Grammar as a Behavioral Biometric: Using Cognitively Motivated Grammar Models for Authorship Verification
 Abstract
-Authorship verification (AV) is the process of analyzing a set of documents to determine whether they were written by a specific author.
-This problem often arises in forensic scenarios, e.g., in cases where the documents in question constitute evidence for a crime.
-Existing state-of-the-art AV methods use computational solutions that are not supported by a plausible scientific explanation for their functioning and that are often difficult for analysts to interpret.
-To address this, we propose a method relying on calculating a quantity we call $\lambda_{G}$ (LambdaG): the ratio between the likelihood of a document given a model of the Grammar for the candidate author and the likelihood of the same document given a model of the Grammar for a reference population.
-These Grammar Models are estimated using $n$-gram language models that are trained solely on grammatical features.
-Despite not needing large amounts of data for training, LambdaG still outperforms other established AV methods with higher computational complexity, including a fine-tuned Siamese Transformer network.
-Our empirical evaluation based on four baseline methods applied to twelve datasets shows that LambdaG leads to better results in terms of both accuracy and AUC in eleven cases and in all twelve cases if considering only topic-agnostic methods.
-The algorithm is also highly robust to important variations in the genre of the reference population in many cross-genre comparisons.
-In addition to these properties, we demonstrate how LambdaG is easier to interpret than the current state-of-the-art.
-We argue that the advantage of LambdaG over other methods is due to fact that it is compatible with Cognitive Linguistic theories of language processing.
+Authorship Verification (AV) is a key area of research in digital text forensics, which addresses the fundamental question of whether two texts were written by the same person.
+Numerous computational approaches have been proposed over the last two decades in an attempt to address this challenge.
+However, existing AV methods often suffer from high complexity, low explainability and especially from a lack of clear scientific justification.
+We propose a simpler method based on modeling the grammar of an author following Cognitive Linguistics principles.
+These models are used to calculate $\lambda_{G}$ (LambdaG): the ratio of the likelihoods of a document given the candidate’s grammar versus given a reference population’s grammar.
+Our empirical evaluation, conducted on twelve datasets and compared against seven baseline methods, demonstrates that LambdaG achieves superior performance, including against several neural network-based AV methods.
+LambdaG is also robust to small variations in the composition of the reference population and provides interpretable visualizations, enhancing its explainability.
+We argue that its effectiveness is due to the method’s compatibility with Cognitive Linguistics theories predicting that a person’s grammar is a behavioral biometric.
 1 Introduction
@@
 g., blackmail letters, suicide notes, letters of confession, wills or theses potentially written by ghostwriters.
-Over the course of time, various research disciplines have emerged to deal with the analysis of the language of authors from different perspectives.
-This has led to an interdisciplinary branch of research that is often referred to as digital text forensics in the literature (e.
+Over time, various research disciplines have emerged to deal with the analysis of the language of authors from different perspectives.
+This has led to an interdisciplinary branch of research that is often referred to as digital text forensics (e.
 g., [103, 14, 49, 109, 47]).
@@
 AV, on the other hand, deals (in its simplest form) with the task of deciding whether two documents were written by the same author.
-Our focus in this paper is on AV, which represents a reformulation of AA, based on the fact that any AA problem can be decomposed into a series of AV problems [67].
+In this paper, we focus on AV, which can be seen as a reformulation of AA, given that any AA problem can be decomposed into a series of AV problems [81].
 1.1 Decision Problems
@@
   2. 2.
-    $\textrm{AV}_{\,\textrm{Batch}}$: Given two sets of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}^{\,(a)}_{1},\mathcal{D}^{\,(a)}_{2},\ldots\}$ and $\mathbb{D}_{\mathcal{B}}=\{\mathcal{D}^{\,(b)}_{1},\mathcal{D}^{\,(b)}_{2},\ldots\}$ each of which written by a single author: Are the documents in $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{D}_{\mathcal{B}}$ written by the same author?
+    $\textrm{AV}_{\,\textrm{Batch}}$: Given two sets of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}^{\,\mathcal{A}}_{1},\mathcal{D}^{\,\mathcal{A}}_{2},\ldots\}$ and $\mathbb{D}_{\mathcal{B}}=\{\mathcal{D}^{\,\mathcal{B}}_{1},\mathcal{D}^{\,\mathcal{B}}_{2},\ldots\}$ each of which written by a single author: Were the documents in $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{D}_{\mathcal{B}}$ written by the same author?
   3. 3.
-    $\textrm{AV}_{\,\textrm{Known}}$: Given a set of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots\}$ by the same author $\mathcal{A}$ as well as a document $\mathcal{D}_{\mathcal{U}}$ of an unknown author $\mathcal{U}$: Has $\mathcal{A}$ also written $\mathcal{D}_{\mathcal{U}}$ such that $\mathcal{A}=\mathcal{U}$ holds?
+    $\textrm{AV}_{\,\textrm{Known}}$: Given a set of documents $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots\}$ by the same author $\mathcal{A}$ as well as a document $\mathcal{D}_{\mathcal{U}}$ of an unknown author $\mathcal{U}$: Has $\mathcal{A}$ also written $\mathcal{D}_{\mathcal{U}}$ (i.
+e., $\mathcal{A}=\mathcal{U}$)?
 In a forensic context the process of assisting the trier-of-facts in finding answers to these problems is referred to as Forensic Text Comparison (FTC) [62].
@@
 In all decision problems, including $\textrm{AV}_{\,\textrm{Core}}$ where the denotation of which document is known or unknown can be arbitrary, an AV method seeks to answer the question of whether $\mathcal{A}=\mathcal{U}$ ($\mathcal{A}$ authored $\mathcal{D}_{\mathcal{U}}$) or $\mathcal{A}\neq\mathcal{U}$ ($\mathcal{A}$ did not author $\mathcal{D}_{\mathcal{U}}$) holds.
-In this context, a known verification case with a matching and non-matching authorship is denoted by Y- and N-case, respectively.
+In this context, a known verification case with matching and non-matching authorship is denoted by Y- and N-case, respectively.
 Moreover, we denote $\neg\mathcal{A}$ as any author other than $\mathcal{A}$.
-1.2 Applications of Authorship Verification
-For all the mentioned forms of decision problems, numerous AV applications can be found across different domains.
-In forensic linguistics, for example, AV can be used to analyze malicious texts such as ransom notes, blackmail letters or other non-malicious texts that are used as evidence in investigatory, criminal or civil contexts.
-In the field of journalism and media, AV can play a vital role when it comes to confirming the authorship of news articles or opinion pieces or investigating cases of ghostwriting and unauthorized publication of content.
-In academia, AV can be used to analyze research articles, dissertations or other written content, for example, to resolve a plagiarism claim, or contract cheating from students, and thereby overall contribute to combating academic misconduct.
-In the context of social media, AV can also be employed to detect fake accounts or multiple accounts belonging to the same user.
-Moreover, as part of email and communication analysis, AV can be used to verify the authorship of emails in cases of fraud, harassment or cybercrime.
-Finally, in the Digital Humanities, AV can be used to resolve matters that deal with disputed historical evidence or to provide evidence in favour or against the origin of literary works.
+1.2 Applications of AV
+AV can be used in many different areas and offers a wide range of possible applications.
+In forensic linguistics, AV is instrumental in analyzing texts related to criminal activity, such as ransom notes, blackmail letters, or other documents used as evidence in investigative, criminal, or civil cases.
+For example, Iqbal et al.
+[60] used a NIST-based speaker recognition system to verify the authorship of malicious emails in forensic contexts.
+Similarly, Remmide et al.
+[107] employed AV to detect phishing attacks.
+In journalism and media, AV plays a vital role in authenticating the authorship of news articles and opinion pieces, as well as detecting ghostwriting [4].
+In academia, AV is crucial for detecting academic misconduct, such as contract cheating [122] and plagiarism [42].
+Social media platforms also benefit from AV, especially when it comes to detecting fake accounts or multiple accounts operated by the same person.
+In this context, Igawa et al.
+[59] and Barbon et al.
+[67] demonstrated the effectiveness of AV regarding the identification of compromised accounts on X/Twitter.
+Lundmark et al.
+[87], on the other hand, used AV to link user accounts on discussion forums, whereas Weerasinghe et al.
+[134] applied AV to mitigate abuse in online communities.
+Finally, in the Digital Humanities, AV has also become an important instrument for clarifying questions about the authorship of historical documents and authenticating literary works.
+Kestemont et al.
+[72], for example, applied AV to authenticate historical writings attributed to Julius Caesar, while Tuccinardi [126] used AV methods to investigate the authenticity of Pliny the Younger’s letter to Emperor Trajan.
 1.3 Categories of AV Methods
-According to Halvani et al.
-[43, 47], AV methods can be divided into three possible categories, namely unary, binary-intrinsic and binary-extrinsic AV methods [^1] A summarized visual representation of these three categories can be seen in [47, Figure 1].
+AV methods can be divided into three different categories: unary, binary-intrinsic and binary-extrinsic [54, 58].
 All three rely on a decision criterion $\theta$, in order to accept or reject $\mathcal{A}$ as the author of $\mathcal{D}_{\mathcal{U}}$.
@@
 These methods assume $\mathcal{D}_{\mathcal{U}}$ to be written by $\mathcal{A}$, if it is similar enough to the documents in $\mathbb{D}_{\mathcal{A}}$, where the decision is only based on these and $\theta$.
-  • Binary-intrinsic AV methods require a given training corpus, consisting of verification cases with a ground truth regarding the Y- (same-authorship) and N-classes (different-authorship), in order to determine $\theta$.
+  • Binary-intrinsic AV methods require a given training corpus, consisting of verification cases with a ground truth regarding the same-authorship (Y) and different-authorship (N) classes, in order to determine $\theta$.
 Binary-intrinsic AV methods treat verification cases that include the unknown and known document(s) of $\mathcal{A}$, as a single unit $X$ (e.
@@
 Koppel and Winter [84], for example, used two approaches: (1) collecting documents using search engine queries and (2) using a fixed set of already crawled documents.
-Binary-extrinsic AV methods assume $\mathcal{D}_{\mathcal{U}}$ to be written by $\mathcal{A}$ only if $\mathcal{D}_{\mathcal{U}}$ is consistently more similar to $\mathcal{D}_{\mathcal{A}}$ (or alternatively, the documents in $\mathbb{D}_{\mathcal{A}}$) despite the presence of distractor impostor documents $\{I_{1},I_{2},\ldots\}$ after calculating a similarity score using a bootstrapping approach.
-In a forensic context, the adherence to the likelihood ratio paradigm for forensic science implies that an AV method does not return $\theta$ but instead returns a likelihood ratio that quantifies the strength of the evidence in favour or against the two propositions.
+Binary-extrinsic AV methods assume $\mathcal{D}_{\mathcal{U}}$ to be written by $\mathcal{A}$ only if $\mathcal{D}_{\mathcal{U}}$ is consistently more similar to $\mathcal{D}_{\mathcal{A}}$ (or alternatively, the documents in $\mathbb{D}_{\mathcal{A}}$) than any of a number of distractor impostor documents $\{I_{1},I_{2},\ldots\}$ using a bootstrapping approach.
+In a forensic context, the adherence to the likelihood ratio paradigm for forensic science implies that an AV method does not return a Y- or N-decision but a likelihood ratio that quantifies the strength of the evidence in favor or against the two propositions.
 In addition to the three types above, therefore, it is also possible to classify an AV method as forensic when the method returns a likelihood ratio.
-In a review of the field, Stamatatos [102] concluded that binary-extrinsic methods tend to outperform the other categories.
-In particular, a binary-extrinsic AV method that has been found to be very successful over several independent studies is the well-known Impostors Method (IM).
-The basic idea behind IM can be traced back to another successful AV method introduced by Koppel and Schler [65] and called Unmasking and later modified by Koppel et al.
+1.4 Existing AV Solutions
+This section describes the advantages and limitations of some well-known existing AV methods that have demonstrated a consistent degree of effectiveness over time.
+This section covers in particular those methods that have found significant validation across the yearly competitions held at the PAN series of conferences focused on stylometry and digital text forensics.
+Since its launch in 2007, PAN has hosted 73 computational challenges, known as shared tasks, in areas such as authorship analysis, computational ethics, and evaluating the originality of writing, as well as curated 57 evaluation datasets, in addition to nine datasets contributed by the community [10].
+1.4.1 The Impostors Method
+In a review of the field, Stamatatos [118] concluded that binary-extrinsic methods tend to outperform other types of methods.
+A widely recognized and successful binary extrinsic AV method that has been repeatedly validated in independent studies is the Impostors Method (IM), which has been proposed by Koppel and Winter [84].
+The basic idea behind IM can be traced back to another successful AV method called Unmasking, initially introduced by Koppel and Schler [79] and later refined by Koppel et al.
 in [82].
-The key insight of IM is to collect a set of texts that are plausible impostors for $\mathcal{D}_{\mathcal{A}}$ and then calculate a similarity score for $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ not once but a large number of times while withholding a random sub-set of the features at every iteration.
-The rationale is that only the real author of the disputed text is so similar to it to withstand random permutations of the features set, which, for IM, is the set of the most frequent character 4-grams.
+Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, the key insight of IM is to collect a set of texts that are plausible impostors for $\mathcal{D}_{\mathcal{A}}$ and then calculate a similarity score for $\mathcal{D}_{\mathcal{U}}$ and $\mathcal{D}_{\mathcal{A}}$ not once but a large number of times while withholding a random sub-set of the features at every iteration.
+The rationale is that only the real author of the disputed text $\mathcal{D}_{\mathcal{U}}$ is so similar to it to withstand random permutations of the features set, which, for IM, is the set of the most frequent character $4$-grams.
 Koppel et al.
-[67] proposed that to choose the impostors the key principle should be to identify a relevant universe, choose linguistically similar texts, and then randomly sample from them.
+[81] suggested that the key principle to choose the impostors should be to identify a relevant population, select linguistically similar texts, and then randomly sample from it.
 IM found immediate external validation in the PAN 2013 competition, where a variant called GenIM (General Impostors Method), proposed by Seidman [113], outperformed all other submitted AV methods.
@@
 [80], GenIM made use of collected impostors document from the web.
-The method then used the relative frequency of single words weighted by their document frequency (a quantity usually called tf-idf) and the Min-Max (aka Ružička or fuzzy Jaccard) similarity.
-Successful results were then found by Koppel and Winter [70] for the $\textrm{AV}_{\,\textrm{Core}}$ decision problem, again using the Min-Max similarity between vectors of character 4-grams tf-idf.
-For the impostors, Koppel and Winter [70] tried various configurations, including with impostors generated on the fly using web searches or sampling from genre-compatible corpora.
-They discovered that genre compatibility permits the use of fewer impostors but that, equally, a large number of impostors ($\approx 200$) generated using web searches can compensates for lack of genre control.
+The method then used the relative frequency of single words weighted by their document frequency (a quantity usually called tf-idf) and the Min-Max (aka Ružička or fuzzy Jaccard) similarity measure.
+Successful results were then found by Koppel and Winter [84] for the $\textrm{AV}_{\,\textrm{Core}}$ decision problem, again using Min-Max applied to vectors of character $4$-grams weighted by tf-idf.
+They explored various configurations for impostors, including dynamically generating them through web searches or sampling from genre-compatible corpora.
+Koppel and Winter discovered that genre compatibility permits the use of fewer impostors but that, equally, a large number of impostors ($\approx 200$) generated through web searches can compensate for the absence of genre control.
 Their results demonstrated that it is possible to solve this challenging two-text only verification problem even for texts as short as 500 words.
@@
 The third issue is not strictly speaking an issue of IM but a well-known problem for both AV and AA, as virtually all methods to solve theses tasks tend to be highly affected by genre variation [8].
-Regarding IM, Halvani [41] found that the performance significantly decreases when the set of impostors is chosen from another genre.
-For example, performing AV on a corpus of chat logs using a corpus of chat logs leads to a performance of 0.92 in accuracy.
+For IM, Halvani [51] found that the performance (in terms of Accuracy [^1] Note that we use the capitalized form ”Accuracy” to clarify that we are talking about a specific machine learning evaluation metric, rather than the general concept of ”accuracy”.) declines significantly when impostors are drawn from a different genre.
+For example, performing AV on a corpus of chat logs achieves an Accuracy of 0.92 when using chat logs as impostors.
 However, when using business emails or academic articles as impostors, this performance drops to almost chance level at 0.54.
-Only when the genres are relatively similar (e.
-g., chat logs compared to blogs) then, although a significant drop in performance is still evident, the results are still above chance (0.74).
+Even with relatively similar genres, such as chat logs and blogs, the performance still decreases notably, though it remains above chance (0.74).
 The last and most significant limitation of IM and of any other AV methods using the frequency of short character $n$-grams is that the most important features of the analysis are not immediately obvious to the human analyst.
@@
 g., $\mathcal{A}$ is likely to be the author of $\mathcal{D}_{\mathcal{U}}$ because $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$ are about “grief therapy”).
-1.4 Effect of Content
-This issue of the effect of content is one of a number of biases identified by Bevendorff et al.
+1.4.2 Content-agnostic Approaches
+This content effect limitation of IM is also one of a number of biases identified by Bevendorff et al.
 [14] on the basis of the 2014 and 2015 PAN competitions.
@@
 However, the problem of using features that can potentially capture topic or content instead of authorship remains in many verification systems, including more recent ones.
-One of the first approaches dealing with content bias in the context of authorship analysis was the topic-masking technique TextDistortion [103], which is simply based on replacing any word that is not among the top most frequent words in a reference corpus with an asterisk.
+One of the first approaches dealing with content bias in the context of authorship analysis was the topic-masking technique TextDistortion [119], which is simply based on replacing any word [^2] As an alternative, Stamatatos [119] proposed an additional variant of replacing not the entire word $w$ but instead each character in it with an asterisk.
+$w$ that is not among the top most frequent words in a reference corpus with an asterisk.
 TextDistortion was found to actually improve the results in cross-topic and cross-genre analyses.
-In 2021, Halvani and Graner [42] tested TextDistortion against another algorithm they introduced called POSNoise, which is a pre-processing method that consists in the replacement of words that are unlikely to be functional with their Part-Of-Speech (POS).
+In 2021, Halvani and Graner [53] tested TextDistortion against another algorithm they introduced called POSNoise (see Section 6.2), which is a pre-processing method that consists in the replacement of words that are unlikely to be functional with their Part-Of-Speech (POS).
 They found that this method clearly improves the algorithms in cross-domain cases while often resulting in poorer performance in cases that are not cross-domain.
@@
 Among the methods that benefited from this pre-processing was IM, which was also one of the most successful ones in their evaluations.
-A similar feature type to avoid content bias is the use of frame $n$-grams [82].
-Similarly to POSNoise, this feature set consists of all function words and with all nouns, verbs, adjectives, adverbs, and pronouns being replaced with their corresponding POS.
 As an alternative to these rule-based content removal methods, Bischoff et al.
@@
 Despite the success of their own neural network method, they discovered that a simpler approach such as TextDistortion works equally well at a fraction of the computational and data cost of a neural network.
-Another approach to counteract content bias, called TAVeer, was proposed by Halvani et al.
+Another approach to counteract content bias, called TAVeer  [^3] TAVeer stands for Topic-agnostic Authorship Verifier based on equal error rate”., was proposed by Halvani et al.
 [55].
-In contrast to TextDistortion and POSNoise, TAVeer represents a stand-alone AV method based on ensembles of topic-agnostic (TA) feature categories (namely punctuation $n$-grams, TA sentence and sentence-initial tokens, TA sentence endings, TA token $n$-grams and TA masking token $n$-grams).
+In contrast to TextDistortion and POSNoise, TAVeer represents a stand-alone AV method based on ensembles of topic-agnostic (TA) feature categories (namely punctuation $n$-grams, TA sentence and sentence-initial tokens, TA sentence endings, TA token $n$-grams and TA masked token $n$-grams).
 The TA feature categories themselves cover a total of $\approx 1000$ words and phrases that originate from 20 linguistic categories such as punctuation marks, conjunctions, determiners, prepositions, contractions, auxiliary verbs, transition words and others, which are known in the literature [100, 17, 124] to be content and topic independent.
-1.5 Language Models and Compression Methods
+Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$ and a set of selected feature categories, TAVeer constructs two feature vectors for each category, which represent the $L^{1}$-normed occurrence frequencies of specific features (such as the conjunction word “and”) in the corresponding document.
+Afterwards, for each category, the Manhattan distance between two vectors is calculated and calibrated into the interval $[0,1]$, where $0.5$ represents a decision boundary.
+The median of the calibrated values over all selected categories, compared to $0.5$ then represents the final decision of $\mathcal{A}=\mathcal{U}$ or $\mathcal{A}\neq\mathcal{U}$.
+The thresholds and the optimal subset of selected TA feature categories are optimized, by selecting the ones that yield the best accuracy on a given training corpus.
+1.4.3 Language Models and Compression Methods
 In addition to IM and its variants, the most successful family of methods in AV are the ones that are based on language modeling and/or compression, which are mathematically related.
-A language model is here defined as a generative statistical model that assigns a probability distribution to stretches of language.
+Here, a language model is defined as a generative statistical model that assigns a probability distribution to stretches of language.
 One of the very first successful examples for this class of methods was the winner of the PAN 2015 competition, Bagnall [12], whose AV method was based on a character-level recurrent neural network language model.
-Another notable neural network language modeling approach is Boenninghoff’s et al.
-[14] use of a Siamese neural network.
-Their approach was found to outperform IM by halving its error rate while also offering a way to visualize the importance of tokens.
-Their approach was designed to mitigate the effect of content but its limitation is that a large and structured comparison dataset is necessary, which is not always available.
-King and Cook [61] also proposed an approach based on language models which works by retraining the final layer of a model based on potentially short samples from an author and then comparing this model against a general model on the questioned text.
-They found that this system outperformed the use of traditional machine learning methods like Support Vector Machines.
+Another noteworthy neural language modeling approach, based on a Siamese neural network, is HRSN  [^4] HRSN stands for Hierarchical Recurrent Siamese Network, proposed by Boenninghoff et al.
+[20].
+This approach was found to outperform IM by reducing its error rate by half.
+However, as noted by the authors [19], it has two key drawbacks: first, it does not utilize linguistically interpretable stylometric features, and second, it lacks a mechanism for directly visualizing the decision-making process.
+Additionally, HRSN is constrained by its reliance on a large and well-structured comparison dataset, which may not always be available.
+Another well-known AV method that can be regarded as an improved version of HRSN is ADHOMINEM  [^5] ADHOMINEM stands for Attention-based Deep Hierarchical cOnvolutional siaMese bIdirectional recurreNt nEural-network Model [19]., which was also proposed by Boenninghoff et al.
+[19].
+ADHOMINEM achieved the highest ranking in the PAN AV competition 2020 [70] and, according to Lundmark et al.
+[87], has shown great promise with respect to real AV applications.
+The method can be summarized as follows: Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, ADHOMINEM preprocesses both documents by hierarchically breaking them down into sentences, words, and characters.
+A Siamese neural network, incorporating a CNN and bi-directional LSTM layers with attention mechanisms, is then employed to derive a neural document embedding for each input.
+Specifically, the CNN layer extracts word representations from character embeddings, which are subsequently passed to an LSTM layer operating at the sentence level.
+The final LSTM layer synthesizes these sentence embeddings to construct a holistic document representation.
+Based on the document embeddings of $\mathcal{D}_{\mathcal{U}}$ and $\mathcal{D}_{\mathcal{A}}$, a nonlinear metric learning scheme is employed, which measures how closely $\mathcal{D}_{\mathcal{U}}$ and $\mathcal{D}_{\mathcal{A}}$ resemble each other.
 Although language models are also inevitably affected by the issue of content bias, attempts to mitigate this problem are scarce.
@@
 Both of them can be used together with POSNoise to avoid content bias.
-Although IM tended to be more robust over the various tests it was subjected to, the advantage of COAV over IM is that the method is deterministic and much faster (for comparison see, for example, [46, Table 4]).
+Although IM tended to be more robust over the various tests it was subjected to, the advantage of COAV over IM is that the method is deterministic and much faster (for comparison see, for example, [57, Table 4]).
+Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, COAV compresses $\mathcal{D}_{\mathcal{U}}$, $\mathcal{D}_{\mathcal{A}}$ and the concatenation of both using a PPM variant called PPMd.
+Based on the length of the three compression results, a dissimilarity value is calculate using Compression-based Cosine (CBC).
+This dissimilarity is then compared to a threshold that is optimized on a training corpus to estimate whether either $\mathcal{A}=\mathcal{U}$ or $\mathcal{A}\neq\mathcal{U}$ holds.
 In addition, Halvani [51] was able to visualize the importance of character sequences for the prediction by producing a color-coded version of the texts, thus making the method more qualitatively explorable by the analyst.
+1.4.4 Feature Vector Difference
+Finally, another AV method that is also worth mentioning, and which was ranked second in the PAN AV competition 2020 [70], is the so-called Feature Vector Difference approach proposed by Weerasinghe and Greenstadt [132].
+Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, the first step of this method is to construct corresponding feature vectors $\mathcal{F}\mbox{}_{\mathcal{U}},\mathcal{F}\mbox{}_{\mathcal{A}}$ using a set of features.
+Then, based on these two, a difference vector $\mathcal{F}\mbox{}=|\mathcal{F}\mbox{}_{\mathcal{U}}-\mathcal{F}\mbox{}_{\mathcal{A}}|$ is then calculated, which represents either a Y- or N-case.
+This procedure is performed for all verification cases in a training corpus, and based on these a classifier is finally trained.
+For this, Weerasinghe and Greenstadt considered two classification models: Logistic Regression (LR) and Multi-Layer Perceptron (MLP).
+Regarding the features, the authors chose tf-idf values of: character $n$-grams (for $n\in\{1,2,\ldots,6\}$), POS tag trigrams, 31 predefined special characters, trigrams of POS tag chunks (extracted from the parse tree representation of the document) as well as noun and verb phrase expansions.
+Moreover, they considered frequencies of 179 stopwords (defined in the NLTK corpus package [^6] https://github.com/nltk/nltk), total number of characters and tokens in the document, average word length (characters per word) per document, distribution of token lengths (relative frequency of tokens with length $\ell$ (for $1\leq\ell\leq 10$) and vocabulary richness measures.
+2 Cognitive Linguistics, Grammar, and Linguistic Individuality
+After reviewing the AV literature, this section introduces the theoretical grounding for the new method that we are proposing, which is motivated and inspired by the field of Cognitive Linguistics.
+In Cognitive Linguistics, language and grammar are seen as complex adaptive systems [13], where regularities emerge from speakers interacting with each other, like a phenomenon of the third kind [40].
+This is partially evidenced by decades of linguistics studies on corpora which have revealed that the traditional understanding of grammar as a set of phrase structure rules acting over basic Part-of-Speech categories (e.
+g., nouns, adjectives) is untenable and that, instead, a better model of grammar is a lexicogrammar continuum that goes from fixed lexicalized expressions (e.
+g., “I don’t know”) to flexible templates (e.
+g., “the” X-er “the” Y-er) up to completely abstract structures (e.
+g., Agent Action Patient).
+This continuum roughly corresponds to the continuum between declarative memory, which is semantic/encyclopedic memory for facts and which tends to be conscious, and procedural memory, which is instead memory for the probabilistic prediction of patterns of sequences and that deals with subconscious knowledge of habits and skills [128].
+The roles that these two memories play in language is clearly evidenced in cases of aphasia.
+When an impairment affects circuits of declarative memory, this gives rise to fluent aphasia, resulting in speech that is mostly constituted by function words (articles, prepositions, pronouns) or morphological forms (“-ed”, “-s”) but with lack of content words that contain meaning, leading to fluent but nonsensical speech, almost resembling POSNoise pre-processed data such as the one in Table 7.
+Instead, when the impairment is on procedural memory, then speech is only consisting of unconnected content words [129].
+These two memory systems work together and, typically, once a habit is formed through repetition then procedural memory takes over, leading to fluency of expression.
+This transition explains why languages tend to be heavily formulaic.
+For example, in English one can say “big mistake” but not “large mistake”, even though a formal phrase structure rule based grammar would seemingly have no problems with the latter expression, which is an adjective modifying a noun.
+What explains this reliance on formulaic expressions is the cognitive tendency to rely as much as possible on habitual and ritualized patterns stored as templates [128, 115].
+The way these units of procedural memory are formed is through the process of chunking: when two items that often go together are found together, then these two items can be treated as one, thus saving space in working memory [48, 27, 34].
+In all effects, chunking is a process of information compression and, if the units of grammar are chunks, then grammar itself can be thought of as a compressed version of a language or, equivalently, as a probability model over these units.
+According to Cognitive Linguistics, therefore, a probability model like a language model is a better model of grammar than a list of phrase structure rules.
+This way of conceiving language also explains why each individual has a different and possibly unique grammar, a statement encapsulated by the Principle of Linguistic Individuality [97].
+The status of unithood given by chunking is not discrete (e.
+g., something is or is not a unit) but a continuum of strength of unithood, where this strength is called entrenchment [39, 111].
+Because entrenchment is dependent on regular exposure and usage, it is fundamentally personal and idiosyncratic.
+It is for this reason that Langacker hypothesized that “the set of units mastered by all members of a speech community might be a rather small proportion of the units constituting the linguistic ability of any given speaker” [85], a hypothesis that is finding increasing empirical support [97].
+For example, for most if not all speakers of English the entrenchment for a sequence such as “of the” is maximal, meaning that they can produce it effortlessly and unconsciously as a single unit.
+In contrast, a sequence such as “the of” would not be a unit for speakers of English because it is virtually never encountered or used.
+If a speaker of English encounters this sequence, they would decompose it as a structure made up of two units, “the” and “of”.
+In between these two extreme examples there is a lot of freedom as to what counts as a unit for a person or another, depending on what each individual has automatized in their procedural memory.
+For example, an individual $\mathcal{A}_{1}$ might have a very entrenched sequence such as “it is known that” when writing academic articles while an individual $\mathcal{A}_{2}$ might have “it is well established that”.
+The higher the entrenchment, the more likely it is that $\mathcal{A}$ is using procedural memory and that it is therefore performing an unconscious routine that constitutes a unit for this individual.
+These sequences will then have open slots ready to be filled in with the appropriate meaning.
+Although there have been other attempts at modeling the grammar of an author in the past for AV or AA purposes [31, 11, 46, 50], these methods relied too much on a conception of grammar in the traditional sense, for example using sequences of POS tags or syntactic trees.
+For the reasons outlined above, these ways of modeling grammar are not compatible with predictions made in Cognitive Science and Cognitive Linguistics.
+The picture of grammar painted by Cognitive Linguistics is much more lexically-based and word-idiosyncratic.
+For example, the preposition “of” in English behaves very oddly compared to other prepositions [115].
+Giving it the same POS tag as to other prepositions is therefore going to obscure a lot of important grammatical information.
+With this theoretical grounding, it is now possible to present our proposed novel approach, LambdaG.
 3 The Likelihood Ratio of Grammar Models
-The method proposed in this paper is a binary-extrinsic AV method influenced by the literature above which offers a general theoretically-motivated framework for AV that can help to better explain the other methods reviewed above.
-This section explains its theoretical framework and the algorithm, which is outlined in Algorithm 1, while most technical details are left to Section 5.
-Input: $\mathbb{D}_{\mathcal{U}}=\{\mathcal{D}_{\mathcal{U}}^{1},\mathcal{D}_{\mathcal{U}}^{2},\dots\}$, $\mathbb{D}_{\mathcal{A}}=\{\mathcal{D}_{\mathcal{A}}^{1},\mathcal{D}_{\mathcal{A}}^{2},\dots\}$, $\mathbb{D}_{\text{ref}}=\{\mathcal{D}_{\text{ref}}^{1},\mathcal{D}_{\text{ref}}^{2},\dots\}$, $N$, $r$
-1 Function sample($k$, $\mathbb{S}$):
-2       Randomly sample a subset of $k$ sentences from the input set $\mathbb{S}$
-3Function sent($\mathbb{D}$):
-4       Construct a set $\mathbb{S}$ of tokenized sentences from the input set $\mathbb{D}$
-// Apply POSNoise to each document set
-$\mathbb{S}_{\mathcal{U}}\mbox{}\leftarrow\textnormal{{sent}}(\textsf{POSNoise}\mbox{}(\mathbb{D}_{\mathcal{U}}))$ $\mathbb{S}_{\mathcal{A}}\mbox{}\leftarrow\textnormal{{sent}}(\textsf{POSNoise}\mbox{}(\mathbb{D}_{\mathcal{A}}))$ $\mathbb{S}_{\textsf{ref}}\mbox{}\leftarrow\textnormal{{sent}}(\textsf{POSNoise}\mbox{}(\mathbb{D}_{\text{ref}}))$ // Build Grammar Model for the known author $\mathcal{A}$
-$G_{\mathcal{A}}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\mathcal{A}},N)$ // Build reference Grammar Models
-5 for $i\leftarrow 1$ to $r$ do
-6       $\mathbb{S}_{\text{i}}\leftarrow$ sample($|\mathbb{S}_{\mathcal{A}}|$, $\mathbb{S}_{\text{ref}}$) $G_{i}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\text{i}},N)$
+The algorithm presented in this paper is a binary-extrinsic AV method.
+This section provides an overview of the algorithm, which is formally outlined in Algorithm 1 and schematically illustrated in Figure 1.
+For a more detailed technical discussion, refer to Section 6.
+Figure 1: Schematic overview of our proposed AV method LambdaG.
+Input: $\mathcal{D}_{\mathcal{U}}$, $\mathcal{D}_{\mathcal{A}}$, $\mathbb{D}_{\text{ref}}=\{\mathcal{D}_{\text{ref}}^{1},\mathcal{D}_{\text{ref}}^{2},\dots\}$, $N$, $r$
+// where $\mathcal{D}_{\mathcal{U}}$, $\mathcal{D}_{\mathcal{A}}$ are two documents, $\mathbb{D}_{\text{ref}}=\{\mathcal{D}_{\text{ref}}^{1},\mathcal{D}_{\text{ref}}^{2},\dots\}$ is a set of reference documents, $N$ is the model order and $r$ is the number of repetitions
+1
+2Function sample($s$, $\mathbb{S}$):
+3       Randomly select a subset of $s$ sentences from the given tokenized input sentence set $\mathbb{S}$
+4
+5
+6Function posnoise($\mathbb{D}$):
+7       Applies POSNoise to each document in $\mathbb{D}$ and returns a set of their posnoise representations
+8
+9Function sent($\mathbb{D}$):
+10       Construct a set $\mathbb{S}$ of tokenized sentences from all documents in $\mathbb{D}$
+11
+// Apply POSNoise to all respective documents and construct tokenized sentences
+12 $\mathbb{S}_{\mathcal{U}}\mbox{}\leftarrow\textnormal{{sent}}(\textnormal{{posnoise}}(\{\mathcal{D}_{\mathcal{U}}\}))$
+13 $\mathbb{S}_{\mathcal{A}}\mbox{}\leftarrow\textnormal{{sent}}(\textnormal{{posnoise}}(\{\mathcal{D}_{\mathcal{A}}\}))$
+14 $\mathbb{S}_{\textsf{ref}}\mbox{}\leftarrow\textnormal{{sent}}(\textnormal{{posnoise}}(\mathbb{D}_{\text{ref}}))$
+15
+// Build Grammar Model for the known author $\mathcal{A}$
+16 $G_{\mathcal{A}}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\mathcal{A}},N)$
+17
+// Build reference Grammar Models
+18 for $j\leftarrow 1$ to $r$ do
+19       $\mathbb{S}_{\text{j}}\leftarrow$ sample($|\mathbb{S}_{\mathcal{A}}|$, $\mathbb{S}_{\text{ref}}$)
+20       $G_{j}\leftarrow\textnormal{{KN}}(\mathbb{S}_{\text{j}},N)$
+21
+22
 // Calculate $\lambda_{G}$ (the log-likelihood ratio of the Grammar Models) over $\mathbb{S}_{\mathcal{U}}\mbox{}$
-7 $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow 0$ for $S\mbox{}\in\mathbb{S}_{\mathcal{U}}$ do
-       // Decompose $S\mbox{}$ into a sequence of tokens
-8       $(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{m})\leftarrow S\mbox{}$ for $j\leftarrow 1$ to $m$ do
-             // Calculate mean $\lambda_{G}$ for $t\mbox{}_{j}$ over reference Grammar Models
-9             for $i\leftarrow 1$ to $r$ do
-10                   $c\leftarrow(t\mbox{}_{j-N+1},\ldots,t\mbox{}_{j-1})$ $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})+\frac{1}{r}\log\frac{P(t\mbox{}_{j}|c;G_{\mathcal{A}})}{P(t\mbox{}_{j}|c;G_{i})}$
-11
-12
-return $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})$
+23 $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow 0$
+24
+25for $S\mbox{}_{i}\in\mathbb{S}_{\mathcal{U}}$ do
+       // Decompose $S\mbox{}_{i}$ into a sequence of tokens
+26       $(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{z})\leftarrow S\mbox{}_{i}$
+27       for $k\leftarrow 1$ to $z$ do
+             // Calculate mean $\lambda_{G}$ for $t\mbox{}_{k}$ over reference Grammar Models
+28             for $j\leftarrow 1$ to $r$ do
+29                   $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})\leftarrow\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})+\dfrac{1}{r}\log\dfrac{P(t\mbox{}_{k}|t_{<k};G_{\mathcal{A}})}{P(t\mbox{}_{k}|t_{<k};G_{j})}$
+30
+31
+32
+33 return $\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})$
 Algorithm 1 Likelihood Ratio of Grammar Models (LambdaG)
-[…]
-$P(S\mbox{};G)=\prod_{j=1}^{m}P(t\mbox{}_{j}|t\mbox{}_{j-N+1},t\mbox{}_{j-N+2},\ldots,t\mbox{}_{j-1};G)$ (1)
-For readability reasons, we rewrite the probability assigned by $G$ to each token $t\mbox{}$ in its context $c$ as $P(t\mbox{}|c;G)$ (as in line 18 in Algorithm 1, where $t\mbox{}_{j}$ is used instead of $t$).
+We define the most basic grammatical unit of analysis as a function token $t\mbox{}\in\mathcal{T}\mbox{}_{L}$, where $\mathcal{T}\mbox{}_{L}$ is the set all function tokens of a language $L$, including all function words, punctuation marks as well as abstract grammatical categories.
+For example, for English, $\mathcal{T}\mbox{}_{E}=\{\texttt{the},\texttt{of},\texttt{at},\dots,\allowbreak\texttt{NOUN},\allowbreak\texttt{VERB},\allowbreak\texttt{ADJECTIVE},\dots,\texttt{!},\texttt{:},\texttt{;}\dots\}$.
+We then define a Grammar Model $G$ as a statistical model that generates a probability distribution over sentences $\mathbb{S}$, which are defined as complete sequences of function tokens.
+To obtain these representations, we first apply the POSNoise algorithm (see Section 6.2) to $\mathcal{D}_{\mathcal{U}}$, $\mathcal{D}_{\mathcal{A}}$, and each $\mathcal{D}$ in the reference corpus $\mathbb{D}_{\text{ref}}$, retaining only function tokens.
+This is indicated in Algorithm 1, lines 7–9, where POSNoise is applied.
+Then, we tokenize each $\mathcal{D}$ into sentences by adding sentence boundaries markers at an end of sentence punctuation mark or new line, thus turning them into their respective set of sentences $\mathbb{S}_{\mathcal{A}}\mbox{}$, $\mathbb{S}_{\mathcal{U}}\mbox{}$ and $\mathbb{S}_{\textsf{ref}}\mbox{}$ (also lines 7–9 in Algorithm 1).
+The boundaries of each sentence could also be set by a grammatical parsing algorithm.
+However, we observed that LambdaG is robust to this variation.
+We then employ $n$-gram models, which allow the probability of any grammatical token to depend only on a finite number of preceding tokens from the same sentence, thereby capturing only the information contained in short-term grammatical correlations.
+$n$-gram models assume that each sentence is independent and we maintain that this is not an unreasonable assumption since grammatical dependencies between tokens are known to be relatively local [34].
+The probability assigned to each sentence $S\mbox{}_{i}=(t\mbox{}_{1},t\mbox{}_{2},\ldots,t\mbox{}_{z})$ by a Grammar Model $G$ would then be calculated as follows:
+$P(S\mbox{}_{i};G)=\prod_{k=1}^{z}P(t\mbox{}_{k}|\underbrace{t\mbox{}_{k-N+1},t\mbox{}_{k-N+2},\ldots,t\mbox{}_{k-1}}_{t\mbox{}_{<k}};G)$ (1)
+For readability reasons, we rewrite the probability assigned by $G$ to each token $t\mbox{}_{k}$ in its context as $P(t\mbox{}_{k}|t\mbox{}_{<k};G)$ (as in line 19 in Algorithm 1).
 We estimate Grammar Models adopting Kneser-Ney smoothing [76, 32], represented in Algorithm 1 as $\textnormal{{KN}}()$, with a fixed single default discount parameter $D=0.75$, as explained in Section 6.3.
-We set $N=10$ as an order of the model to ensure that all grammatical relations among the tokens of a sentence are captured, even in cases of long sentences.
-However, the order of the model $N$ could be considered as a hyperparameter of the algorithm (lines 8 and 11 in Algorithm 1).
+We set $N=10$ as an order of the model to ensure that most grammatical relations among the tokens of a sentence are captured, even in cases of long sentences.
+However, the order of the model $N$ could be considered as a hyperparameter of the algorithm (lines 10 and 13 in Algorithm 1).
 Assuming the Principle of Linguistic Individuality that at any time $\tau$, for any language $L_{\tau}$, there do not exist two individuals, $\mathcal{A}$ and $\mathcal{B}$, for whom their grammar is identical [97], we expect that in a large number of cases the two Grammar Models $G_{\mathcal{A}}$ and $G_{\mathcal{B}}$ would assign two different probabilities to the same token in context.
-$P(t\mbox{}|c;G_{\mathcal{A}})$ can be interpreted as the extent to which token $t\mbox{}$ is ’grammatical’ or, according to Cognitive Linguistic theories, entrenched for author $\mathcal{A}$ when in context $c$.
-For this reason, we would expect that, if the sequence $s=t\mbox{}_{1},t\mbox{}_{2},t\mbox{}_{3}$ was produced by $G_{A}$, then $P(s;G_{\mathcal{A}})>P(s;G_{\mathcal{B}})$ or, more generally according to the Principle of Linguistic Individuality, $P(s;G_{\mathcal{A}})>P(s;G_{\neg\mathcal{A}})$.
-The model $G_{\mathcal{A}}$ is estimated from $\mathbb{S}_{\mathcal{A}}\mbox{}$ (line 8 in Algorithm 1).
-Instead, to generate an approximation to $G_{\neg\mathcal{A}}$ we estimate a set of reference Grammar Models $\mathbb{S}_{\textsf{ref}}\mbox{}=\{G_{1},G_{2},\dots,G_{r}\}$ where each $G_{r}$ is generated by uniformly sampling randomly from $\mathbb{S}_{\textsf{ref}}\mbox{}$ a set of sentences $\mathbb{S}_{r}$ of the same size as the set of sentences in the candidate author’s corpus, $\mathbb{S}_{\mathcal{A}}\mbox{}$, so that $|\mathbb{S}_{r}|=|\mathbb{S}_{\mathcal{A}}\mbox{}|$ (line 10 in Algorithm 1).
+$P(t\mbox{}_{k}|t\mbox{}_{<k};G_{\mathcal{A}})$ can be interpreted as the extent to which token $t\mbox{}_{k}$ is ’grammatical’ or, according to Cognitive Linguistic theories, entrenched for author $\mathcal{A}$ when in context $t\mbox{}_{<k}$.
+For this reason, we would expect that, if the sequence $\sigma=(t\mbox{}_{1},t\mbox{}_{2},t\mbox{}_{3})$ was produced by $G_{A}$, then $P(\sigma;G_{\mathcal{A}})>P(\sigma;G_{\mathcal{B}})$ or, more generally according to the Principle of Linguistic Individuality, for any $\mathcal{A}^{\prime}\neq\mathcal{A}$, $P(\sigma;G_{\mathcal{A}})>P(\sigma;G_{\mathcal{A}^{\prime}})$.
+The model $G_{\mathcal{A}}$ is estimated from $\mathbb{S}_{\mathcal{A}}\mbox{}$ (line 10 in Algorithm 1).
+To approximate a comparison population of $G_{\mathcal{A}^{\prime}}$, we estimate a set of reference Grammar Models $\mathbb{S}_{\textsf{ref}}\mbox{}=\{G_{1},G_{2},\dots,G_{r}\}$ where each $G_{r}$ is generated by uniformly sampling randomly from $\mathbb{S}_{\textsf{ref}}\mbox{}$ a set of sentences $\mathbb{S}_{r}$ of the same size as the set of sentences in the candidate author’s corpus, $\mathbb{S}_{\mathcal{A}}\mbox{}$, so that $|\mathbb{S}_{\text{ref}}|=|\mathbb{S}_{\mathcal{A}}\mbox{}|$ (line 12 in Algorithm 1).
 We set $r=100$ but this value can also be considered a hyperparameter of the algorithm.
 A comparison of hyperparameter settings is reported in Section 4.
-Taking an approach consistent with the likelihood ratio framework adopted in forensic science, similarly to Ishihara [48], it is therefore possible to calculate (lines 16-18 in Algorithm 1) a mean log-likelihood ratio of the probability assigned to each token in context by $G_{\mathcal{A}}$ and by each $G_{r}$, which we call $\lambda_{G}$:
-$\lambda_{G}(t\mbox{}|c)=\frac{1}{r}\sum_{i=1}^{r}{\log\frac{P(t\mbox{}|c;G_{\mathcal{A}})}{P(t\mbox{}|c;G_{i})}}$ (2)
+Taking an approach consistent with the likelihood ratio framework adopted in forensic science, similarly to Ishihara [61], it is therefore possible to calculate (lines 17–19 in Algorithm 1) a mean ratio of the likelihoods assigned to each token in context by $G_{\mathcal{A}}$ and by each $G_{r}$, which we call $\lambda_{G}$:
+$\lambda_{G}(t\mbox{}_{k}|t\mbox{}_{<k})=\frac{1}{r}\sum_{j=1}^{r}{\log\frac{P(t\mbox{}_{k}|t\mbox{}_{<k};G_{\mathcal{A}})}{P(t\mbox{}_{k}|t\mbox{}_{<k};G_{j})}}$ (2)
 In Eq.
-2 the numerator can be seen as a measure of similarity, how much the token is grammatical or entrenched for $G_{\mathcal{A}}$ given its context, while the denominator is a measure of typicality, how ’ungrammatical’ or atypical this token in context is in the reference population.
-A value of $\lambda_{G}$ for each $S\mbox{}\in\mathbb{S}_{\mathcal{U}}$ can then be calculated (line 14-18 in Algorithm 1) as:
-$\lambda_{G}(S\mbox{})=\sum_{j=1}^{m}\lambda_{G}(t\mbox{}_{j}|c)$ (3)
-Since we model each $S\mbox{}\in\mathbb{S}_{\mathcal{U}}\mbox{}$ as independent from each other, we can then calculate (line 12-18 in Algorithm 1):
-$\begin{split}\lambda_{G}(\mathcal{D}_{\mathcal{U}})=\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})=\sum_{S\mbox{}\in\,\mathbb{S}_{\mathcal{U}}\mbox{}}\lambda_{G}(S\mbox{})=\sum\limits_{S\mbox{}\,\in\,\mathbb{S}_{\mathcal{U}}\mbox{}}\sum\limits_{j=1}^{m}{\frac{1}{r}\sum\limits_{i=1}^{r}\log\frac{P(t\mbox{}_{j}|c;G_{\mathcal{A}})}{P(t\mbox{}_{j}|c;G_{i})}}\end{split}$ (4)
+2 the numerator can be seen as a measure of similarity, how much the token is grammatical or entrenched for $G_{\mathcal{A}}$ given its context, while the denominator is a measure of typicality, how atypical this token is in this context in the reference population.
+A value of $\lambda_{G}$ for each $S\mbox{}_{i}\in\mathbb{S}_{\mathcal{U}}\mbox{}=\{S\mbox{}_{1},S\mbox{}_{2},\dots,S\mbox{}_{m}\}$, can then be calculated (line 15–19 in Algorithm 1) as:
+$\lambda_{G}(S\mbox{}_{i})=\sum_{k=1}^{z}\lambda_{G}(t\mbox{}_{k}|t\mbox{}_{<k})$ (3)
+Since we model each $S\mbox{}_{i}\in\mathbb{S}_{\mathcal{U}}\mbox{}$ as independent from each other, we can then calculate (line 14–20 in Algorithm 1):
+$\begin{split}\lambda_{G}(\mathbb{S}_{\mathcal{U}}\mbox{})=\sum_{i=1}^{m}\lambda_{G}(S\mbox{}_{i})=\sum\limits_{i=1}^{m}{\sum\limits_{k=1}^{z}\lambda_{G}(t\mbox{}_{k}|t\mbox{}_{<k})}\end{split}$ (4)
 An advantage of this formula is that the final $\lambda_{G}$ can be decomposed into $\lambda_{G}$ values for each single sentence and each single token in each single sentence.
-This means that the analyst can reconstruct step by step what features contribute to the final score, for example, by ranking each sentence for importance and, for each sentence, color-code or highlight the most important tokens.
+This means that the analyst can reconstruct step by step what features contribute to the final score, for example, by ranking each sentence for importance and, for each sentence, color-code or highlight the most important tokens, as demonstrated in Section 7.
 In addition to this property, the $\lambda_{G}$ value is also effectively an uncalibrated log-likelihood ratio score that includes information about both similarity and typicality and that can then be calibrated into a forensic log-likelihood ratio using a calibration method such as a logistic regression.
@@
 More details on the Likelihood Ratio Framework for forensic science and its evaluation is described in Section 6.4.
-Based on $\Lambda_{G}$, we classify a verification case as Y if $\Lambda_{G}>0$, else as N.
-Finally, we note that another important aspect that is not discussed in detail in this section is the fact that several of the key components of LambdaG are consistent with Cognitive Linguistics theories.
-More details about this aspect of LambdaG are left to Section 4.
+Based on $\Lambda_{G}$, we classify a new verification case $c_{\textrm{new}}$ as Y if $\Lambda_{G}>0$ (that is, if $\text{sigmoid}(\Lambda_{G})>0.5$), else as N.
 4 Results
 Algorithm 1 was applied to a set of twelve corpora that simulate various real-life AV problems and scenarios, described in Section 6.
-The full results are presented in Table 1, where the performance of each method is evaluated using five standard performance measures for classification: Accuracy, the Area Under the Curve (AUC), F1, Precision, and Recall.
+The complete results are presented in Table LABEL:tab:EvalResults, where the performance of each method is evaluated using five standard performance measures for classification: Accuracy, the Area Under the Curve (AUC), $F_{\text{1}}$, Precision, and Recall.
 These are values that were calculated using a threshold $\theta$, which is derived from $\Lambda_{G}$.
 The split between training and test datasets for each corpus tested is shown in Table 5.
-To assess the performance of $\Lambda_{G}$, instead, the $C_{\text{llr}}\mbox{}$ measure is used, which expresses the amount of information in the likelihood ratio.
+To further assess the performance of $\Lambda_{G}$, the $C_{\text{llr}}\mbox{}$ measure is also used, which is a typical measure of accuracy for a likelihood ratio.
 This metric is explained in Section 6.4.
-In Table 1, LambdaG is compared against four well-known and tested AV methods: IM, COAV, TAVeer and a neural language model using a Siamese network, SiamBERT, which is explained in Section 5.5.
-IM and COAV were run on POSNoise pre-processed corpora and they are therefore content-agnostic, similarly to LambdaG.
-TAVeer was run on the unprocessed corpora but because the method only measures topic-agnostic features, this is also comparable to LambdaG.
-Since SiamBERT uses a pre-trained model on data that it is not POSNoise-processed, fine-tuning it on POSNoise-processed data would be an unfair comparison.
-Equally, however, SiamBERT fine-tuned on unprocessed data risks using content information and is also, in a different way, an unfair comparison.
-For this reason, we present the results of both versions.
-The results of SiamBERT on POSNoise-processed data is indicated with ${}^{\,(\bm{\dagger})}$.
-The results in Table 1 show that LambdaG is superior to the other AV topic-agnostic baselines in all cases.
-LambdaG is also the best performing method when compared to SiamBERT despite the fact that any correlation between author and topic could have been exploited in each corpora.
-The only exception is for $\mathcal{C}_{\mathrm{AllNews}}$, where instead LambdaG comes second to SiamBERT.
-We do not have an explanation for this result but we advance the plausible hypothesis that $\mathcal{C}_{\mathrm{AllNews}}$ contains data with a significant author/topic bias such that the task is made easier when content words are included.
-This hypothesis is supported by the extremely poor performance of ${}^{\,(\bm{\dagger})}$ on $\mathcal{C}_{\mathrm{Stack}}$, which is a corpus explicitly designed to penalize AV methods that exploit topic information.
-Table 1: Evaluation results of LambdaG and the four baseline methods.
-${}^{\,(\bm{\dagger})}$ indicates that the method was applied to the original (i.
+When running the analysis on the training corpora, the reference corpus $\mathbb{D}_{\text{ref}}$ was the remaining of the corpus after taking out the authors of $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$.
+When analyzing the test corpora, the reference corpus was simply the corresponding training corpus.
+A robustness test using different corpora as reference data was also carried out (see Figure 4).
+Given that LambdaG is non-deterministic, each analysis was repeated five times, and the results presented in Table LABEL:tab:EvalResults are the average of these five repetitions.
+In Table LABEL:tab:EvalResults, LambdaG is compared with seven well-known and tested baseline AV methods: IM, COAV, TAVeer, ADHOMINEM, FeVecDiff (LR), FeVecDiff (MLP), and our own developed baseline SiamBERT, which is explained in Section 6.6.
+IM, COAV and FeVecDiff (LR and MLP) were run on POSNoise pre-processed corpora and are therefore content-agnostic, similarly to LambdaG.
+TAVeer, on the other hand, was applied to the unprocessed corpora.
+However, since this method is inherently designed to consider only topic-agnostic features [^7] Note that both TAVeer and POSNoise rely on the same list of topic-agnostic words and phrases shown in Table 6., its results remain comparable to those of LambdaG.
+With regard to SiamBERT and ADHOMINEM, we wish to point out that both methods are based on pre-trained neural network models, which were originally trained on POSNoise-unprocessed text data.
+Using pre-trained models is essential because training neural networks from scratch requires datasets that are larger than our AV corpora.
+Instead, we therefore fine-tuned the original pre-trained models on our AV corpora.
+Since pre-training was performed on POSNoise unprocessed text, the standard approach would be to fine-tune on similarly unprocessed AV corpora (denoted as ${}^{\,(\bm{\dagger})}$ and ${}^{\,(\bm{\dagger})}$ in Table LABEL:tab:EvalResults).
+However, this could introduce topic-related distortions, making the results less comparable with the other AV methods used.
+To mitigate this, we also fine-tuned and applied SiamBERT and ADHOMINEM on our POSNoise pre-processed corpora.
+Table 1: Evaluation results of LambdaG and the seven introduced baselines.
+$(\bm{\dagger})$ is marked next to a pre-trained neural network approach to indicate that it was fine-tuned using the original (i.
 e., non-POSNoise-processed) corpora.
 In terms of Accuracy and AUC, bold and underlined values represent the best and second-best results, respectively.
+For better readability, the confusion matrix outputs for each tested method have been grayed out.
 Method Acc. AUC F1 Prec. Rec. TP FN FP TN
@@
 TAVeer 0.729 0.842 0.735 0.750 0.720 36 12 14 34
+FeVecDiff (LR) 0.781 0.855 0.784 0.776 0.792 38 10 11 37
+FeVecDiff (MLP) 0.750 0.839 0.765 0.722 0.812 39 9 15 33
 SiamBERT 0.646 0.673 0.721 0.595 0.917 44 4 30 18
 $\bm{\mathcal{C}_{\mathrm{Enron}}}$ ${}^{\,(\bm{\dagger})}$ 0.750 0.841 0.778 0.700 0.875 42 6 18 30
+ADHOMINEM 0.583 0.594 0.474 0.643 0.375 18 30 10 38
+${}^{\,(\bm{\dagger})}$ 0.573 0.607 0.438 0.640 0.333 16 32 9 39
 LambdaG 0.876 0.938 0.872 0.841 0.905 95 18 10 103
@@
 TAVeer 0.788 0.858 0.789 0.796 0.783 90 23 25 88
+FeVecDiff (LR) 0.739 0.840 0.678 0.886 0.549 62 51 8 105
+FeVecDiff (MLP) 0.735 0.809 0.737 0.730 0.743 84 29 31 82
 SiamBERT 0.580 0.627 0.676 0.550 0.876 99 14 81 32
 $\bm{\mathcal{C}_{\mathrm{Wiki}}}$ ${}^{\,(\bm{\dagger})}$ 0.757 0.909 0.799 0.681 0.965 109 4 51 62
+ADHOMINEM 0.584 0.641 0.510 0.620 0.434 49 64 30 83
+${}^{\,(\bm{\dagger})}$ 0.575 0.618 0.575 0.575 0.575 65 48 48 65
 LambdaG 0.873 0.933 0.864 0.807 0.929 92 22 7 107
@@
 TAVeer 0.693 0.763 0.685 0.667 0.704 76 38 32 82
+FeVecDiff (LR) 0.939 0.977 0.941 0.910 0.974 111 3 11 103
+FeVecDiff (MLP) 0.882 0.957 0.883 0.872 0.895 102 12 15 99
 SiamBERT 0.605 0.648 0.643 0.587 0.711 81 33 57 57
 $\bm{\mathcal{C}_{\mathrm{Stack}}}$ ${}^{\,(\bm{\dagger})}$ 0.535 0.576 0.579 0.529 0.640 73 41 65 49
-LambdaG 0.782 0.861 0.775 0.75 0.802 105 35 26 114
+ADHOMINEM 0.504 0.516 0.464 0.505 0.430 49 65 48 66
+${}^{\,(\bm{\dagger})}$ 0.596 0.594 0.603 0.593 0.614 70 44 48 66
+LambdaG 0.782 0.861 0.775 0.750 0.802 105 35 26 114
 IM 0.775 0.851 0.755 0.693 0.829 97 43 20 120
@@
 TAVeer 0.707 0.748 0.689 0.650 0.734 91 49 33 107
+FeVecDiff (LR) 0.714 0.752 0.750 0.667 0.857 120 20 60 80
+FeVecDiff (MLP) 0.564 0.632 0.527 0.576 0.486 68 72 50 90
 SiamBERT 0.593 0.641 0.639 0.574 0.721 101 39 75 65
 $\bm{\mathcal{C}_{\mathrm{ACL}}}$ ${}^{\,(\bm{\dagger})}$ 0.607 0.664 0.654 0.584 0.743 104 36 74 66
+ADHOMINEM 0.554 0.575 0.449 0.586 0.364 51 89 36 104
+${}^{\,(\bm{\dagger})}$ 0.496 0.512 0.598 0.498 0.750 105 35 106 34
 LambdaG 0.955 0.990 0.955 0.955 0.955 149 7 7 149
@@
 TAVeer 0.885 0.939 0.889 0.923 0.857 144 12 24 132
+FeVecDiff (LR) 0.878 0.889 0.886 0.835 0.942 147 9 29 127
+FeVecDiff (MLP) 0.865 0.930 0.878 0.803 0.968 151 5 37 119
 SiamBERT 0.865 0.920 0.873 0.828 0.923 144 12 30 126
 $\bm{\mathcal{C}_{\mathrm{PJ}}}$ ${}^{\,(\bm{\dagger})}$ 0.740 0.805 0.771 0.690 0.872 136 20 61 95
+ADHOMINEM 0.673 0.787 0.625 0.733 0.545 85 71 31 125
+${}^{\,(\bm{\dagger})}$ 0.590 0.593 0.642 0.569 0.737 115 41 87 69
+Method Acc. AUC F1 Prec. Rec. TP FN FP TN
 LambdaG 0.909 0.961 0.911 0.929 0.893 158 12 19 151
@@
 TAVeer 0.829 0.891 0.822 0.788 0.859 134 36 22 148
+FeVecDiff (LR) 0.829 0.920 0.839 0.795 0.888 151 19 39 131
+FeVecDiff (MLP) 0.821 0.895 0.828 0.795 0.865 147 23 38 132
 SiamBERT 0.629 0.655 0.652 0.615 0.694 118 52 74 96
 $\bm{\mathcal{C}_{\mathrm{Apric}}}$ ${}^{\,(\bm{\dagger})}$ 0.800 0.871 0.814 0.760 0.876 149 21 47 123
-Method Acc. AUC F1 Prec. Rec. TP FN FP TN
+ADHOMINEM 0.553 0.610 0.510 0.564 0.465 79 91 61 109
+${}^{\,(\bm{\dagger})}$ 0.526 0.592 0.410 0.544 0.329 56 114 47 123
 LambdaG 0.760 0.822 0.758 0.750 0.766 180 60 55 185
@@
 TAVeer 0.690 0.729 0.655 0.588 0.738 141 99 50 190
+FeVecDiff (LR) 0.533 0.534 0.533 0.533 0.533 128 112 112 128
+FeVecDiff (MLP) 0.477 0.475 0.516 0.480 0.558 134 106 145 95
 SiamBERT 0.529 0.551 0.555 0.526 0.588 141 99 127 113
 $\bm{\mathcal{C}_{\mathrm{Trip}}}$ ${}^{\,(\bm{\dagger})}$ 0.581 0.586 0.623 0.567 0.692 166 74 127 113
+ADHOMINEM 0.494 0.497 0.475 0.493 0.458 110 130 113 127
+${}^{\,(\bm{\dagger})}$ 0.502 0.525 0.505 0.502 0.508 122 118 121 119
 LambdaG 0.765 0.839 0.769 0.783 0.755 188 52 61 179
@@
 TAVeer 0.671 0.749 0.686 0.721 0.655 173 67 91 149
+FeVecDiff (LR) 0.508 0.514 0.561 0.507 0.629 151 89 147 93
+FeVecDiff (MLP) 0.508 0.507 0.543 0.507 0.583 140 100 136 104
 SiamBERT 0.571 0.594 0.611 0.559 0.675 162 78 128 112
 $\bm{\mathcal{C}_{\mathrm{Yelp}}}$ ${}^{\,(\bm{\dagger})}$ 0.658 0.727 0.703 0.622 0.808 194 46 118 122
+ADHOMINEM 0.537 0.568 0.456 0.554 0.388 93 147 75 165
+${}^{\,(\bm{\dagger})}$ 0.521 0.588 0.498 0.523 0.475 114 126 104 136
 LambdaG 0.796 0.874 0.791 0.772 0.811 618 182 144 656
@@
 TAVeer 0.691 0.752 0.668 0.621 0.722 497 303 191 609
+FeVecDiff (LR) 0.618 0.644 0.639 0.606 0.675 540 260 351 449
+FeVecDiff (MLP) 0.563 0.593 0.566 0.562 0.569 455 345 354 446
 SiamBERT 0.621 0.672 0.652 0.602 0.711 569 231 376 424
 $\bm{\mathcal{C}_{\mathrm{IMDB}}}$ ${}^{\,(\bm{\dagger})}$ 0.739 0.817 0.755 0.712 0.804 643 157 260 540
-LambdaG 0.891 0.954 0.891 0.892 0.89 803 97 99 801
+ADHOMINEM 0.541 0.553 0.525 0.544 0.507 406 394 341 459
+${}^{\,(\bm{\dagger})}$ 0.581 0.638 0.552 0.594 0.515 412 388 282 518
+LambdaG 0.891 0.954 0.891 0.892 0.890 803 97 99 801
 IM 0.827 0.908 0.812 0.750 0.886 675 225 87 813
@@
 TAVeer 0.781 0.867 0.778 0.767 0.789 690 210 185 715
+FeVecDiff (LR) 0.839 0.921 0.832 0.873 0.794 715 185 104 796
+FeVecDiff (MLP) 0.810 0.882 0.811 0.805 0.818 736 164 178 722
 SiamBERT 0.749 0.843 0.770 0.711 0.840 756 144 308 592
 $\bm{\mathcal{C}_{\mathrm{Blog}}}$ ${}^{\,(\bm{\dagger})}$ 0.701 0.776 0.719 0.679 0.763 687 213 325 575
+ADHOMINEM 0.601 0.647 0.550 0.630 0.488 439 461 258 642
+${}^{\,(\bm{\dagger})}$ 0.647 0.718 0.612 0.680 0.557 501 399 236 664
+Method Acc. AUC F1 Prec. Rec. TP FN FP TN
 LambdaG 0.921 0.977 0.921 0.922 0.919 1107 93 97 1103
@@
 TAVeer 0.842 0.915 0.839 0.822 0.857 986 214 164 1036
+FeVecDiff (LR) 0.901 0.968 0.904 0.876 0.934 1121 79 159 1041
+FeVecDiff (MLP) 0.863 0.937 0.864 0.855 0.874 1049 151 178 1022
 SiamBERT 0.856 0.937 0.864 0.821 0.912 1094 106 239 961
 $\bm{\mathcal{C}_{\mathrm{Amazon}}}$ ${}^{\,(\bm{\dagger})}$ 0.864 0.941 0.871 0.830 0.916 1099 101 225 975
+ADHOMINEM 0.586 0.619 0.578 0.590 0.568 681 519 474 726
+${}^{\,(\bm{\dagger})}$ 0.671 0.759 0.622 0.732 0.540 648 552 237 963
 LambdaG 0.705 0.777 0.689 0.652 0.729 868 463 322 1009
@@
 TAVeer 0.635 0.682 0.614 0.581 0.651 773 558 414 917
+FeVecDiff (LR) 0.663 0.729 0.696 0.634 0.772 1028 303 593 738
+FeVecDiff (MLP) 0.634 0.693 0.637 0.632 0.643 856 475 499 832
 SiamBERT 0.603 0.647 0.631 0.590 0.677 901 430 626 705
 $\bm{\mathcal{C}_{\mathrm{AllNews}}}$ ${}^{\,(\bm{\dagger})}$ 0.807 0.886 0.809 0.799 0.820 1092 239 275 1056
+ADHOMINEM 0.506 0.511 0.572 0.505 0.660 879 452 862 469
+${}^{\,(\bm{\dagger})}$ 0.551 0.549 0.610 0.539 0.702 935 396 799 532
+In addition to Table LABEL:tab:EvalResults, Figure 2 shows 95% confidence intervals for Accuracy, calculated using the classical standard error estimate for a proportion, under the simplifying assumption of independent and identically distributed (i.i.d.) test data:
+$\text{SE}=\sqrt{\frac{1}{n}\Bigl{(}\text{Accuracy}\times(1-\text{Accuracy})}\Bigr{)}$ (5)
+Here, $n$ indicates the total number of tests performed.
+Even though there is no obvious reason to take the i.i.d.
+assumption seriously due to the non-trivial sampling process of our test datasets (see Section 6.1) and the relatively high number of potential comparisons involved, we can consider the resulting standard errors as a rough indication of the potential size of statistical fluctuations and an assessment of whether the observed performance improvements really suggest the presence of actual effects.
+Figure 2: 95% Confidence Intervals for Accuracy.
+When considering the overall pattern of results, we can conclude that LambdaG is superior to the other AV baselines as LambdaG is the only method that is consistently ranked at the top for Accuracy for most corpora.
+This consistency across many corpora representing different AV scenarios in terms of lengths and genres is a strong suggestion that this effect is genuine.
+When judging the robustness of this difference by looking at the overlap of the confidence intervals, LambdaG’s superiority is clear for $\mathcal{C}_{\mathrm{Yelp}}$, $\mathcal{C}_{\mathrm{Blog}}$ and $\mathcal{C}_{\mathrm{Amazon}}$.
+IM is the only method that tends to be a close second in many corpora, often with substantial overlap in the error bars with LambdaG ($\mathcal{C}_{\mathrm{Enron}}$, $\mathcal{C}_{\mathrm{Wiki}}$, $\mathcal{C}_{\mathrm{ACL}}$, $\mathcal{C}_{\mathrm{PJ}}$, $\mathcal{C}_{\mathrm{Apric}}$, and $\mathcal{C}_{\mathrm{IMDB}}$).
+For $\mathcal{C}_{\mathrm{ACL}}$, although LambdaG is ranked first, there is a substantial overlap in the confidence intervals with both IM and COAV but all these methods are clearly superior to all the neural network approaches.
+There are only two corpora for which LambdaG is not ranked as first: $\mathcal{C}_{\mathrm{Stack}}$ and $\mathcal{C}_{\mathrm{AllNews}}$.
+For $\mathcal{C}_{\mathrm{Stack}}$, the best method is FeVecDiff (LR) but LambdaG is a close second, with the confidence interval substantially overlapping with FeVecDiff (MLP).
+For $\mathcal{C}_{\mathrm{AllNews}}$, LambdaG comes second to ${}^{\,(\bm{\dagger})}$ and this difference is significant.
+However, it is important to note that ${}^{\,(\bm{\dagger})}$ could have exploited any topic bias in the data.
+For all other corpora, LambdaG always performs better than ${}^{\,(\bm{\dagger})}$ and ${}^{\,(\bm{\dagger})}$, despite the fact that any correlation between author and topic could have been exploited in each corpora.
+We do not have an explanation for why ${}^{\,(\bm{\dagger})}$ outperforms all other methods only in $\mathcal{C}_{\mathrm{AllNews}}$, but we advance the hypothesis that $\mathcal{C}_{\mathrm{AllNews}}$ contains data with a significant author/topic bias such that the task is made easier when content words are included.
+This hypothesis is supported by the extremely poor performance of ${}^{\,(\bm{\dagger})}$ on $\mathcal{C}_{\mathrm{Stack}}$, which is a corpus explicitly designed to penalize AV methods that exploit topic information.
 Table 2 instead shows the values of $C_{\text{llr}}$ and $C_{\text{llr}}^{\text{min}}$ for both $\lambda_{G}$ and $\Lambda_{G}$, since both are log-likelihood ratios.
@@
 When $\lambda_{G}$ is turned into $\Lambda_{G}$ by fitting a logistic regression on training data, however, the lower values of $C_{\text{llr}}\mbox{}(\Lambda_{G})$ are evidence that $\Lambda_{G}$ is well calibrated.
-All results in Table 1 and Table 2 were obtained using LambdaG with parameters $r=100$ and $N=10$.
-Figure 1 shows how changes in these hyper-parameters affects Accuracy, revealing that the method is robust in the choice of these hyper-parameters.
-It is clear from the graph that $r=1$, meaning that only one reference Grammar Model $G_{\text{ref}}$ is extracted from the reference corpus, is not enough to achieve optimal results.
-However, in most corpora $r=100$ is not needed as performance stabilize at $r=30$.
-Regarding the other hyperparameter $N$, the order of the $n$-gram model, the value $N=10$ was chosen to ensure that grammatical dependencies are captured even for long sentences and Figure 1 does suggest that this assumption is correct.
-For example, for a corpus such as $\mathcal{C}_{\mathrm{PJ}}$ consisting of chat messages or $\mathcal{C}_{\mathrm{Enron}}$ consisting of emails, $N=2$ is enough.
-In contrast, for the academic papers in $\mathcal{C}_{\mathrm{ACL}}$ and the news in $\mathcal{C}_{\mathrm{AllNews}}$ a larger order of the model clearly has an advantage.
-In all corpora, however, there is no gain in increasing $N$ to 20.
-Figure 1: Accuracy results for different numbers of repetitions $r$ and orders of the $n$-gram model $N$ for each corpus.
-Figure 2: The loss in Accuracy (top) and $C_{\text{llr}}$ (bottom) results for cross-corpus comparison, i.e.
-evaluating on Base Corpus while using reference texts $\mathbb{D}_{\text{ref}}$ from Reference Corpus.
-Diagonal bold values denote the original Accuracy and $C_{\text{llr}}$, respectively.
-Darker shades denote a greater loss in performance.
-Another important robustness test that we performed is the selection of the reference data.
-Reliance on external data is a potential limitation for any binary extrinsic AV methods, like IM.
-As explained in Section 1, IM, suffers a substantial loss in performance if the impostors are selected from another genre.
-LambdaG also has the crucial requirement that a reference corpus $\mathbb{D}_{\text{ref}}$ is collected and it is presumably important that $\mathbb{D}_{\text{ref}}$ is as compatible as possible, especially in genre but generally speaking in as many parameters as possible, with $\mathbb{D}_{\mathcal{A}}$ and $\mathbb{D}_{\mathcal{U}}$.
-To test the robustness of LambdaG to variations in $\mathbb{D}_{\text{ref}}$ type we ran 144 tests by pairing each of the twelve corpora to every other corpora.
-For example, we performed the analysis for $\mathcal{C}_{\mathrm{Enron}}$ but we used as $\mathbb{D}_{\text{ref}}$ $\mathcal{C}_{\mathrm{ACL}}$.
-We then calculated Accuracy and $C_{\text{llr}}$ for all combinations and we produced a plot for each metric displaying the resulting drop in performances when using a different $\mathbb{D}_{\text{ref}}$ (Figure 2).
 Corpus $C_{\text{llr}}\mbox{}(\lambda_{G})$ $C_{\text{llr}}\mbox{}(\Lambda_{G})$ $C_{\text{llr}}^{\text{min}}$
@@
 $C_{\text{llr}}^{\text{min}}$ is identical for both $\lambda_{G}$ and $\Lambda_{G}$.
-The diagonal of Figure 2 (top) shows the same results reported in Table 1, which function as the benchmark of performance for each row of the plot, corresponding to the corpus being tested.
-Overall the results suggest that LambdaG is extremely robust in many cases of genre variation, including some unrealistic genre combinations.
-This result is unexpected given how IM is instead very much dependent on the set of impostors [41].
-For example, cross-corpus comparisons such as using $\mathcal{C}_{\mathrm{Yelp}}$ as reference for problems in $\mathcal{C}_{\mathrm{Amazon}}$ lead to very small losses in Accuracy and this is clearly explainable by the fact that both corpora contain reviews.
-However, certain results are more surprising, such as that using the academic papers from $\mathcal{C}_{\mathrm{ACL}}$ as reference data for problems in a corpus of chat logs such as $\mathcal{C}_{\mathrm{PJ}}$ leads to a performance loss of only 0.27 Accuracy, which is substantial but much less than expected given the disparity.
+All results shown in Table LABEL:tab:EvalResults and Table 2 were obtained with LambdaG using the hyperparameters $r=100$ and $N=10$.
+Figure 3 shows how changes in these hyperparameters affects Accuracy, revealing that the method is robust in the choice of these hyperparameters.
+It is clear from the graph that $r=1$, meaning that only one reference Grammar Model $G_{\text{ref}}$ is extracted from the reference corpus, is not enough to achieve optimal results.
+However, in most corpora $r=100$ is not needed as performance stabilize at $r=30$.
+Regarding the other hyperparameter $N$, the order of the $n$-gram model, the value $N=10$ was chosen to ensure that grammatical dependencies are captured even for long sentences and Figure 3 does suggest that this assumption is correct.
+For example, for a corpus such as $\mathcal{C}_{\mathrm{PJ}}$ consisting of chat messages or $\mathcal{C}_{\mathrm{Enron}}$ consisting of emails, $N=2$ is enough.
+In contrast, for the academic papers in $\mathcal{C}_{\mathrm{ACL}}$ and the news in $\mathcal{C}_{\mathrm{AllNews}}$ a larger order of the model clearly has an advantage.
+In all corpora, however, there is no gain in increasing $N$ to 20.
+Figure 3: Variation in Accuracy depending on the number of repetitions, $r$, and orders of the $n$-gram model, $N$.
+Another important robustness test that we performed is the selection of the reference data.
+Reliance on external data is a potential limitation for all binary extrinsic AV methods such as IM.
+As explained in Section 1, IM, suffers a substantial loss in performance if the impostors are selected from another genre.
+LambdaG has the same crucial requirement that a reference corpus $\mathbb{D}_{\text{ref}}$ is collected and it is presumably important that $\mathbb{D}_{\text{ref}}$ is as compatible as possible, especially in genre, with $\mathcal{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$.
+To test the robustness of LambdaG to variations in $\mathbb{D}_{\text{ref}}$ type we ran 144 tests by pairing each of the twelve corpora to every other corpora.
+For example, we performed the analysis for $\mathcal{C}_{\mathrm{Enron}}$ but we used as $\mathbb{D}_{\text{ref}}$ $\mathcal{C}_{\mathrm{ACL}}$.
+We then calculated Accuracy and $C_{\text{llr}}$ for all combinations and we produced a plot for each metric displaying the resulting drop in performances when using a different $\mathbb{D}_{\text{ref}}$ (Figure 4).
+The diagonal of Figure 4 (top) shows the same results reported in Table LABEL:tab:EvalResults, which function as the benchmark of performance for each row of the plot, corresponding to the corpus being tested.
+These results show that, as expected, the performance of LambdaG can be strongly affected by genre variations in the reference corpus.
+However, the performance loss overall is not as dramatic as it would be expected for a method such as IM [51].
+Firstly, for reasonable cross-corpus comparisons such as using $\mathcal{C}_{\mathrm{Yelp}}$ as reference for problems in $\mathcal{C}_{\mathrm{Amazon}}$, where both corpora are reviews, the loss in Accuracy is small.
+Secondly and more surprisingly, ‘unreasonable’ comparisons such using the academic papers from $\mathcal{C}_{\mathrm{ACL}}$ as reference data for problems in a corpus of chat logs such as $\mathcal{C}_{\mathrm{PJ}}$ leads to a reduction of, relatively speaking, only 0.27 Accuracy.
+Such a loss is clearly substantial but the expectation given the literature is that the method should perform at random in such a case.
 This is a comparison that a human analyst would be unlikely to perform and therefore it is a real test of the lower limits in terms of robustness to reference corpora.
-Looking at Figure 2 (top) row-wise, the most notable pattern is that the two corpora that suffer the most performance loss overall are $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$.
-For $\mathcal{C}_{\mathrm{PJ}}$ the explanation is likely to be that it is the only corpus with sentences mostly corresponding to chat messages, and therefore any match of this corpus with other corpora results in at least this incompatibility.
+Examining Figure 4 (top) row-wise, the most notable pattern is that the two corpora that suffer the most performance loss overall are $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$.
+For $\mathcal{C}_{\mathrm{PJ}}$ the explanation is likely to be that it is the only corpus with sentences corresponding to chat messages, and therefore any match of this corpus with other corpora suffers from this important difference.
 The chat logs in $\mathcal{C}_{\mathrm{PJ}}$ are also the texts with the lowest level of formality and highest degree of linguistic freedom among all corpora.
@@
 When looking at Figure 4 column-wise, instead, the most notable pattern is that using $\mathcal{C}_{\mathrm{ACL}}$ or $\mathcal{C}_{\mathrm{AllNews}}$ as reference corpora tends to affect performance more drastically for the majority of corpora, which again could be explained by their relatively higher degree of formality compared to the other corpora.
-All in all this analysis suggests that, for $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$ the reference corpus has a moderate impact on performance while for all other corpora it is only the use of $\mathcal{C}_{\mathrm{ACL}}$ or $\mathcal{C}_{\mathrm{AllNews}}$ as reference where this modest impact can be observed.
-For all other cross-corpus comparisons for most corpora the effect of the reference corpus is tolerable.
+All in all, this analysis suggests that, for $\mathcal{C}_{\mathrm{PJ}}$ and $\mathcal{C}_{\mathrm{Stack}}$ the reference corpus has a moderate impact on performance, while for all other corpora it is only the use of $\mathcal{C}_{\mathrm{ACL}}$ or $\mathcal{C}_{\mathrm{AllNews}}$ as a reference where this modest impact can be observed.
+For most other cross-corpus comparisons, the effect of using a different reference corpus is tolerable.
 We also present the results for the $C_{\text{llr}}$ metric in Figure 4 (bottom).
-As explained in Section 5.4, a $C_{\text{llr}}$ value close to or equal to one means that the system is not providing any useful information for the case, while a $C_{\text{llr}}$ value higher than 1 means that results tend to be misleading (e.g.
-negative $\Lambda_{G}$ for a Y-problem).
-None of the cross-genre comparisons produce $C_{\text{llr}}$ higher than one, meaning that in all cases the values of $\Lambda_{G}$ are not misleading the analyst but can be, at worst, inconclusive with $\Lambda_{G}\approx 0$.
+As explained in Section 6.4, a $C_{\text{llr}}$ value close to or equal to one means that the system is not providing any useful information, while a $C_{\text{llr}}$ value higher than 1 means that results tend to be misleading (e.
+g., negative $\Lambda_{G}$ for a Y-problem).
+None of the cross-genre comparisons produce $C_{\text{llr}}$ higher than one, meaning that in all cases, on average, the values of $\Lambda_{G}$ are not misleading the analyst but can be, at worst, inconclusive with $\Lambda_{G}\approx 0$.
 These results support the principle that an analyst should always compile a reference data set that is as close as possible to the case data [65].
-However, the results do also suggest that LambdaG is stable to small variations in genre.
-The final test we carried out is a comparison of LambdaG as implemented in Algorithm 1 vs.
+However, the results do also suggest that LambdaG is certainly stable to small variations in genre or reference corpus composition.
+Figure 4: The loss in Accuracy (top) and $C_{\text{llr}}$ (bottom) results for cross-corpus comparison, i.
+e., evaluating on Base Corpus while using reference texts $\mathbb{D}_{\text{ref}}$ from Reference Corpus.
+Diagonal bold values denote the original Accuracy and $C_{\text{llr}}$, respectively.
+Darker shades denote a greater loss in performance.
+The final tests we carried out relate to the contribution of POSNoise to the algorithm.
+We firstly performed a comparison of LambdaG as implemented in Algorithm 1 vs.
 LambdaG without applying the POSNoise function.
 The rationale for this test is to understand whether more authorial information can be extracted and exploited using this method beyond just grammatical structure.
-The results in Table LABEL:tab:EvalResultsPosnoise show that, regarding Accuracy and AUC, in all cases the proposed version of Algorithm 1 with the POSNoise pre-processing is better performing than a version of the algorithm without this pre-processing.
-The difference in Accuracy is not large but the method is clearly superior when the focus is entirely on grammar to the extent that adding information about meaning instead leads to adding noise in the authorial signal.
+The results in Table 3 show that, regarding Accuracy and AUC, in all cases the proposed version of Algorithm 1 with the POSNoise pre-processing is better performing than a version of the algorithm without this pre-processing.
+The difference in Accuracy is not large but it is systematic across all corpora, meaning that the method is clearly superior when the focus is entirely on grammar.
+$\Delta$Acc. $\Delta$AUC $\Delta$F1 $\Delta$Prec. $\Delta$Rec. $\Delta$TP $\Delta$FN $\Delta$FP $\Delta$TN
+$\mathcal{C}_{\mathrm{Enron}}$ 0.031 0.039 0.039 -0.017 0.083 4 -4 1 -1
+$\mathcal{C}_{\mathrm{Wiki}}$ 0.026 0.003 0.029 0.022 0.036 4 -4 -2 2
+$\mathcal{C}_{\mathrm{Stack}}$ 0.005 0.001 -0.002 0.047 -0.044 -5 5 -6 6
+$\mathcal{C}_{\mathrm{ACL}}$ 0.057 0.050 0.031 0.106 -0.050 -7 7 -23 23
+$\mathcal{C}_{\mathrm{PJ}}$ 0.009 0.005 0.010 0.001 0.019 3 -3 0 0
+$\mathcal{C}_{\mathrm{Apric}}$ 0.077 0.042 0.084 0.038 0.129 22 -22 -4 4
+$\mathcal{C}_{\mathrm{Trip}}$ 0.037 0.049 0.048 0.022 0.071 17 -17 -1 1
+$\mathcal{C}_{\mathrm{Yelp}}$ 0.057 0.046 0.063 0.043 0.083 20 -20 -7 7
+$\mathcal{C}_{\mathrm{IMDB}}$ 0.080 0.101 0.091 0.070 0.108 87 -87 -42 42
+$\mathcal{C}_{\mathrm{Blog}}$ 0.063 0.038 0.066 0.048 0.083 75 -75 -38 38
+$\mathcal{C}_{\mathrm{Amazon}}$ 0.046 0.030 0.046 0.041 0.050 60 -60 -49 49
+$\mathcal{C}_{\mathrm{AllNews}}$ 0.043 0.049 0.059 0.034 0.076 101 -101 -15 15
+Table 3: Differences in evaluation results when applying POSNoise in LambdaG compared to omitting it.
+Positive values (shaded in blue) indicate results improvement, while negative values (shaded in red) indicate results degradation by using POSNoise.
+$\Delta$Acc. $\Delta$AUC $\Delta$F1 $\Delta$Prec. $\Delta$Rec. $\Delta$TP $\Delta$FN $\Delta$FP $\Delta$TN
+$\mathcal{C}_{\mathrm{Enron}}$ 0.010 0.010 0.014 -0.019 0.042 2 -2 1 -1
+$\mathcal{C}_{\mathrm{Wiki}}$ 0.026 0.014 0.027 0.036 0.018 2 -2 -4 4
+$\mathcal{C}_{\mathrm{Stack}}$ $0$0.004 -0.001 0.008 -0.009 -1 1 -1 1
+$\mathcal{C}_{\mathrm{ACL}}$0.022 -0.014 -0.019 -0.033 -0.007 -1 1 5 -5
+$\mathcal{C}_{\mathrm{PJ}}$ 0.016 0.000 0.017 0.001 0.032 5 -5 0 0
+$\mathcal{C}_{\mathrm{Apric}}$0.003 -0.004 -0.003 $0$0.006 -1 1 0 0
+$\mathcal{C}_{\mathrm{Trip}}$ 0.014 0.010 0.019 0.007 0.029 7 -7 0 0
+$\mathcal{C}_{\mathrm{Yelp}}$0.014 -0.010 -0.019 -0.003 -0.038 -9 9 -2 2
+$\mathcal{C}_{\mathrm{IMDB}}$ 0.015 0.013 0.014 0.019 0.010 8 -8 -16 16
+$\mathcal{C}_{\mathrm{Blog}}$ 0.006 0.002 0.006 0.005 0.006 6 -6 -5 5
+$\mathcal{C}_{\mathrm{Amazon}}$0.002 -0.001 -0.003 0.003 -0.009 -10 10 -5 5
+$\mathcal{C}_{\mathrm{AllNews}}$ 0.027 0.031 0.040 0.015 0.057 76 -76 5 -5
+Table 4: Differences in evaluation results when applying POSNoise in LambdaG compared to applying the POSNoise without POS-Tagging (i.
+e., using a simple star).
+Positive values (shaded in blue) indicate results improvement, while negative values (shaded in red) indicate results degradation by using full POSNoise.
+Secondly, we evaluated LambdaG using a variant of POSNoise similar to TextDistortion, but instead of utilizing POS tags, we replaced all non-functional items with asterisks.
+The results, presented in Table 4, indicate that the POS tag labels contribute very little information overall, implying that a POS tagger might not be necessary.
 As explained in Section 6.3, the advantage of LambdaG is that it can be better interpreted than other methods.
-As a first step, the analyst can produce a ranking of all sentences in $\mathcal{D}_{\mathcal{U}}$ and then, additionally, each token can be color-coded according to its value of $\lambda_{G}$, similarly to COAV.
+The analyst can produce a ranking of all sentences in $\mathcal{D}_{\mathcal{U}}$ and each token can be color-coded according to its value of $\lambda_{G}$, similarly to COAV.
 Some of these visualizations explaining the features discovered are shown in Section 7.
-$\Delta$Acc. $\Delta$AUC $\Delta$F1 $\Delta$Prec. $\Delta$Rec. $\Delta$TP $\Delta$FN $\Delta$FP $\Delta$TN
-$\mathcal{C}_{\mathrm{Enron}}$
-0.031 0.039 0.039 -0.017 0.083 4 -4 1 -1
-$\mathcal{C}_{\mathrm{Wiki}}$
-0.026 0.003 0.029 0.022 0.036 4 -4 -2 2
-$\mathcal{C}_{\mathrm{Stack}}$
-0.005 0.001 -0.002 0.047 -0.044 -5 5 -6 6
-$\mathcal{C}_{\mathrm{ACL}}$
-0.057 0.050 0.031 0.106 -0.050 -7 7 -23 23
-$\mathcal{C}_{\mathrm{PJ}}$
-0.009 0.005 0.010 0.001 0.019 3 -3 0 0
-$\mathcal{C}_{\mathrm{Apric}}$
-0.077 0.042 0.084 0.038 0.129 22 -22 -4 4
-$\mathcal{C}_{\mathrm{Trip}}$
-0.037 0.049 0.048 0.022 0.071 17 -17 -1 1
-$\mathcal{C}_{\mathrm{Yelp}}$
-0.057 0.046 0.063 0.043 0.083 20 -20 -7 7
-$\mathcal{C}_{\mathrm{IMDB}}$
-0.080 0.101 0.091 0.070 0.108 87 -87 -42 42
-$\mathcal{C}_{\mathrm{Blog}}$
-0.063 0.038 0.066 0.048 0.083 75 -75 -38 38
-$\mathcal{C}_{\mathrm{Amazon}}$
-0.046 0.030 0.046 0.041 0.050 60 -60 -49 49
-$\mathcal{C}_{\mathrm{AllNews}}$
-0.043 0.049 0.059 0.034 0.076 101 -101 -15 15
 5 Discussion
-Our proposed LambdaG approach matches many desiderata of an AV method:
-  1. 1.
-    It works on both short and long documents;
-  2. 2.
-    It is not affected or minimally affected by topic or content;
-  3. 3.
-    It is stable to small genre variations in the reference corpus;
-  4. 4.
-    It is more interpretable by an analyst;
-  5. 5.
-    It is not computationally intensive and does not have a high run time;
-  6. 6.
-[…]
-  7. 7.
-    It has a theoretical foundation based on Cognitive Linguistic theories of language processing.
-Section 3 showed evidence of the first six of these points as well as evidence that LambdaG outperforms other AV state-of-the-art methods.
-We believe that a possible reason to explain these results, especially considered the simplicity of the method compared to approaches such as SiamBERT, is precisely that LambdaG has a number of properties that make it consistent with knowledge about language processing gained from studies in Cognitive Science and Cognitive Linguistics.
-Since the pioneering study of Mosteller and Wallace [79], it is well known that the frequency of functional items distinguish authors and this conclusion is also evident in the case of LambdaG.
+We believe that our proposed LambdaG approach matches many desiderata of an AV method.
+The evaluation on our set of 12 corpora demonstrates that LambdaG works consistently well for relatively short texts in a variety of different genres.
+In addition, because the method only considers information in functional items, LambdaG is also minimally affected by potential content biases in the data, such as named entities or topic.
+In contrast to the other very successful binary-extrinsic AV method, IM, LambdaG tends to be much more robust to variations in the reference corpus.
+Moreover, unlike IM, which relies on the similarity between vectors of thousands of character $n$-grams, LambdaG generates text heat maps that aid the analyst’s understanding of the output of the method.
+Finally, LambdaG is unique among other AV methods in being motivated by Cognitive Linguistics and in having a theoretical explanation for its mechanisms.
+We believe that its greater scientific plausibility justifies LambdaG over other state-of-the-art methods.
+Since the pioneering study of Mosteller and Wallace [94], it is well known that the frequency of functional items distinguish authors and this conclusion is also evident in our results.
 The reason why this is the case, however, has always been seen as a mystery in the field of authorship analysis and stylometry [69].
 A connection between the unreasonable effectiveness of these functional items and Cognitive Linguistics has been proposed by Nini [97] and we believe that the success of LambdaG is additional evidence for these explanations.
-Cognitive Linguistics models language as a complex adaptive system [9], where regularities emerge from the complexity of language used in interactions, like a phenomenon of the third kind [33].
-This is partially evidenced by decades of linguistics studies on corpora which have revealed that the traditional conception of a grammar as an abstract set of phrase structure rules is untenable and that, instead, a better model of grammar is a lexicogrammar continuum that goes from fixed lexicalized expressions (e.
-g., “I don’t know”) to flexible templates (e.
-g., “the” X-er “the” Y-er) up to completely abstract structures (e.
-g., Doer Action Object).
-This continuum roughly corresponds to the continuum between declarative memory, which is semantic/encyclopedic memory for facts and which tends to be conscious, and procedural memory, which is instead memory for the probabilistic prediction of patterns of sequences and that deals with subconscious knowledge of habits and skills [108].
-The roles that these two memories play in language is clearly evidenced in cases of aphasia.
-When an impairment affects circuits of declarative memory, this gives rise to fluent aphasia, resulting in speech that is mostly constituted by function words like articles, prepositions, pronouns, and morphological forms but with lack of content words that contain meaning, leading to fluent but nonsensical speech comparable to POSNoise pre-processed data such as the one in Table 5.
-Instead, when the impairment is on procedural memory, then speech is only consisting of content words that are not connected through grammar [109].
-These two memory systems work together and, typically, once a habit is formed through repetition then procedural memory takes over, leading to fluency of expression.
-This transition explains why languages tend to be heavily formulaic.
-For example, in English one can say big mistake but not large mistake, even though the formal grammar would seemingly have no problems with the latter expression, which is an adjective modifying a noun.
-What explains this reliance on formulaic expressions is the cognitive tendency to recur to memory as much as possible rather than creating something new every time we use language [108].
-The way these units of procedural memory are formed is through the process of chunking: when two items that often go together are found together, then these two items can be treated as one, thus saving space in working memory [39, 21, 28].
-In all effects, chunking is a process of information compression and, if the units of grammar are chunks, then the logical inference is that grammar itself can be thought of as a compressed version of a language or, equivalently, as a probability model.
-A probability model of only functional items is therefore quite like the Grammar Models used in this study.
-Cognitive Linguistics also explains why each individual has a different and possibly unique grammar, a statement encapsulated by the Principle of Linguistic Individuality [82].
-The status of unithood given by chunking is not discrete (e.
-g., something is or is not a unit) but a continuum of strength of unithood, where this strength is called entrenchment [32, 97].
-Because entrenchment is dependent on exposure and usage, it is fundamentally personal and idiosyncratic.
-It is for this reason that Langacker hypothesized that “the set of units mastered by all members of a speech community might be a rather small proportion of the units constituting the linguistic ability of any given speaker” [71], a hypothesis that is finding increasing empirical support [82].
-For example, for most if not all speakers of English the entrenchment for a sequence such as “of the” is maximal, meaning that they can produce it effortlessly and unconsciously as a single unit.
-In contrast, a sequence such as “the of” would not be a unit for speakers of English because it is virtually never encountered or used.
-If a speaker of English encounters this sequence, they would decompose it as a structure made up of two units, “the” and “of”.
-In between these two extreme examples there is a lot of freedom as to what counts as a unit for a person or another, depending on what each individual has automatized in their procedural memory.
-For example, an individual $\mathcal{A}_{1}$ might have a very entrenched sequence such as “it is known that” when writing academic articles while an individual $\mathcal{A}_{2}$ might have “it is well established that”.
-These sequences will then have open slots ready to be filled in with the appropriate meaning.
-This kind of phenomenon theorized by Langacker can be seen at play in this method, especially when looking at the examples of idiosyncratic structures in Section 6.
-As briefly mentioned in Section 5.3, therefore, the likelihood of a token given a context assigned by an author $\mathcal{A}$’s Grammar Model, $P(t|c;G_{\mathcal{A}})$, can be seen as a way to measure the level of entrenchment of token $t$ in context $c$ for $\mathcal{A}$ The higher the entrenchment, the more likely it is that $\mathcal{A}$ is using procedural memory and that it is therefore performing an unconscious routine that constitutes a unit for this individual.
-There have been other attempts at modeling the grammar of an author in the past for AV or AA purposes [24, 7, 37] but these methods relied too much on Part of Speech tags or on syntactic trees which are often not compatible with predictions made in Cognitive Science and Cognitive Linguistics.
-As explained above, the picture of grammar painted by Cognitive Linguistics is much more lexically-based and word-idiosyncratic.
-For example, the preposition “of” in English behaves very oddly compared to other prepositions [100].
-Giving it the same Part of Speech label as to other prepositions is therefore going to obscure a lot of important grammatical information.
-Instead, a topic-masking method such as POSNoise is much more consistent with the predictions of Cognitive Linguistics because it leaves the function items unaffected while only obscuring those items are more likely to change from context to context depending on the communicative needs.
-To sum up, the model of an individual’s grammar that emerges from Cognitive Science and Cognitive Linguistics is one that is very much compatible with the Grammar Models used for LambdaG.
-Although language modeling is seen in computational linguistics as a tool and not necessarily a realistic representation of language, evidence from Cognitive Linguistics suggests that, in reality, a language model, while not being realistic as a model of linguistic production, is instead a plausible representation of how knowledge of grammar is stored in the mind.
+As outlined in Section 2, the mental representation of the grammar for an individual can be seen as a a probability model over lexicogrammatical chunks quite like the Grammar Models used in LambdaG.
+Although language modeling is seen in computational linguistics as a tool and not necessarily a realistic representation of language, evidence from Cognitive Linguistics suggests that, in reality, a language model, while not being realistic as a model of linguistic production, is instead a plausible representation of how knowledge of grammar is stored in the mind, as well as an in silico model of the brain language system [127].
 The mechanisms of such a sequential processing can be efficiently represented through the short-term correlations captured by $n$-gram models as done in this work.
 A deeper connection between LambdaG and approaches to AV based on language modeling or, as for COAV, on compression algorithms can also be drawn, thus offering a plausible explanation for their success.
-Through the lenses of Information Theory, the average $P(s;G)$ for each $s\in\mathcal{D}$ can be also seen as an estimate of the cross-entropy rate of the Grammar Model $G$ against the probability distribution of grammatical tokens that generates $\mathcal{D}$.
+Through the lenses of Information Theory, the average $P(S;G)$ for each sentence $S$ in $\mathcal{D}$ can be also seen as an estimate of the cross-entropy rate of the Grammar Model $G$ against the probability distribution of grammatical tokens that generates $\mathcal{D}$.
 In language modeling literature, the cross-entropy is often exponentiated to calculate the perplexity of a document given a model.
@@
 In addition to the successful results of LambdaG, we note that the qualitative explorations reported in Section 7 also constitute evidence that LambdaG captures linguistic behavior compatible with these Cognitive Linguistic explanations.
+That individuals vary greatly in their mental grammar, as theorized by Langacker, can be seen at play in these text heat maps.
+The theoretical interpretation of these results is that a stronger shade of red effectively means a greater likelihood that the token is entrenched for that individual’s unique grammar in contrast to the population.
 Although it is impossible to know for certain the effective mental status of the sequences found in the examples in Section 7, we propose that it is plausible to believe that the authors of these texts do not know that they use the sequences of tokens highlighted in red or that these sequences identify them in the general population.
 These sequences, which are often seemingly unremarkable (e.
-g.,enough to be ADVERB VERB” or “so PRONOUN cant”), can be extremely idiosyncratic and uniquely used only by the candidate author.
+g.,enough to be ADVERB VERB” or “so PRONOUN cant”[sic]), can be extremely idiosyncratic and uniquely used only by the candidate author.
 Given the fact that the problems analyzed in this paper involve relatively short texts (the longest texts are just around 2000 tokens), we must conclude that these idiosyncratic units are easy to find and this conclusion therefore strongly supports Langacker’s proposal that most units in a language are not shared by many individuals.
@@
 If individual grammars are so different from each other, then the role of the reference corpus could be to just weight down the those units that are well-entrenched for most individuals (e.
-g., of the) and that therefore carry almost no authorial information.
+g.,of the”) and that therefore carry almost no authorial information.
 If this explanation is correct, then this means that once these common units are taken into account, the rest is idiosyncratic enough to be identifying.
@@
 We propose that this is the explanation underlying most if not all AV and AA methods that, by measuring the frequency of functional items, are indirectly analyzing the idiosyncratic grammar of an individual.
-LambdaG tends to be more successful probably because it is measuring the same phenomenon more directly.
+It is possible that LambdaG is more successful because it is measuring this same phenomenon more directly.
 5.1 Limitations and Future Research
 We argue that future research should therefore continue in this direction of extracting better representations of the grammar of an individual.
-The POSNoise algorithm used here is language dependent and suffers from the use of general part of speech categories that are not compatible with Cognitive Linguistics, which instead predicts the existence of categories that are more semantic in nature and at lower degrees of abstractness.
+The POSNoise algorithm used here is language dependent and suffers from the use of general part of speech categories that are actually not compatible with Cognitive Linguistics, which instead predicts the existence of categories that are more semantic in nature and at lower degrees of abstraction.
 New future methods should embrace the close relationship between Cognitive Linguistics and Information Theory to gain more insight into grammar, which could consequently lead to further breakthroughs in AV.
@@
 Despite these limitations, future methods based on a better understanding of grammar could still lead to more progress.
-In this respect, the reliance on a reference corpus could also be seen as a limitation of any binary AV method.
-The tendency for future AV methods should be to increase the power of unary methods that only need a test dataset.
+In this respect, the reliance on a reference corpus could also be seen as a limitation of any binary-extrinsic AV method.
+The tendency for future AV methods should be to increase the power of unary methods that consider only the documents within a verification case.
+LambdaG also suffers from two limitations that also apply to IM because of their bootstrapping component: both algorithms are stochastic and a higher run time due to the number of repetitions.
+In our experiment we noted that LambdaG is faster than IM, although this depends on the specific implementation.
+Our test of the $r$ hyperparameter also shows that 30 repetitions are enough for best performance, in contrast to the 100 iterations in IM.
+In addition, the computational complexity of LambdaG is still very limited compared to any method using neural networks and running on GPUs.
+To mitigate the effect of randomness in the bootstrapping, all experiments in this paper were run five times.
+However, a deterministic algorithm would definitely be preferable.
 Finally, another limitation of the present paper is that the research was carried out entirely on English.
@@
 Different tokenizers might have to be employed as well as different algorithms such as POSNoise to focus the model on the grammatical aspects of the text.
-Another goal for the future would therefore be to find a language-independent variant of LambdaG.
+Our test showing that POS tags are not necessary is also encouraging for the replicability of this algorithm for languages with less computational resources.
+In general, though, another goal for the future would therefore be to find a language-independent variant of LambdaG.
 5.2 Final Remarks
@@
 Secondly, knowledge of Linguistics is needed to understand and potentially explain in a court of law the reasons why these patterns identify or exclude a certain author.
-We believe that the work in this paper has made some significant advances in AV while also proposing and stimulating a paradigm change to move on from the impasse described by Stamatatos [104] towards new AV methods that are not only well-performing but also scientifically meaningful.
+We believe that the work in this paper has made some significant advances in AV while also proposing and stimulating a paradigm change to move on from the impasse described by Stamatatos [121] towards new AV methods that are not only well-performing but also scientifically plausible.
 6 Methods
-In this section, we present the more technical details of our work, first with an introduction to the corpora we have compiled to evaluate LambdaG, then by providing more details about the POSNoise content-masking algorithm as well as the details of the Kneser-Ney smoothing algorithm, the likelihood ratio framework for forensic sciences and, finally, the Siamese neural network baseline method, SiamBERT, developed for this analysis.
+In this section, we present the more technical details of our work, first with an introduction to the corpora we have compiled to evaluate LambdaG, then by providing more details about the POSNoise content-masking algorithm as well as the details of the Kneser-Ney smoothing algorithm, the likelihood ratio framework for forensic sciences and, finally, our own baseline method, SiamBERT, developed for this analysis.
 6.1 Corpora
@@
 Each $c\in\mathcal{C}$ denotes a verification case $(\mathcal{D}_{\mathcal{U}},\mathbb{D}_{\mathcal{A}})$, where $\mathcal{D}_{\mathcal{U}}$ is an unknown document and $\mathbb{D}_{\mathcal{A}}$ is a set of example documents of the known author $\mathcal{A}$.
-To counteract population homogeneity, a known bias in authorship verification described by Bevendorff [11], we ensured that for each author $\mathcal{A}$ there are exactly two (one same-authorship (Y) and one different-authorship (N)) verification cases.
+To counteract population homogeneity, a known bias in AV described by Bevendorff [14], we ensured that for each author $\mathcal{A}$ there are exactly two verification cases i.
+e., one with same-authorship (Y) and one with different-authorship (N).
 Furthermore, we constructed all corpora so that the number of Y- and N-cases is equal, in other words, all training and test corpora are (strictly) balanced.
-In what follows, we describe all corpora in detail, underline their respective challenges and explain the pre-processing [^2] The pre-processing of the documents was carried out using TextUnitLib https://github.com/Halvani/TextUnitLib steps we have performed.
+In what follows, we describe all corpora in detail, underline their respective challenges and explain the respective pre-processing steps (using TextUnitLib [52]) we have performed for each corpus.
 A compact overview of the 12 corpora, including their partitioning into training and test corpora, as well as further information can be found in Table 5.
@@
 $\mathcal{C_{\,\textrm{test}}}$ 2662 680 (3665) 942 (5071)
-Table 4: All used corpora and their key statistics.
+Table 5: All corpora used and their key statistics.
 Notation: $|\mathcal{C}|$ denotes the number of verification cases in each corpus $\mathcal{C}$, while avg$(\mathbb{D}_{\mathcal{A}})$ and avg$(\mathcal{D}_{\mathcal{U}})$ denotes the average token count (and character count in parentheses) of all $\mathcal{D}_{\mathcal{A}}\in\mathbb{D}_{\mathcal{A}}$ and $\mathcal{D}_{\mathcal{U}}$, respectively.
 6.1.1 $\mathcal{C}_{\mathrm{Enron}}$
-The $\mathcal{C}_{\mathrm{Enron}}$ corpus was derived from the well-known Enron Email Dataset333available at https://www.cs.cmu.edu/~enron, which has been used in many AV and AA studies, including [17, 18, 19, 27, 31, 72, 83, 112, 4], as well as in many other research areas.
+The $\mathcal{C}_{\mathrm{Enron}}$ corpus was derived from the well-known Enron Email Dataset [75], which has been used in many AV and AA studies, including [23, 24, 25, 33, 38, 86, 98, 136, 6], as well as in many other research areas.
 The Enron email dataset has gained considerable popularity in the authorship analysis community, likely due to the fact that it is (to the best of our knowledge) the only publicly available corpus of real-world emails with a guaranteed ground truth.
@@
 6.1.2 $\mathcal{C}_{\mathrm{Wiki}}$
-The $\mathcal{C}_{\mathrm{Wiki}}$ corpus consists of 752 excerpts from 288 Wikipedia talk page editors taken from the Wikipedia sockpuppets dataset, which was released by Solorio et al.
+The $\mathcal{C}_{\mathrm{Wiki}}$ corpus consists of 752 excerpts from 288 Wikipedia talk page editors taken from the Wikipedia sockpuppets dataset, released by Solorio et al.
 [116].
@@
 6.1.3 $\mathcal{C}_{\mathrm{Stack}}$
-The $\mathcal{C}_{\mathrm{Stack}}$ corpus consists of 567 posts from 189 users crawled from the question-and-answer (Q & A) network Stack Exchange [^4] https://stackexchange.com in 2019.
+The $\mathcal{C}_{\mathrm{Stack}}$ corpus consists of 567 posts from 189 users crawled from the question-and-answer (Q & A) network Stack Exchange [117] in 2019.
 The network comprises 173 Q & A communities, with each community focusing on a specific topic (e.
@@
 e., N-cases are more likely to be classified as Y and vice versa) while AV methods that use topic-agnostic features behave more robustly.
-Another challenge of $\mathcal{C}_{\mathrm{Stack}}$ is that the corpus is very unbalanced with respect to the length of the documents, with lengths between 3014 and 20476 characters.
-To avoid stylistic distortions from the outset, we only used the original posts of the respective authors for all texts in $\mathcal{C}_{\mathrm{Stack}}$, while subsequent edits by third parties (for the purpose of spelling or stylistic improvement) were excluded [^5] Stack Exchange provides metadata to show exactly which parts of the text have been edited, and also allows the original unedited post to be retrieved..
+Another challenge of $\mathcal{C}_{\mathrm{Stack}}$ is that the corpus is highly unbalanced in terms of document length, with lengths between 3,014 and 20,476 characters.
+To avoid stylistic distortions from the outset, we only used the original posts of the respective authors for all texts in $\mathcal{C}_{\mathrm{Stack}}$, while subsequent edits by third parties (for the purpose of spelling or stylistic improvement) were excluded (Stack Exchange provides metadata to show exactly which parts of the text have been edited, and also allows the original unedited post to be retrieved.).
 In addition, we replaced equations/formulas, numerical values (e.
 g., percentages and currency values), dates, references and URLs with appropriate placeholders.
-The reason why we did not omit the affected sentences and instead opted for this approach was that we wanted to keep as much text content per author as possible to ensure a sufficient number of authors.
+The reason that we did not leave out the affected sentences and instead chose this approach was that we wanted to retain as much text content per author as possible to ensure a sufficient number of authors.
 6.1.4 $\mathcal{C}_{\mathrm{ACL}}$
-The $\mathcal{C}_{\mathrm{ACL}}$ corpus comprises 466 excerpts from scientific papers by a total of 233 researchers, taken from the computational linguistics archive ACL Anthology [^6] https://aclanthology.org/.
-The corpus was constructed in such a way that there are exactly two papers for each author [^7] Based on the metadata of the papers (more precisely, the author fields within the corresponding BibTex entries), we ensured that each paper is from exactly one author., dating from different time periods.
+The $\mathcal{C}_{\mathrm{ACL}}$ corpus comprises 466 excerpts from scientific papers by a total of 233 researchers, taken from the computational linguistics archive ACL Anthology [^8] https://aclanthology.org.
+The corpus was constructed in such a way that there are exactly two papers for each author [^9] Using the metadata of the papers (more precisely, the author fields in the corresponding BibTex entries), we ensured that each paper was written by exactly one author., dating from different time periods.
 The average time span between the two documents of each author is approximately 15.6 years, while the minimum and maximum time spans are 8 and 31 years, respectively.
@@
 In contrast to $\mathcal{C}_{\mathrm{Enron}}$, we have automatically pre-processed the texts in this corpus.
-Among other things, we sorted out chat lines with less than 5 tokens and 10 characters or those containing usernames, timestamps, URLs and words with many repeating characters (e.
+Among other things, we removed chat lines with less than 5 tokens and 10 characters or those containing usernames, timestamps, URLs and words with many repeating characters (e.
 g., “mmmhhhmm”, “heeheehee” and “:-*:-*:-*:-*:-*:-*:-*:-*”).
@@
 For this purpose, we applied a string similarity function to each pair $(x_{i},x_{j})$ with $i\neq j$.
-As a metric for string similarity, we chose the Levenshtein distance [^9] Regarding the fuzzy string matching, we used FuzzyWuzzy https://github.com/seatgeek/thefuzz based on a threshold of 0.25, which we considered appropriate for our purposes.
+As a metric for string similarity, we chose the Levenshtein distance (using TheFuzz library [112]) based on a threshold of 0.25, which we considered appropriate for our purposes.
 All pairs of lines that exceeded this threshold were discarded with respect to the text fragments.
 6.1.6 $\mathcal{C}_{\mathrm{Apric}}$
-The $\mathcal{C}_{\mathrm{Apric}}$ corpus comprises 1,395 forum posts from 284 users crawled from the portal The Apricity - A European Cultural Community [^10] https://theapricity.com.
+The $\mathcal{C}_{\mathrm{Apric}}$ corpus comprises 1,395 forum posts from 284 users crawled from the portal The Apricity - A European Cultural Community [7].
 To compile the $\mathcal{C}_{\mathrm{Apric}}$, we ensured that all posts within each verification case were collected from different sub-forums covering a variety of topics, including anthropology, genetics, race and society and ethno-cultural discussion.
@@
 6.1.7 $\mathcal{C}_{\mathrm{Trip}}$
-The $\mathcal{C}_{\mathrm{Trip}}$ corpus comprises reviews from 300 bloggers extracted from the Webis TripAdvisor Corpus 2014 published by Trenkmann and Spiel [107].
-As the original corpus was not intended for the purposes of AV, we reformatted the data accordingly.
+The $\mathcal{C}_{\mathrm{Trip}}$ corpus consists of reviews from 300 bloggers, which were extracted from the Webis TripAdvisor Corpus 2014, compiled by Trenkmann and Spiel [125].
+Given that the original corpus was not intended for AV purposes, we reformatted the data accordingly.
 While inspecting the data, we found that most of the included reviews were already of reasonable quality, so we only carried out moderate pre-processing.
 Among other things, we adjusted spaces next to punctuation marks (e.
-g., “( $\rightarrow$(alight”)) and normalized many character repetitions to a maximum of three characters (e.
+g., “( $\rightarrow$(alight”) and normalized many character repetitions to a maximum of three characters (e.
 g., “(!!!!$\rightarrow$!!!”).
@@
 6.1.8 $\mathcal{C}_{\mathrm{Yelp}}$
-The $\mathcal{C}_{\mathrm{Yelp}}$ corpus comprises 2,400 reviews written by 400 so-called “Yelplers” extracted from the well-known Yelp dataset [^11] https://www.yelp.com/dataset.
+The $\mathcal{C}_{\mathrm{Yelp}}$ corpus comprises 2,400 reviews written by 400 so-called “Yelplers” extracted from the well-known Yelp dataset [137].
 The corpus was developed with several challenges in mind.
@@
 Fraud.
-Lies.”) as well as review-type summaries such as “Quality: 10/10; Service: 9/10; Ambience/Location: 10/10; Overall: 9.5/10”.
+Lies.”) as well as review-type summaries such as “Quality: 10/10; Service: 9/10;” “Ambience/Location: 10/10; Overall: 9.5/10”.
 6.1.9 $\mathcal{C}_{\mathrm{IMDB}}$
@@
 6.1.10 $\mathcal{C}_{\mathrm{Blog}}$
-The $\mathcal{C}_{\mathrm{Blog}}$ corpus comprises 6,000 documents written by 1,500 bloggers, taken from the well-known Blog Authorship Corpus released by Schler et al.
+The $\mathcal{C}_{\mathrm{Blog}}$ corpus comprises 6,000 documents written by 1,500 bloggers, taken from the well-known Blog Authorship Corpus, released by Schler et al.
 [110].
@@
 The $\mathcal{C}_{\mathrm{AllNews}}$ corpus consists of 4,438 news articles written by 2,219 authors.
-The articles stem from the large well-known All the News 2.0 dataset [^12] The dataset can be accessed at https://components.one/datasets/all-the-news-2-news-articles-dataset, which consists of 2.7 million news articles and essays from 27 American news agencies and podcasting companies.
+The articles stem from the large well-known portal All the News 2.0 [35], which consists of 2.7 million news articles and essays from 27 American news agencies and podcasting companies.
 From these, we limit ourselves to a single source, namely Reuters, which contains the largest number of news articles (840,094 in total).
@@
 Next, we sorted all articles of each author by their publication year and kept only four articles (the two oldest and the two newest).
-The main reasons for this were firstly to maximize the stylistic variation of each author and secondly to ensure sufficient length of the resulting documents, which in some cases were massively shortened by the pre-processing.
-Based on the four documents per author, the pre-processing was performed as follows.
+The main reasons for this were, firstly, to maximize the stylistic diversity of each author and, secondly, to ensure a sufficient length of the resulting documents, which in some cases were heavily shortened due to the pre-processing.
+Based on the four documents per author, the pre-processing was carried out as follows.
 First, we removed all quotations (which are very common due to the nature of news articles) to mitigate the degree of stylistic distortion.
@@
 Unlike the Stack Exchange network, All the News 2.0 does not provide metadata indicating which specific parts of the original article have been edited.
-Therefore, it was not possible for us to remove these afterwards.
+Therefore, it was not possible for us to remove these retrospectively.
 6.2 Topic-Masking via POSNoise
+Figure 5: The POSNoise algorithm. Details on the notation can be found in [53].
 Authorship Verification constitutes a similarity detection problem in which the subject of the similarity determination is the language of the author rather than other document aspects such as the topic [53].
-However, a large proportion of existing AV methods including [1, 17, 19, 22, 69, 70, 72, 80, 88] use features that are directly influenced by the topic and thus run the risk of biasing their predictions accordingly.
-To counteract the problem of biased predictions regarding LambdaG, which also employs topic-affected features, we apply topic masking to all documents in our corpora.
-Topic masking is an approach to pre-process texts in such a way that topic-dependent text units (e.
-g., words, phrases, etc.) are masked by appropriate placeholders, while grammatical text units (e.
+However, a large number of existing AV methods including [1, 23, 25, 29, 83, 84, 86, 95, 102] use features that are directly influenced by the topic and thus run the risk of biasing their predictions accordingly.
+To counteract the problem of biased predictions regarding LambdaG, which also employs topic-affected features, we apply topic masking to all involved documents (see Algorithm 1).
+Topic masking is an approach to pre-process texts in such a way that topic-dependent text units such as words or phrases are masked by appropriate placeholders, while grammatical text units (e.
 g., determiners, conjunctions, contractions, prepositions, pronouns, auxiliary verbs, quantifiers, transitional phrases, etc.) are retained within the texts.
 Of the few topic masking approaches published in the literature, we chose POSNoise presented by Halvani and Graner [53], which proved to outperform the previously published topic masking approach called TextDistortion [119] in 34 out of 42 cases, with an increase in Accuracy of up to 10%.
+Category Examples
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 6: All categories of topic agnostic words and phrases whitelisted by POSNoise with some examples.
 Corpus Original / topic-masked sentences
@@
 I got a @ # that was @ to # #
-Table 5: Comparison between the original texts and their resulting topic-masked representations generated by POSNoise [42, Algorithm: 1].
-The semantics of the placeholders can be looked up in [42, Table: 2].
-5.3 N-Gram Language Modelling with Kneser-Ney Smoothing
+Table 7: Comparison between the original texts and their resulting topic-masked representations generated by POSNoise [53, Algorithm: 1].
+The semantics of the placeholders ©, Ø, @, # and § can be looked up in [53, Table: 2].
+6.3 N-Gram Language Modeling with Kneser-Ney Smoothing
 This section provides a self-contained introduction to (modified) Kneser-Ney smoothing [76, 32].
@@
 In plain words, these functions provide the number of different prefixes or suffixes a $n$-gram has been observed with exactly, or at least, $r$ times.
-Notice that all definitions automatically extend to the case where $g=\emptyset$, and we will use the shorthand notations $N_{x}(\bullet)\equiv N_{x}(\emptyset\bullet)$ and $N_{x}(\bullet\bullet)\equiv N_{x}(\bullet\emptyset\bullet)$.
+Notice that all definitions automatically extend to the case where $g=\emptyset$, and we will use the shorthand notations $N_{r}(\bullet)\equiv N_{r}(\emptyset\bullet)$ and $N_{r}(\bullet\bullet)\equiv N_{r}(\bullet\emptyset\bullet)$.
 The juxtaposition of the empty 00-gram to other $n$-grams is understood as no operation, e.
@@
 $P(\mathcal{D}_{m}|m)=\prod_{i=1}^{m}P(S\mbox{}_{i})$ (13)
-where, in turn, the probability of a sentence $S\mbox{}_{i}=t\mbox{}_{i,1}t\mbox{}_{i,2}\cdots t\mbox{}_{i,n_{i}}$ is:
-$P(S\mbox{}_{i})=\prod_{j=1}^{n_{i}+1}P(t\mbox{}_{i,j}|t\mbox{}_{i,j-N+1}t\mbox{}_{i,j-N+2}\cdots t\mbox{}_{i,j-1})$ (13)
-and we make the conventions $t\mbox{}_{i,n_{i}+1}=\langle\text{EOS}\rangle$, $t\mbox{}_{i,-1}=t\mbox{}_{i,-2}=\cdots=t\mbox{}_{i,-N+1}=\langle\text{BOS}\rangle$.
-A detail that is often overlooked in the literature is that the probability in Eq.
+where, in turn, the probability of a sentence $S\mbox{}_{i}=t\mbox{}_{i,1}t\mbox{}_{i,2}\cdots t\mbox{}_{i,z_{i}}$ is:
+$P(S\mbox{}_{i})=\prod_{j=1}^{z_{i}+1}P(t\mbox{}_{i,j}|t\mbox{}_{i,j-N+1}t\mbox{}_{i,j-N+2}\cdots t\mbox{}_{i,j-1})$ (14)
+and we make the conventions $t\mbox{}_{i,z_{i}+1}=\langle\text{EOS}\rangle$, $t\mbox{}_{i,-1}=t\mbox{}_{i,-2}=\cdots=t\mbox{}_{i,-N+1}=\langle\text{BOS}\rangle$.
+As a parenthetical remark, we observe that the probability in Eq.
 (13) is conditional on the number $m$ of sentences that compose the document $\mathcal{D}_{m}$, as reflected by our notation.
@@
 $p_{\text{KN}}(t\mbox{}|\emptyset)$ $=\alpha(t\mbox{}|\emptyset)+\gamma(\emptyset)\cdot\dfrac{1}{|\mathcal{T}\mbox{}|+1}.$ (16)
-The functions $\alpha$ and $\gamma$ are defined by:
+The functions $\alpha(\cdot)$ and $\gamma(\cdot)$ are defined by:
 $\alpha(t\mbox{}|g)=\begin{cases}\dfrac{\max(c_{\text{KN}}(gt\mbox{})-D(c_{\text{KN}}(gt\mbox{})),\,0)}{\sum_{t\mbox{}^{\prime}\in\mathcal{T}\mbox{}^{*}}c_{\text{KN}}(gt\mbox{}^{\prime})}&\text{if }c(g)>0,\\ 0&\text{otherwise},\end{cases}$ (17)
@@
 $c_{\text{KN}}(t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})=\begin{cases}c(t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})&\text{if }n=N,\\ N_{1+}(\bullet t\mbox{}_{1}t\mbox{}_{2}\cdots t\mbox{}_{n})&\text{otherwise}.\end{cases}$ (19)
-Finally, $D(1),\ D(2),\,\dots$ are free parameters satisfying $10<D(k)<$ 1, whose optimal values can be estimated from data.
+Finally, $D(1),\,D(2),\ldots$ are free parameters satisfying $10<D(k)<$ 1, whose optimal values can be estimated from data.
 The parameter space is usually made finite by assuming $D(k)=D(r)$ for some $r$ and $k\geq r$, in which case we can rewrite the infinite sum in (18) as:
@@
 In our work we use $r=3$, as suggested by the original reference [32]. We conclude the Section with a few remarks:
+  • As shown in Eq.
+(16), the $1$-gram distribution is interpolated with the uniform distribution over the extended dictionary $\mathcal{T}\mbox{}^{*}$ (see next point).
+This is relevant to our application, since documents featuring tokens not seen during training would otherwise be assigned zero probability, i.e.
+infinite perplexity.
+  • The conditional probability $p_{\text{KN}}(t\mbox{}|g)$ is defined for $t\mbox{}\in\mathcal{T}\mbox{}^{*}$, which is the reason why the denominator in the second term of (16) is $|\mathcal{T}\mbox{}|+1$.
+In particular $p_{\text{KN}}(\langle\text{EOS}\rangle|g)$ gives the probability that the $N-1$ gram $g$ terminates the sentence, whereas $p_{\text{KN}}(\langle\text{BOS}\rangle|g)$ is not defined (and indeed has no meaningful interpretation).
   • The modified count (19) treats differently $n$-grams of top and lower orders, respectively.
@@
 g., $\frac{N_{1+}(\bullet t\mbox{})}{N_{1+}(\bullet\bullet)}$ estimates the probability that, out of all distinct bigrams $t\mbox{}_{1}t\mbox{}_{2}$ found in a text corpus, the second token $t\mbox{}_{2}=t\mbox{}$.
-  • The conditional probability $p_{\text{KN}}(t\mbox{}|g)$ is defined for $t\mbox{}\in\mathcal{T}\mbox{}^{*}$, which is the reason why the denominator in the second term of (15) is $|\mathcal{T}\mbox{}|+1$.
-In particular $p_{\text{KN}}(\langle\text{EOS}\rangle|g)$ gives the probability that the $N-1$ gram $g$ terminates the sentence, whereas $p_{\text{KN}}(\langle\text{BOS}\rangle|g)$ is not defined (and indeed has no meaningful interpretation).
   • Kneser-Ney smoothing as originally formulated [76] uses a constant discount function $D(k)\equiv D$, corresponding to $r=1$ in the present formulation.
@@
   • The Prosecution Hypothesis ($H_{p}$): The disputed document ($\mathcal{D}_{\mathcal{U}}$) was written by the author of the known document(s) (the defendant, $\mathcal{A}$).
-  • the Defence Hypothesis ($H_{d}$): The disputed document ($\mathcal{D}_{\mathcal{U}}$) was not written by the author of the known document(s) (the defendant, $\mathcal{A}$), but was written by someone else, $\neg\mathcal{A}$
+  • The Defense Hypothesis ($H_{d}$): The disputed document ($\mathcal{D}_{\mathcal{U}}$) was not written by the author of the known document(s) (the defendant, $\mathcal{A}$), but was written by someone else, $\neg\mathcal{A}$
 The police’s or stakeholders’ interest in the given evidence ($E$), comprising $\mathcal{D}_{\mathcal{U}}$ and $\mathcal{D}_{\mathcal{A}}$, is probabilistically expressed as the ratio of the following two Bayesian conditional probabilities:
@@
 Equation 21 states that posterior odds can be estimated as the product of the prior odds and the likelihood ratio.
-The prior odds represents the trier-of-fact’s belief about the two competing hypotheses ($H_{p}$ and $H_{d}$) before observing the evidence (i.e., linguistic text evidence).
+The prior odds represents the trier-of-fact’s belief about the two competing hypotheses ($H_{p}$ and $H_{d}$) before observing the evidence (i.
+e., linguistic text evidence).
 The forensic scientist’s task is to evaluate the strength of the evidence, quantified in the form of a likelihood ratio.
@@
 They cannot estimate it logically because they are not privy of the trier-of-fact’s belief regarding the hypotheses and do not have access to other pieces of evidence presented prior to the linguistic text evidence.
-Legally, they must not refer to it because the posterior odds represents the ultimate issue of being guilty or innocence [73, p.
+Legally, they must not refer to it because the posterior odds represents the ultimate issue of being guilty or innocence [88, p.
 96] and this determination is a specific responsibility reserved only for the trier-of-fact.
@@
 Metrics such as precision, recall, and equal error rate (EER) are typical in this regard.
-While these metrics offer a quick overview of performance and are intuitively easy to comprehend, they are not suitable for assessing likelihood ratio-based inference systems [78, p.
+While these metrics offer a quick overview of performance and are intuitively easy to comprehend, they are not suitable for assessing likelihood ratio-based inference systems [93, p.
 93].
@@
 Firstly, metrics relying on categorical thresholding implicitly assess how effectively the system correctly identifies the source of evidence.
-In other words, these metrics implicitly pertain to the strength of the hypothesis rather than the strength of the evidence [78, p.
+In other words, these metrics implicitly pertain to the strength of the hypothesis rather than the strength of the evidence [93, p.
 93].
-Secondly, categorical metrics fail to capture the gradient nature of likelihood ratios [78, p.
+Secondly, categorical metrics fail to capture the gradient nature of likelihood ratios [93, p.
 93].
-For instance, in a same-author comparison, if one system returns $\log_{10}(\text{LR})=3$ while another yields $\log_{10}(\text{LR})=2$, the former is superior as it supports the correct hypothesis more strongly.
+For instance, in a same-author comparison, if one system returns $\log_{10}(\text{\emph{LR}})=3$ while another yields $\log_{10}(\text{\emph{LR}})=2$, the former is superior as it supports the correct hypothesis more strongly.
 However, using a categorical metric would assess both systems in the same way, disregarding the differing strengths of evidence supporting the correct hypothesis.
 The conventional metric employed for evaluating the performance of a system within the Likelihood Ratio Framework is the log-likelihood ratio cost ($C_{\text{llr}}\mbox{}$).
-The computation of the $C_{\text{llr}}\mbox{}$ involves utilising a set of LRs for same-source scenarios (e.g., same-author, Y) and a set for different-source scenarios (e.
-g., different-author, N).
+The computation of the $C_{\text{llr}}\mbox{}$ involves utilizing a set of LRs for same-source scenarios (e.
+g., same-authorship, Y) and a set for different-source scenarios (e.
+g., different-authorship, N).
 The computation is performed using the following equation [26, 104, 131]
-$C_{\text{llr}}\mbox{}=\frac{1}{2}\left(\left[\frac{1}{n}\sum_{i}^{n}\log_{2}\left(1+\frac{1}{LR_{i}}\right)\right]+\left[\frac{1}{m}\sum_{j}^{m}\log_{2}\bigg{(}1+LR_{j}\bigg{)}\right]\right)$ (21)
-where, $n$ and $m$ are the number of Y- and N-cases, and $LR_{i}$ and $LR_{j}$ are the LRs estimated for the Y- and N-cases, respectively.
-$C_{\text{llr}}\mbox{}$ considers the magnitude of calculated LR values, imposing appropriate penalties based on their values.
-Within the $C_{\text{llr}}\mbox{}$ metric, LRs supporting counter-factual hypotheses, or contrary-to-fact LRs (with $\log_{10}(\text{LR})<0$ for Y-problems and $\log_{10}(\text{LR})>0$ for N-problems), incur significant penalties.
-The extent of the penalty is proportional to the deviation of the LRs from unity.
-Optimal performance is attained when $C_{\text{llr}}\mbox{}$ equals 0, and performance gradually diminishes as $C_{\text{llr}}\mbox{}$ approaches and surpasses 1.
-Therefore, a lower $C_{\text{llr}}\mbox{}$ value corresponds to better performance.
+$C_{\text{llr}}\mbox{}=\frac{1}{2}\left(\frac{1}{\gamma}\sum_{i}^{\gamma}\log_{2}\left(1+\frac{1}{\emph{LR}_{i}}\right)+\frac{1}{\eta}\sum_{j}^{\eta}\log_{2}\bigg{(}1+\emph{LR}_{j}\bigg{)}\right)$ (22)
+[…]
 $C_{\text{llr}}\mbox{}$ serves as an assessment metric for the overall performance of an LR-based inference system and can be decomposed into two distinct metrics: $C_{\text{llr}}^{\text{min}}\mbox{}$ and $C_{\text{llr}}^{\text{cal}}\mbox{}$.
 The former evaluates the discrimination performance and the latter the calibration performance of the system.
-The $C_{\text{llr}}^{\text{min}}\mbox{}$ is determined by computing the losses for the optimised LRs through the non-parametric pool-adjacent-violators algorithm.
+The $C_{\text{llr}}^{\text{min}}\mbox{}$ is determined by computing the losses for the optimized LRs through the non-parametric pool-adjacent-violators algorithm.
 The disparity between $C_{\text{llr}}\mbox{}$ and $C_{\text{llr}}^{\text{min}}\mbox{}$ is expressed as $C_{\text{llr}}^{\text{cal}}\mbox{}$; in other words, $C_{\text{llr}}\mbox{}$ equals the sum of discrimination loss ($C_{\text{llr}}^{\text{min}}\mbox{}$) and calibration loss ($C_{\text{llr}}^{\text{cal}}\mbox{}$).
@@
 The $C_{\text{llr}}\mbox{}$ has already been effectively utilized for evaluating LR-based forensic text comparison systems [62, 63, 64, 97].
-5.5 SiamBERT: Neural Net Baseline
-Recognising the paradigm shift towards neural network techniques that is also evident in the research field of AV, we consider it important to compare LambdaG with contemporary methods utilizing deep learning techniques.
-For this reason, we incorporate in our evaluation a baseline method based on a Siamese BERT neural network and functionally very similar to the Sentence-BERT architecture [93] and Ishihara’s et al [53] method, in which likelihood-ratios were estimated via cosine similarity and probabilistic linear discrimination analysis.
-We call this method from here on SiamBERT.
-Note that, in our experiments, we use the BERT-Large, Cased model with its corresponding tokenizer [30], which are pretrained on English text and considers character casing.
-Given a document $\mathcal{D}$, we tokenize it into a sequence of tokens.
-Since the BERT model we are using is only able to handle up to 512 tokens, we split the sequence of tokens into a set of adjacent chunks $\varsigma_{1},\ldots,\varsigma_{m}$, where $m$ is the number of chunks.
-Each $\varsigma_{i}$ comprises 512 tokens, except for $\varsigma_{m}$, which may comprise less.
-These chunks are then embedded as feature vectors $F_{1},\ldots,F_{m}$, by encoding each $\varsigma_{i}$ with the BERT model and taking the average of the resulting embedding vectors over the tokens in a chunk.
-Finally, the feature vectors are then again averaged, resulting in a final feature vector $F$, which represents the document $\mathcal{D}$ in an embedding space.
-Given a single verification case, we consider an unknown document $\mathcal{D}_{\mathcal{U}}$ and a known document $\mathcal{D}_{\mathcal{A}}$ for SiamBERT.
-Both are then embedded as described above, resulting in two feature vectors $F_{\mathcal{U}}$ and $F_{\mathcal{A}}$.
-Based on these vectors, we then calculate their cosine similarity $s$.
-If $s>\theta$, where $\theta$ is a threshold, we assume $\mathcal{U}=\mathcal{A}$ (i.
-e., same authorship).
-To train SiamBERT, we further split the training set $\mathcal{C_{\,\textrm{train}}}$ into $\mathcal{C_{\,\textrm{train}*}}$ and $\mathcal{C_{\,\textrm{val}}}$ based on a 80/20 ratio, in order to optimize hyper-parameters.
-We trained the parameters of the BERT model on $\mathcal{C_{\,\textrm{train}}}$ with MSE loss function (calculating the mean squared error between the cosine similarity and the binary label, where 1 represents a Y case), the AdamW optimizer, a batch size of $4$ and a learning rate of $0.00005$ for $100$ epochs (scheduled with a linear warmup in the first $10$ epochs and a cosine decay to zero in the remaining).
-In order to optimize the hyper-parameter epoch count, we measured the AUC on each epoch of the corresponding model on $\mathcal{C_{\,\textrm{val}}}$ and selected the model with the best AUC.
-Finally, we calculated the cosine similarities of the verification cases of $\mathcal{C_{\,\textrm{train}}}$ using the selected model and determined the optimal threshold $\theta$, which would yield the best accuracy on $\mathcal{C_{\,\textrm{train}}}$.
-6 Appendix A: Qualitative Data Exploration
+6.5 Baseline AV Methods
+For the evaluation of LambdaG, we selected seven baseline AV methods that were covered in our literature review (see Section 1).
+Specifically, we chose IM [84], COAV [57], TAVeer [55], FeVecDiff (LR [133] + MLP [132]) and ADHOMINEM [19].
+The selection of these methods was guided by several key considerations.
+ADHOMINEM, FeVecDiff (LR + MLP) and TAVeer were the three top approaches in the PAN AV 2020 competition [70].
+Furthermore, FeVecDiff (LR) ranked third in the PAN AV 2021 competition [71], closely followed by the second-placed method.
+Regarding IM and COAV, Halvani [51] demonstrated that, across ten different corpora, IM achieved the highest average performance, while COAV ranked second (both in terms of Accuracy and AUC) among twelve evaluated AV methods.
+COAV has also attracted attention in the AV community.
+In independent studies, Halvani et al.
+[57] and Bevendorff et al.
+[15] came to the conclusion that COAV is on par with Bagnall’s well-known RNN-based AV method [12], the winning approach of the PAN AV 2015 competition [120].
+Moreover, COAV was selected as one of six baseline methods in the latest PAN AV 2024 competition, where it outperformed four of them [16, Table 5].
+In addition to these six (published) baselines, we have also decided to implement an additional baseline that is used as a common conceptual AV approach in the literature (e.
+g., [5, 135]).
+This baseline is described in Section 6.6.
+Note that with respect to COAV, TAVeer, FeVecDiff (LR + MLP) and ADHOMINEM, we used the original implementations [^11] The original implementations were downloaded from the following Github repositories:
+ADHOMINEM $\rightarrow$ https://github.com/boenninghoff/AdHominem
+FeVecDiff (MLP) $\rightarrow$ https://github.com/janithnw/pan2020_authorship_verification
+FeVecDiff (LR) $\rightarrow$ https://github.com/janithnw/pan2021_authorship_verification
+of the authors.
+Regarding IM, we have used the same implementation as used in [51].
+The hyperparameter optimization for FeVecDiff (LR + MLP) and ADHOMINEM was performed by splitting each training corpus $\mathcal{C_{\,\textrm{train}}}$ into a training subset $\mathcal{C_{\,\textrm{train}*}}$ and a validation set $\mathcal{C_{\,\textrm{val}}}$, following an 80/20 ratio.
+6.6 SiamBERT
+As briefly explained in the previous section, we decided to add another baseline method to our evaluation, which we refer to as SiamBERT.
+For this purpose, we have implemented an AV method based on a Siamese BERT neural network with a structure similar to the well-known sentence BERT architecture [106].
+SiamBERT has a similar functionality to the method of Ishihara et al.
+[66], in which likelihood-ratios are estimated via cosine similarity and probabilistic linear discrimination analysis.
+SiamBERT is based on the BERT-Large, Cased [^12] https://huggingface.co/google-bert/bert-large-cased model with the corresponding tokenizer [37], which are pre-trained on English texts and take into account character-sensitive capitalization.
+For a given document $\mathcal{D}$, we first tokenize it into a sequence of tokens.
+Since our BERT model is constrained to processing a maximum of 512 tokens per input, we segment the token sequence into a set of adjacent chunks $\varsigma_{1},\ldots,\varsigma_{m}$, where each chunk $\varsigma_{i}$ consists of 512 tokens, except for the final chunk, which may contain fewer.
+These chunks are then embedded as feature vectors, by encoding each $\varsigma_{i}$ with the BERT model and taking the average of the resulting token-level embedding vectors over the tokens in a chunk.
+Finally, we obtain the document’s overall representation $\mathcal{F}\mbox{}$ by averaging these $m$ chunk-level feature vectors, embedding the document $\mathcal{D}$ in a continuous vector space.
+Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, we first embed both documents into their corresponding feature vectors $\mathcal{F}\mbox{}_{\mathcal{U}}$ and $\mathcal{F}\mbox{}_{\mathcal{A}}$ (as described above) and then compute their cosine similarity $s$.
+Finally, $c$ is classified as either Y (i.
+e., $\mathcal{A}=\mathcal{U}$) if $s>\theta$, otherwise as N (i.
+e., $\mathcal{A}\neq\mathcal{U}$), where $\theta$ is a predefined threshold.
+To train SiamBERT, we first split the training corpus $\mathcal{C_{\,\textrm{train}}}$ into $\mathcal{C_{\,\textrm{train}*}}$ and $\mathcal{C_{\,\textrm{val}}}$ based on an 80/20 ratio to optimize the hyperparameters.
+Based on $\mathcal{C_{\,\textrm{train}*}}$, we then train the model parameters using the MSE loss function (calculating the mean squared error between the cosine similarity and the binary label, where 1 represents a Y-case), the AdamW optimizer, a batch size of $4$ and a learning rate of $0.00005$ for $100$ epochs (scheduled with a linear warm up in the first $10$ epochs and a cosine decay to zero in the remaining).
+To optimize the hyperparameter for the epoch count, we evaluated the AUC at each epoch on $\mathcal{C_{\,\textrm{val}}}$ and selected the model with the highest AUC.
+Using this selected model, we then computed the cosine similarities for the verification cases in $\mathcal{C_{\,\textrm{train}}}$ and determined the optimal threshold $\theta$ that maximized Accuracy on $\mathcal{C_{\,\textrm{train}}}$.
+7 Qualitative Data Exploration
+This Section contains a series of examples of how an analyst can visualize the results of LambdaG.
+To aid readability, we replaced the original placeholders used in the POSNoise paper (see [53, Table 2]) within the visualizations with their corresponding POS tags, according to the following mapping:
+[…]
+[…]
+[…]
+[…]
+[…]
 […]
@@
 This token is marked as very important because the overarching sequence that identifies this author is the use of “can’t” in combination with “so” with an intervening pronoun, a sequence only found in one other author in the corpus.
+8 Acknowledgments
+This research work was supported by the National Research Center for Applied Cybersecurity ATHENE.
+ATHENE is funded jointly by the German Federal Ministry of Education and Research and the Hessian Ministry of Higher Education, Research and the Arts.
+We thank Dmitry Nikolaev, Łukasz Dębowski, Petar Milin, Ricardo Bermúdez-Otero, and Richard Zimmermann for their useful comments and feedback.
+The authors acknowledge that they received no funding in support for this research.
+Author contributions:
+  • Conceptualization: AN
+  • Methodology: AN, OH, LG, VG
+  • Implementation: AN, LG, VG, OH
+  • Validation: LG, VG
+  • Formal analysis: LG, VG, SI
+  • Investigation: AN, LG, OH
+  • Resources: OH, LG
+  • Data Curation: OH, LG
+  • Writing—original draft: AN, OH, LG, VG
+  • Writing—review and editing: AN, OH, LG, VG, SI
+  • Visualization: LG, OH
+  • Supervision: AN
+  • Project Administration: AN
+Authors declare that they have no competing interests.